### v3

# Structured Generative Modeling with The Thermodynamic Kolmogorov-Arnold Model

Abstract
Learning an energy-based model (EBM) in the latent space of a top-down generative model offers an expressive and interpretable framework for text and image generation. However, it remains unclear how this interpretability can be used to guide model design, improve generative quality, and reduce training time. Moreover, the reliance on Langevin Monte Carlo (LMC) sampling presents challenges in efficiency and sampling multimodal latent distributions. In this work, we propose a novel adaptation of the Kolmogorov–Arnold representation theorem for generative modeling and introduce the Thermodynamic Kolmogorov–Arnold Model (T-KAM) to take advantage of structural and inductive biases. By constraining the prior to univariate relationships, T-KAM enables fast and exact inference via the inverse transform method. With the low dimensionality of the latent space and suitable inductive biases encoded, we demonstrate that importance sampling becomes a viable, unbiased, and highly efficient training strategy. We also introduce a training criterion using population-based LMC, which decomposes posterior sampling into a sequence of annealed distributions to improve multimodal exploration. T-KAM elegantly balances common trade-offs in generative modeling, offering fast inference, high sample quality, and stable training, while being naturally suited to upcoming Zettascale Computing Co. hardware and extendable to other high-impact research directions in generative intelligence.
1 Introduction
Deep generative modeling is a core subfield of machine learning that enables systems to learn complex data distributions, generate novel data, and support a wide range of tasks including simulation, compression, and representation learning. Two prominent and widely studied approaches are Variational Autoencoders (VAEs) (Kingma & Welling, 2022) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). Both frameworks assume that each data point is generated from a low-dimensional latent vector via a top-down neural network, where the latent vector is typically drawn from a simple, non-informative prior, such as a uniform or isotropic Gaussian distribution.
Latent energy-based prior models (EBMs) (Pang et al., 2020) extend this framework by learning an expressive, data-dependent prior directly in the latent space. This prior can be trained jointly with the top-down generator using maximum likelihood estimation (MLE), where the EBM learns by the statistical discrepancy between samples from the prior and posterior distributions, following a paradigm similar to Empirical Bayes. Each learning iteration requires Markov Chain Monte Carlo (MCMC) sampling from both the prior and posterior, specifically the Unadjusted Langevin Algorithm (ULA) (Brooks et al., 2011; Rossky et al., 1978b; Roberts & Stramer, 2002), which offers greater modeling flexibility compared to the amortized inference methods of VAEs.
Despite being extendable to several high-impact research directions, such as multimodal generative modeling (Yuan et al., 2024), and interpretable classification and meta-learning (Kong et al., 2021), (which involves rapid adaptability and few-shot generalization to out-of-distribution tasks), latent EBMs face critical limitations due to their reliance on MCMC sampling.
Firstly, MCMC is inherently iterative, and algorithms such as ULA require repeated gradient evaluations through the model, introducing significant computational overhead during both training and inference. Secondly, ULA is best suited for sampling from smooth, unimodal distributions and tends to struggle when exploring the multimodal latent distributions that are typical in latent variable generative modeling. Moreover, separate ULA samplers are required for prior and posterior sampling, which is problematic given that ULA is sensitive to hyperparameter tuning: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias due to its use of the Euler–Maruyama method to evolve the underlying Langevin equation.
Additionally, it is not evident how the unique separation between the prior and generator networks can be leveraged for interpretability or training efficiency. Unlike the framework introduced by Karniadakis et al. (2021), which describes how scientific inductive and learning biases can be embedded into neural networks to reduce their reliance on data and improve training efficiency, generalization, and explainability, (all of which are critical for ensuring trust in Artificial Intelligence), it remains unclear how the EBM prior can be effectively aligned with the posterior to achieve similar effects. Although ideas from mechanistic interpretability, as reviewed by Bereska & Gavves (2024), could potentially be applied, there is currently no established method to leverage the unique strengths of the EBM prior framework for this purpose.
1.1 Inductive biases and importance sampling
To address these limitations, we reimagine the EBM prior framework into a novel univariate or element-wise model, drawing inspiration from Kolmogorov–Arnold Networks (KANs) (Liu et al., 2024) and deep spline networks Bohra et al. (2020), which replace the sigmoidal activations used in multilayer perceptrons Murtagh (1991) with spline interpolations. By representing the latent space with univariate functions and distributions, we improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning.
This allows the EBM prior to be aligned with the posterior before training begins. We show that improved prior–posterior alignment enables efficient training with importance sampling (IS) (Goertzel, 1949), a technique for estimating posterior expectations that is otherwise avoided in the literature due to its degeneracy under severe prior–posterior mismatch and the curse of dimensionality, (Dupuis & Wang, 2004). Moreover, the trained prior can be visualized or recovered, offering new opportunities for transfer learning and scientific discovery.
This is tested in Sec. 3.1. Multiple KAN bases and reference priors are used to verify the effect of inductive bias. However, Gaussian Radial Basis Functions (RBF KANs) (Li, 2024), were used instead of the cubic B-spline bases proposed by Liu et al. (2024), as RBFs are more efficient for GPU implementation, and Poggio & Girosi (1990) previously demonstrated that they provide an optimal solution to the regularized approximation/interpolation problem.
To derive our new model, which we call the Thermodynamic Kolmogorov-Arnold Model (T-KAM), we first introduce a novel adaptation of the Kolmogorov–Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces. Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on heuristics and automated tuning. This reinterpretation allows us to frame inverse transform sampling (ITS) (Devroye, 1986) as a measurable transformation applied to uniform noise, thus preserving the deterministic structure of KART while admitting a probabilistic interpretation that aligns with MLE.
To address the potential inflexibility of strict adherence to KART and the limitations of using independent energy-based priors, we introduce a scalable strategy based on mixture distributions. Beyond introducing dependencies between latent dimensions to capture inter-dimensional relationships, this also improves sampling efficiency by enabling component-wise sampling.
1.2 Population-based Langevin sampling and Thermodynamic training
However, while ITS enables fast inference, posterior sampling during training must still be conducted using ULA when the IS estimator is insufficient. Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by Cui & Han (2024) and the diffusion-assisted EBM by Zhang et al. (2023). To maintain prior interpretability and fast inference post-training, we instead introduce posterior annealing as a novel and alternative strategy. This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and sampling with population-based MCMC methods, specifically Parallel Tempering (PT) (Swendsen & Wang, 1986) (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996).
Although this may reduce expressivity compared to hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training. Our training criterion is then derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al. (2019) to ensure the impact of annealing on learning dynamics and parameter updates.
While ULA is computationally efficient and enables more flexible chain dynamics, it does not preserve detailed balance or reversibility. This can be recovered by incorporating Metropolis-Hastings (MH) adjustments (Metropolis & Ulam, 1949) via the Metropolis-adjusted Langevin Algorithm (MALA) (Rossky et al., 1978a). The codebase in Sec. A.1 optionally supports MALA with locally adaptive step-size tuning, (autoMALA) (Biron-Lattes et al., 2024), to assist when a single step size fails to accommodate the varying geometries of power posteriors, and when the MH criterion deteriorates under the curse of dimensionality, (Hairer et al., 2014).
MALA and autoMALA are discussed under Secs. A.6.3 and A.6.4. However, ULA remains the preferred method in our work, motivated by its demonstrated efficiency and effectiveness by Pang et al. (2020), and further supported by the findings of Nijkamp et al. (2019b; a), with regards to how short-run, biased MCMC methods can still be highly effective for sampling from EBMs.
These scaling strategies are compared in Sec. 3.2. Despite our optimizations, the element-wise and localized nature of KAN functions complicate their parallelization on GPUs, which are best suited to matrix arithmetic. As a result, forward passes through T-KAM and the gradient evaluations required for ULA were found to be slow, necessitating reduced experimental complexity.
We ultimately found that our mixture prior model outperformed a deep KAN prior architecture closely aligned with the model used in the study by Pang et al. (2020), while also offering greater computational efficiency. In contrast, the use of annealing could not be justified due to its significant computational overhead and limited gains in quality.
However, this assessment may change as a greater number of power posteriors and more sophisticated posterior sampling algorithms become feasible with the emerging hardware advances proposed by Zetta (2024). Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting T-KAM at scale, while improving the potential of univariate priors with its function fusion features.
We present T-KAM as an initial step to inspire further adaptations of KART in machine learning. This work aims to lay the groundwork for more exploration amongst the research community, with the far-reaching hope of eventually advancing towards a broader conceptualization: “The Kolmogorov-Arnold Representation Theorem Is All You Need.“
2 The Thermodynamic Kolmogorov-Arnold Model
This section introduces the model and its associated sampling procedures. We begin by presenting a novel interpretation of the Kolmogorov–Arnold Representation Theorem (KART) to define the model’s baseline architecture. Our key practical contribution is the introduction of a univariate energy-based prior, detailed in Sec. 2.1. The top-down generator can be instantiated in various forms, including Kolmogorov–Arnold Networks (KANs), which enables strict adherence to KART.
We then propose scalable extensions to the model, including mixture distributions in Sec. 2.3.2, which introduce dependencies between latent dimensions. This represents a novel contribution to the literature on EBM priors. The final sections address prior and posterior sampling: we introduce inverse transform sampling (ITS) as a new approach for efficient EBM prior sampling, and introduce the posterior annealing strategy in Sec. 2.5 to improve multimodal sampling during training.
The Kolmogorov–Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of T-KAM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
The Kolmgorov-Arnold Representation theorem 1
For any integer $n_{z}>1$, there are continuous univariate functions $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ and $\psi_{q,p}:[0,1]\to\mathbb{R}$, such that any continuous multivariate function, $g:[0,1]^{n_{z}}\to\mathbb{R}$, can be represented as a superposition of those functions:
$g(u_{1},\dots,u_{n_{z}})=\sum_{q=1}^{2n_{z}+1}\Phi_{q}\left(\sum_{p=1}^{n_{z}}\psi_{q,p}(u_{p})\right),$ (1)
Eq. 1 can provide structural bias and guide the design of T-KAM’s architecture. To ensure adherence to it, any transformations must be derived to be continuous, measurable, and deterministic. For readers less familiar with the technical background, the two most practically significant equations of Sec.. 2.1 are Eqs. 46.
2.1 Univariate energy-based prior
We interpret $u_{p}\sim\mathcal{U}(u_{p};\,0,1)$ in Eq. 1 as a uniform random variable in the probability space, $\left([0,1],\mathcal{B}([0,1]),\mu\right)$, where $\mathcal{B}\left([0,1]\right)$ is the Borel $\sigma$-algebra on $[0,1]$ and $\mu$ is the Lebesgue measure on $[0,1]$. To ensure that the transformation $u_{p}\mapsto\psi_{q,p}(u_{p})$ is well defined in a probabilistic setting, we require it to be a Markov kernel. For all Borel measurable sets $A\subseteq\mathcal{Z}$, where $\mathcal{Z}\subset\mathbb{R}$ is introduced as the latent space of T-KAM, the map
$u_{p}\ \mapsto\ K(u_{p},A)=\delta_{\psi_{q,p}(u_{p})}(A)$ (2)
is measurable. Here, $\delta_{\psi_{q,p}(u_{p})}$ denotes the Dirac measure centred at $\psi_{q,p}(u_{p})$, ensuring the kernel is deterministic: i.e., once $u_{p}$ is sampled, the outcome in $\mathcal{Z}$ is fixed. This measurability requirement guarantees that the pushforward from $([0,1],\mathcal{B}([0,1]),\mu)$ to $(\mathcal{Z},\mathcal{B}(\mathcal{Z}),\pi)$ is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$. Here, $\mathcal{B}(\mathcal{Z})$ is the Borel $\sigma$-algebra on the latent space, $\mathcal{Z}$, and $\pi$ is the latent measure.
The functions $\psi_{q,p}$ can therefore be conveniently interpreted as inverse cumulative distribution functions (CDFs) corresponding to independent univariate priors:
$\psi_{q,p}(u_{p})=F^{-1}(\pi_{q,p})(u_{p}).$ (3)
Here, $F^{-1}(\pi_{q,p}):[0,1]\to\mathcal{Z}$ is the generalized inverse CDF associated with a parameterized latent distribution, $\pi_{q,p}$. This is valid under standard regularity assumptions on $F$, since it is a measurable, monotone, increasing function, and thus $F^{-1}(u)=\inf\left\{z\in\mathcal{Z}:F(z)\geq u\right\}$ is a measurable mapping between the two probability spaces. We parameterize the latent distribution $\pi_{q,p}$ via exponential tilting of a base prior $\pi_{0}$:
$\pi_{q,p}(z)=\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),$ (4)
where $f_{q,p}(z)$ is a learned energy function and $Z_{q,p}$ is the normalizing constant (partition function) that ensures $\pi_{q,p}$ defines a valid probability. The inspiration for this exponential tilting form can be traced to the energy-based prior introduced by Pang et al. (2020), adapted to the univariate case. The base prior $\pi_{0}$ serves as a reference measure that guides the form of the learned distribution $\pi_{q,p}$, allowing for flexible reweighting via the energy function $f_{q,p}$.
For $\pi_{q,p}$ to define a valid probability distribution, it must satisfy countable additivity while being non-negative and normalized. This is guaranteed since $f_{q,p}:\mathcal{Z}\to\mathbb{R}$ is Borel-measurable and integrable with respect to the reference density $\pi_{0}$, ensuring the finiteness of the partition function:
$Z_{q,p}=\int_{\mathcal{Z}}\exp(f_{q,p}(z))\,d\pi_{0}(z)<\infty.$ (5)
In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov–Arnold Networks (KANs) (Liu et al., 2024). For instance, smoothing splines, such as those used by KANs or deep spline networks, (Bohra et al., 2020). Splines are typically defined on bounded domains, allowing the support of $\pi_{q,p}$ to be explicitly controlled by the user.
The partition function $Z_{q,p}$ is often intractable in energy-based models. However, in this univariate setting, where the energy function is smooth and defined on a bounded domain $\left[z_{\text{min}},z_{\text{max}}\right]$, it can be accurately approximated using Gauss–Kronrod quadrature, (Laurie, 1997):
$Z_{q,p}$ $=\int_{z_{\text{min}}}^{z_{\text{max}}}\exp(f_{q,p}(z))\pi_{0}(z)dz\approx\sum_{i=1}^{N_{\text{quad}}}w_{i},G(z_{i}^{\text{node}}),$ (6)
where $G(z)=\exp(f_{q,p}(z))\pi_{0}(z)$, and $\{z_{i}^{\text{node}},w_{i}\}_{i=1}^{N_{\text{quad}}}$ are the quadrature nodes and weights for the interval $\left[z_{\text{min}},z_{\text{max}}\right]$. The smoothness of $G$, together with its bounded support, makes this approximation highly accurate.
With an valid form for the normalized CDF, inverse transform sampling (ITS) can be used to draw samples from $\pi_{q,p}$ as deterministic transformations of uniform noise. This preserves the deterministic structure of KART while enabling flexible, learnable latent distributions. Inverse transform sampling remains effective even for functions that are not Riemann integrable, provided they are integrable in the Lebesgue sense, typical for the functions in KAN literature.
As a result, the model retains a coherent probabilistic interpretation of KART that is consistent with the energy-based framework proposed by Pang et al. (2020), which is well suited to maximum likelihood training. However, this approach is significantly limited, as it ignores inter-dimensional dependencies in the latent space and results in an axis-aligned joint distribution. We address this inflexibility in Sec. 2.3.2 by relaxing our adherence to KART.
2.2 Kolmogorov-Arnold Network generator
Having sampled the intermediary latent variable with ITS such that $z_{q,p}=F^{-1}(\pi_{q,p})(u_{p})$, the generator can be realized flexibly as $\mathcal{G}:\mathcal{Z}\to\mathcal{X}$:
$\bm{\tilde{x}}^{(s)}$ $=\mathcal{G}\left(\bm{z}^{(s)}\right)+\bm{\epsilon}^{(s)},\quad\text{where}\quad z_{q,p}^{(s)}\sim\pi_{q,p}(z),\quad\bm{\epsilon}^{(s)}\sim\mathcal{N}(\bm{\epsilon};\;\bm{0},\sigma_{\epsilon}^{2}\bm{I}),$ (7)
where $\mathbf{\tilde{x}}^{(s)}=\{\tilde{x}_{o}^{(s)}\}_{o=1}^{n_{x}}$ represents a generated data sample corresponding to a single latent sample indexed by $s$, and $n_{x}$ is used to specify the dimension of $\bm{\tilde{x}}^{(s)}$. The additive noise is a Gaussian sample with hyperparameter $\sigma_{\epsilon}$. However, the architecture may be made strictly adherent to KART using a Kolmogorov-Arnold Network:
$\tilde{x}_{o}^{(s)}$ $=\sum_{q=1}^{2n_{z}+1}\Phi_{o,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)+\epsilon_{o}^{(s)},\quad\mathbf{\tilde{x}}^{(s)}=\begin{pmatrix}\sum_{q=1}^{2n_{z}+1}\Phi_{1,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)\\[5.0pt] \sum_{q=1}^{2n_{z}+1}\Phi_{2,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)\\[5.0pt] \vdots\\[5.0pt] \sum_{q=1}^{2n_{z}+1}\Phi_{n_{x},q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)\end{pmatrix}+\bm{\epsilon}^{(s)}.$ (8)
The functions, $\Phi_{o,q}:\mathbb{R}\to\mathbb{R}$, are realizable as basis or KAN functions similar to $f_{q,p}:\mathcal{Z}\to\mathbb{R}$.
2.3 Scaling T-KAM
The trainable components of the prior and inner functions are shared across the features of each input $\bm{x}^{(b)}$, while the outer generator functions remain separate for each output dimension. Specifically, the learnable functions are defined as:
$\bm{f}$ $=\begin{pmatrix}\left(f_{1,1},\dots,f_{1,n_{z}}\right)\\ \left(f_{2,1},\dots,f_{2,n_{z}}\right)\\ \vdots\\ \left(f_{2n_{z}+1,1},\dots,f_{2n_{z}+1,n_{z}}\right)\end{pmatrix},\quad\bm{\Phi}=\begin{pmatrix}\left(\Phi_{1,1},\dots,\Phi_{1,2n_{z}+1}\right)\\ \left(\Phi_{2,1},\dots,\Phi_{2,2n_{z}+1}\right)\\ \vdots\\ \left(\Phi_{n_{x},1},\dots,\Phi_{n_{x},2n_{z}+1}\right)\end{pmatrix}$, (9)
where $n_{x}$ denotes the dimensionality of the observed data.
From this point forward, we formulate our derivations over a multivariate probability space $(\mathcal{Z},\mathcal{B}(\mathcal{Z}),P)$, where $\mathcal{Z}$ no longer represents a subinterval of the real line, but rather a subspace of $\mathbb{R}^{(2n_{z}+1)\times n_{z}}$. The measure $P$ defines a product distribution over the latent variables:
$P(\bm{z}\mid\bm{f})=\pi_{1,1}\otimes\cdots\otimes\pi_{2n_{z}+1,n_{z}},$
where each component $\pi_{j,k}$ is defined via exponential tilting, as introduced in Sec. 2.1. The full latent vector $\bm{z}$ is a matrix-valued random variable composed of the independent latent features.
2.3.1 Generator
The generator, $\mathcal{G}$, can be considered a separate generator network allowing for more flexible realizations, such as convolutional layers when adherence to Eq. 1 is not required. Deepening this generator, as guided by Liu et al. (2024) for the Kolmogorov-Arnold Network (KAN), does not replace KART as the innermost foundation of T-KAM:
$x_{o}^{(s)}=$ $\sum_{l_{n}=1}^{L_{n}}h_{o,l_{n}}\Bigg{(}\dots h_{l_{3},l_{2}}\Bigg{(}\sum_{l_{1}=1}^{L_{1}}h_{l_{2},l_{1}}\Bigg{(}\sum_{q=1}^{2n_{z}+1}\Phi_{l_{1},q}\Bigg{(}\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\Bigg{)}\Bigg{)}\Bigg{)}\Bigg{)}+\epsilon_{o}^{(s)}.$ (10)
Here, each $h$ represents a hidden layer introduced between the outer sum of Eq. 1, and the output feature space to allow for more expressivity when required. In fact, each successive instance of $h$ can be interpreted as a distinct outer sum of Eq. 1, where the inner sum corresponds to the previous layer. This is valid provided that:
$L_{1}$ $=2\left(2n_{z}+1\right)+1,\quad L_{2}=2L_{1}+1,\quad\dots\quad L_{n}=2L_{n-1}+1.$ (11)
Thus, a smooth transition between the latent space and data space can be achieved by progressively doubling layer widths, eliminating the need for arbitrary choices while maintaining relaxed compliance with KART. Strict adherence to KART can be achieved for hidden layer $h_{l_{k},l_{k-1}}$ by applying either a sigmoid or affine transformation after layer $h_{l_{k-2},l_{k-3}}$ to shift its range back to $0,1]^{L_{k-1}}$. However, using sigmoid may lead to vanishing gradients and we contend that function domains are less critical to KART than its structural properties and should not be prioritized over maintaining gradient flow and avoiding saturation.
Adopting a KAN-based generator also allows for training with L1 regularization, followed by pruning with the method proposed by Liu et al. (2024). This enables further model compression beyond the baseline structure defined by KART.
2.3.2 Mixture prior
As noted in Sec. 2.1, the independence of the priors $\pi_{q,p}$ limits expressivity. Moreover, the summation over $p$ in Eq. 8 accumulates the variances of sampled $z_{q,p}$, which can lead to overly diffuse representations. To address this, particularly for more complex generative tasks such as those in Sec. 3.2, we relax the strict adherence to KART by reformulating the prior as a mixture distribution. This involves moving the summation over $p$ out of the generator and absorbing it into the prior, yielding a mixture model:
$\sum_{p=1}^{n_{z}}\pi_{q,p}(z)$ $=\sum_{p=1}^{n_{z}}\alpha_{q,p}\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),\quad\text{where}\quad\sum_{p=1}^{n_{z}}\alpha_{q,p}=1,\quad\alpha_{q,p}\geq 0.$ (12)
Here, $\alpha_{q,p}$ denote mixture weights, practically enforced via a softmax function. The resulting distribution defines a mixture of energy-based priors. Sampling becomes more efficient than in the previous formulation, as it proceeds component-wise using ITS for discrete variables, (Devroye, 2006). However, this violates adherence to KART, since $F$ now corresponds to a discrete rather than continuous CDF.
To encourage balanced weighting and reduce model complexity, L1 regularization can be applied to $\alpha_{q,p}$ and small values can be pruned. The mixture proportions are learned by adding $\log\alpha_{q,p}$ to the gradients of the prior in Sec. 2.4.1. Component selection follows a categorical distribution over $p=1,\dots,n_{z}$, determined by $\alpha_{q,p}$. Let $u_{\text{mix}}\sim\mathcal{U}(u_{\text{mix}};\;0,1)$. Guided by Devroye (2006) regarding how ITS applied to discrete variables, the selected component $p^{*}$ is given by:
$p^{*}=\min\left\{j\mid\sum_{p=1}^{j}\alpha_{q,p}\geq u_{\text{mix}}\right\}.$ (13)
This ensures that each component $p^{*}$ is selected with probability $\alpha_{q,p^{*}}$. ITS can then be used to sample from the chosen component, (see  2.4.3).
2.3.3 Deep KAN prior
If required, we may further deepen the energy functions $f_{q,p}$ using the KAN approach. In fact, each $f_{q,p}$ can itself be a realization of Eq. 1 when its domain is restricted to $[0,1]$. This extension introduces new challenges:
  1. 1.
    Prior interpretability is reduced, complicating the recovery of trained priors.
  2. 2.
    The component-wise structure is compromised, undermining the validity of ITS and KART.
Instead, we propose removing the summation over $p$ in Eq. 8 and reparameterizing the prior using a multilayer KAN energy function, which can be deepened as needed:
$P\left(\bm{z}\mid\bm{f}\right)=\frac{1}{\bm{Z}}\cdot\exp\left(\sum_{q=1}^{2n_{z}+1}\Psi_{q,p}\left(\sum_{p=1}^{n_{z}}f_{q,p}(z)\right)\right)\cdot\pi_{0}(\bm{z}),$ (14)
where $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{n_{z}}$, and is constrained to $0,1]^{n_{z}}$ when strict adherence to KART is required. Sampling from this prior can be carried out using the Unadjusted Langevin Algorithm (ULA), as described in Sec. 2.5.2.
2.4 Maximum likelihood training
With $N_{x}$ training examples in a batch, the optimization objective is the minimization of the negative log-marginal likelihood:
$\min_{\bm{f},\bm{\Phi}}\mathcal{L}(\bm{x}^{(1)},\dots,\bm{x}^{(N_{x})})=\min_{\bm{f},\bm{\Phi}}\left[-\sum_{b=1}^{N_{x}}\log\left(P(\bm{x}^{(b)}\mid\bm{f},\mathbf{\Phi})\right)\right].$ (15)
The conditional dependency on the collection, $\bm{f},\mathbf{\Phi}$, represents all learnable elements of T-KAM. For each sample of $\bm{x}^{(b)}$, the log-marginal likelihood in Eq. 15 can be separated out into two log-distributions by first introducing the latent vector, $\bm{z}$, otherwise absent upon marginalization over the latent distribution. This is shown under Sec. A.4.2, and the resulting expectation exhibits a lack of dependence on sampling, as proven under Sec. A.4.3:
$\nabla_{\bm{f},\bm{\Phi}}\left[\log P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right]$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\Bigg{[}\nabla_{\bm{f}}\left[\log P\left(\bm{z}\mid\bm{f}\right)\right]+\nabla_{\bm{\Phi}}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\Bigg{]},$ (16)
2.4.1 Log-prior
Given the independence of the prior’s parameters, the log-prior can be modeled as a sum over the univariate cases:
$\log\Big{(}P\left(\bm{z}\mid\bm{f}\right)\Big{)}$ $=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(\pi_{q,p}(z_{q,p})\right)=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)$ (17)
Tractable normalization is available through Eq. 6, however it may be more efficient to discard normalization in the learning gradient. In Sec. A.4.4, we derive the contrastive divergence learning gradient for T-KAM, which faster to train and more typical of energy-based modeling:
$\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg{[}\nabla_{\bm{f}}\left[\log P\left(\bm{z}\mid\bm{f}\right)\right]\Bigg{]}=\;$ $\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg{[}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg{]}$
-\; $\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg{[}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg{]}.$ (18)
Here, $\bm{z}$ is sampled from $P(\bm{z}\mid\bm{f})$ and $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$ and reflects a univariate element of the prior model. The expectation with respect to the posterior can be approximated using Importance Sampling (IS), outlined under Secs. 2.4.4 &  A.6.1, or a standard Monte Carlo estimator when using Langevin sampling methods. The prior expectation may always be approximated with a Monte Carlo estimator taken across samples:
$\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg{[}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg{]}\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\quad\text{where}\quad z_{q,p}^{(s)}\sim\pi_{q,p}(z).$ (19)
The gradient of Eq. 18 can therefore be viewed as a element-wise correction to the prior, incorporating information from the observations to align the prior with the posterior, thereby following the philosophy of empirical Bayes. The gradient can then be evaluated via autodifferentiation.
2.4.2 Log-likelihood
Depending on the task the log-likelihood can be modeled in multiple ways. For example, a Gaussian model might be adopted for images, and an autoregressive model for text. It can also be realized into a multimodal generative modeling framework using the principles outlined by Yuan et al. (2024).
$\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)$ $=\begin{cases}-\frac{\left\lVert\bm{x}^{(b)}-\bm{\tilde{x}}\right\rVert^{2}_{2}}{2\sigma_{\text{llhood}}^{2}}+\text{const},&\text{(static continuous data)}\\[5.0pt] \sum_{o=1}^{n_{x}}\sum_{c=1}^{K}x_{o,c}^{(b)}\log\tilde{x}_{o,c}+\text{const},&\text{(static discrete data)}\\[5.0pt] \sum_{t=1}^{T}\log\left(P(\bm{x}^{(b,t)}\mid\bm{x}^{(b,1)},\dots,\bm{x}^{(b,t-1)},\bm{z},\bm{\Phi})\right),&\text{(sequential data)}\end{cases}$ (20)
Here, $\bm{\tilde{x}}$ is a generated sample from Eq. 8 corresponding to a single batch item, $\bm{x}^{(b)}$, $\sigma_{\text{llhood}}$ is a hyperparameter subject to assumptions regarding the variance of the data, and $K$ denotes the number of discrete classes. The learning gradient attributed to the log-likelihood can also be evaluated using autodifferentiation, after first approximating the posterior expectation with IS or MC estimators.
2.4.3 Prior inverse transform sampling
In Eqs. 8 &  19, $z_{q,p}$ is drawn from $\pi_{q,p}$ with inverse transform sampling, (ITS) (Devroye, 1986). ITS can be used to sample from a component of the latent distribution, $\pi_{q,p}$, made possible given the unique characteristics of T-KAM. Specifically, the univariate, finite-grid structure of the constituent functions enables quadrature/interpolating integral approximations.
  1. 1.
    CDF inversion: Sampling occurs using the inverse cumulative distribution function (CDF): The CDF for a single component, $F_{\pi_{q,p}}$, is:
    $F_{\pi_{q,p}}(z^{\prime})$ $=\frac{\int_{z_{\text{min}}}^{z^{\prime}}\exp(f_{q,p}(z))\pi_{0}(z)\,dz}{Z_{q,p}},\quad Z_{q,p}=\int_{z_{\text{min}}}^{z_{\text{max}}}\exp(f_{q,p}(z))\pi_{0}(z)\,dz.$ (21)
    These two integrals can be conveniently and efficiently approximated using Gauss-Kronrod quadrature, (Laurie, 1997), given that $f_{q,p}$ is univariate and defined on a bounded grid:
    $G(z)$ $=\exp(f_{q,p}(z))\pi_{0}(z),\quad\int_{z_{\text{min}}}^{z^{\prime}}G(z)\,dz\approx\sum_{i=1}^{N_{\text{quad}}}w_{i}\,G(z_{i}^{\text{node}}),$ (22)
    where $z_{i}^{\text{node}}$ and $w_{i}$ are the Gauss-Legendre quadrature nodes and weights for the interval $z_{\text{min}},z^{\prime}]$. A random variable is uniformly drawn and shared across $q$, $u_{p}\sim\mathcal{U}(u_{p};\;0,1)$. The first interval where $F_{\pi_{q,p}}(z)$ is greater than $u_{p}$ scaled by $Z_{q,p}$ is identified:
    $k^{*}=\min\Bigg{\{}j\mid\sum_{i=1}^{j}w_{i}\,G(z_{i}^{\text{node}})\geq Z_{q,p}\cdot u_{p}\Bigg{\}}.$ (23)
  2. 2.
    Linear interpolation: The sample $z_{p}$ is interpolated within the subset, $z_{k^{*}-1}^{\text{node}},z_{k^{*}}^{\text{node}}]$:
    $z_{p}$ $\approx z_{k^{*}-1}^{\text{node}}+\left(z_{k^{*}}^{\text{node}}-z_{k^{*}-1}^{\text{node}}\right)\cdot\frac{u_{p}-F_{\pi_{q,p}}(z_{k^{*}-1}^{\text{node}})}{F_{\pi_{q,p}}(z_{k^{*}}^{\text{node}})-F_{\pi_{q,p}}(z_{k^{*}-1}^{\text{node}})}.$ (24)
The procedure ensures an efficient, exact and unbiased generation of samples, $z_{q,p}$, distributed according to $\pi_{q,p}$. It can be vectorized to generate samples from all prior elements.
2.4.4 Posterior Expectations with Importance Sampling
The log-marginal likelihood and its learning gradient presented in Eqs. 15 &  16, involve expectations with respect to the posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is potentially multimodal and complex. We propose Importance Sampling (IS) (Goertzel, 1949) as an efficient method for estimating its expectation when prior task knowledge can be incorporated into the model to improve prior-posterior alignment and the quality of the IS estimator.
IS provides a quick approximation of the posterior expectation without requiring extensive exploration. In contrast, ULA, (introduced in Sec. 2.5.2) is more stochastic and exploratory, often yielding higher-quality samples from the posterior, when the prior is poorly matched.
The following procedure is elaborated in more detail under Sec. A.6.1:
  1. 1.
    Proposal distribution: $N_{s}$ independent samples, $\{\bm{z}^{(s)}\}_{s=1}^{N_{s}}$, are drawn from the proposal distribution, $Q(\bm{z})=P(\bm{z}\mid\bm{f})$, which the target distribution, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$ must be absolutely continuous with respect to. Sampling from $Q$ is tractable with the procedure outlined in Sec. 2.4.3.
  2. 2.
    Importance weights: The normalized importance weights are evaluated:
    $w_{\text{norm}}(\bm{z}^{(s)})=\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big{(}\log(P(\bm{x}^{(b)}\mid\bm{z}^{(s)},\bm{\Phi}))\Big{)}.$ (25)
  3. 3.
    Effective Sample Size: To prevent weight degeneracy, resampling is conducted when the Effective Sample Size (ESS) falls below a threshold specified by a hyperparameter, $\gamma$:
    $ESS=\frac{1}{\sum^{N_{s}}_{s=1}w_{\text{norm}}(\bm{z}^{(s)})}<\gamma\,\cdot\,N_{s}.$ (26)
    Otherwise, the weights are taken without alteration, (i.e. the next step is skipped). When the weights are uniformly spread the ESS evaluates to $N_{s}$. High-variance weights with a low ESS indicate that many samples contribute minimally, resulting in inefficient sampling. Resampling focuses learning on samples that are more representative of the posterior distribution, thereby maximizing the usefulness of each parameter update.
  4. 4.
    Resample: If required, the population, $\{\bm{z}^{(s)},w_{\text{norm}}(\bm{z}^{(s)})\}_{s=1}^{N_{s}}$, is resampled from. There are multiple ways to accomplish this, (Douc et al., 2005). Systematic, stratified, and residual resampling are provided in the codebase under Sec. A.1, parallelized on the CPU. In this study, we adopted residual resampling, (explained under Sec. A.6.2).
  5. 5.
    Expectation estimation: The posterior expectation of an arbitrary function $\rho(\bm{z})$, (e.g., the log-likelihood), is estimated with a Monte Carlo estimator:
    $\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}[\rho(\bm{z})]\approx{\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})\cdot w_{\text{norm}}(\bm{\bar{z}_{\text{resampled}}}^{(s)})}.$ (27)
This estimator is used for computing the log-marginal likelihood and its gradient in Eqs. 15 and 16. However, unlike ULA (discussed in Sec. 2.5.2 and demonstrated in Sec. 3.2), IS fails to support posterior exploration when the prior is poorly aligned, making it unsuitable for general-purpose tasks beyond those similar to the datasets presented in Secs. 3.1.
2.5 Population-based and Langevin Monte Carlo Posterior Sampling
To mitigate the poor convergence of multimodal ULA sampling, we introduce a training criterion that facilitates posterior annealing. Details on the cyclic annealing schedule adopted in this study, along with practical guidance on hyperparameter selection, are provided in Sec. A.3.
2.5.1 Thermodynamic Training Criterion
We inform our derivations using the Thermodynamic Integral. In place of the log-marginal likelihood outlined in Eq. 15, this exploits the following representation derived in Sec. A.4.5:
$\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)\right]\,dt$ (28)
Here, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)$ has been introduced as the power posterior, which can be considered a representation of the original posterior distribution, with a likelihood tempered by a temperature parameter, $t$. This is introduced by Friel & Pettitt (2008) and organized for our purposes as follows:
$P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)$ $=\frac{P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\cdot P(\bm{z}\mid\bm{f})}{\bm{Z}_{t}},$
$\quad\text{where}\quad\bm{Z}_{t}=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]$ $=\int_{\mathcal{Z}}P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\,\,dP(\bm{z}\mid\bm{f})$ (29)
Given the form of Eq. 29, setting $t=0$ results in the power posterior distribution assuming the form of the prior distribution. As $t$ increments, the posterior gradually adopts a shape more closely resembling the true Bayesian posterior measure. At $t=1$, the posterior recovers its original form and converges to the original posterior distribution.
The effect of tempering is illustrated in Fig. 1, starting with a Gaussian prior that is entirely uninformed of $\bm{x}^{(b)}$. As $t$ increases from 0 to 1, the contribution of the likelihood term in Eq. 29 progressively grows, ultimately leading to the posterior at $t=1$. This final posterior can be interpreted as the prior updated by information from the samples of $\bm{x}^{(b)}$.
Figure 1: Demonstration of power posterior evolution with $t$. A simple Gaussian prior is sculpted into forms that resemble a more complex, multi-modal posterior as $t$ is incremented from 0 to 1.
The exact calculation of the integral in Eq. 28 is feasible given that the bounds are $t=0$, corresponding to the expectation being taken with respect to the prior, and $t=1$, where it is taken with respect to the posterior. However, the integral can also be expressed exactly without collapsing the temperatures, offering greater flexibility in how the integral evolves through the temperature schedule, $\{t_{k}\}_{k=0}^{N_{t}}$. In particular, the temperature schedule is discretized as:
$\bm{t}=\left(t_{0},t_{1},\ldots t_{N_{\text{t}}}\right),\quad\quad t_{k}\in[0,1]$ (30)
In Sec. A.4.7, the Steppingstone estimator (SE) introduced by Annis et al. (2019) is derived from the discretized Thermodynamic Integral presented by Calderhead & Girolami (2009), (and provided under Sec. A.4.6). This establishes a connection between the two, which we use to derive the learning gradient and inform our approach to the clustering the temperatures in Eq. 30:
[…]
The log-partition ratios are estimated with Monte Carlo estimators using all annealed samples:
$\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)=\log(\bm{Z}_{t_{N_{t}}})^{\text{SS}}=\sum_{k=1}^{N_{t}}\log\Bigg{(}\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg{)}$
[…]
SE and its variants, such as Annealed Importance Sampling (AIS) (Neal, 1998), are commonly used to estimate a model’s evidence without a tractable way to learn the prior, as they lack dependence on prior gradients beyond sampling. However, by recognizing the direct connection between the SE and MLE learning gradients, as proven in Sec. A.4.7, one can simply add the contrastive divergence learning gradient from Eq. 18 to the SE estimator.
This is valid because the SE gradient matches the MLE gradient, and the likelihood gradient is independent of the prior gradient. The posterior expectation of contrastive divergence in Eq. 18 can be estimated with a Monte Carlo estimator using the final samples of ULA.
2.5.2 Unadjusted Langevin Algorithm (ULA) and Replica Exchange
In their implementation of Thermodynamic Integration, Calderhead & Girolami (2009) estimated the Thermodynamic Integral using the discretization in Sec. A.4.6, using population-based MCMC and Parallel Tempering (PT) (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996) to simultaneously sample from all power posteriors.
Specifically, parallel chains were maintained at each temperature, (or Replica), using Metropolis-Hastings (MH) (Metropolis & Ulam, 1949), to drive local moves for each power posterior. They also proposed a geometric path between the prior and posterior, enabling global swaps between adjacent temperatures. This approach allowed higher temperatures to leverage the efficient mixing of lower temperatures, thereby facilitating a more thorough exploration of the posterior landscape.
We adopt a similar strategy but follow Pang et al. (2020) in using the Unadjusted Langevin Algorithm (ULA) (Rossky et al., 1978b; Brooks et al., 2011) without MH correction. Specifically, we update samples using the forward Euler–Maruyama discretization of the Langevin SDE, as detailed by Roberts & Stramer (2002), with the transition kernel described in Sec. A.6.3:
$\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{(i,\;t_{k})}+\eta\nabla_{\bm{z}^{(i,\;t_{k})}}\log\gamma\left(\bm{z}^{(i,\;t_{k})}\right)+\sqrt{2\eta}\bm{\epsilon}^{(i,\;t_{k})},$ (33)
where $\eta$ is the step size, $\bm{\epsilon}^{(i,\;t_{k})}\sim\mathcal{N}(\bm{0},\bm{I})$, and $\gamma_{t_{k}}$ denotes the target power posterior:
$\log\gamma\left(\bm{z}^{(i,\;t_{k})}\right)=\log P(\bm{z}^{(i,\;t_{k})}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\propto t_{k}\log P(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})},\bm{\Phi})+\log P(\bm{z}^{(i,\;t_{k})}\mid\bm{f}).$
ULA is best suited for smooth, unimodal distributions, but annealing with power posteriors might improve sampling from multimodal distributions. Mixing can be further improved by periodically swapping populations between randomly selected adjacent temperatures, subject to:
$r=\frac{P\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k+1})},\bm{\Phi}\right)^{t_{k}}P\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})},\bm{\Phi}\right)^{t_{k+1}}}{P\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})},\bm{\Phi}\right)^{t_{k}}P\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k+1})},\bm{\Phi}\right)^{t_{k+1}}},$ (34)
where the proposed swap, $\bm{z}^{(i,\;t_{k+1})}\leftrightarrow\bm{z}^{(i,\;t_{k})}$, is accepted with probability: $\min\left(1,r\right)$.
3 Experiments
Our experiments are designed to highlight the unique advantages of T-KAM and to compare the proposed strategies for scaling and training. In Sec. 3.1, we demonstrate the effects of introducing different inductive biases through the choice of reference distributions in Eq. 4, as well as through the KAN basis functions used for $f_{q,p}$ and $\Phi_{o,q}$ with comparisons between Radial Basis Functions (RBFs) (Li, 2024) and Fourier bases (FFTs) (Xu et al., 2024). We show that, when appropriate biases are encoded, importance sampling becomes a viable and highly efficient training strategy. We also demonstrate that the model remains effective when strictly adhering to KART with independent energy-based priors, thereby validating the use of KART as a source of structural inductive bias.
In Sec. 3.2, we compare the mixture EBM prior sampled using ITS against a deep EBM prior sampled using ULA. The latter shares an architectural design similar to that used by Pang et al. (2020), but with Chebyshev KAN (SS et al., 2024) energy functions. We further compare T-KAM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion. These comparisons serve to evaluate the practical strategies proposed for scaling and training T-KAM, as well as to compare ITS and our mixture EBM against ULA and deep EBMs, previously used by Pang et al. (2020).
To quantify the fidelity of generated data in a more reliable manner than previous studies in the literature, we used effectively unbiased metrics based on the principles outlined by Chong & Forsyth (2020). By applying linear regression to extrapolate the values of standard metrics to an infinitely sized sample set, Monte Carlo error was mitigated, (otherwise associated with finite sample sizes).
For Sec. 3.2, we gathered $\overline{\text{FID}}_{\infty}$ and $\overline{\text{KID}}_{\infty}$, which are effectively unbiased estimators of the Fréchet Inception Distance and Kernel Inception Distance, introduced by Heusel et al. (2018); Bińkowski et al. (2021) respectively. The set of sample sizes used for the linear regression were: 1000, 1200, 1400, 1600, 1800, 2000, 2200, 2400, 2600, 2800, 3000.
However, we recognise that quantifying the fidelity of generated data is inherently challenging and not an exact science. While the metrics used in this study have limitations, they align with those commonly found in the literature. Our conclusions are also informed by our subjective perception of the images. Furthermore, no metrics were applied in Secs. 3.1.1 &  3.1.1, as vision metrics rely on the Inception V3 model trained on ImageNet (Szegedy et al., 2015), potentially introducing bias against other datasets.
More image samples are provided in Sec. A.7.
3.1 Inductive biases enable Importance Sampling
3.1.1 MNIST & FMNIST
The MNIST (Deng, 2012) and FMNIST (Xiao et al., 2017) datasets are grayscale datasets that can be modeled with the first likelihood model from Eq. 20. The first collection of benchmarks presented in Sec. A.2 were obtained using the hyperparameters and architecture specified for this experiment in Tab. 4. T-KAM was trained for 10 epochs with 20,000 training examples, an importance sample size of $N_{s}=100$, and a batch size of $N_{x}=100$, (corresponding to 2,000 parameter updates) Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
(a) Uniform
(b) Lognormal
(c) Gaussian
Figure 2: Exponentially-tilted priors, extracted after training on MNIST with RBFs. These plots illustrate how T-KAM can be visualized and interpreted by how its priors are adjusted relative to their reference distributions. The lognormal prior remains untrained due to posterior collapse, while the uniform prior has been skewed towards $z_{1,2}=1$. Gaussian is only slightly tilted.
(a) MNIST
(b) FMNIST
Figure 3: Generated NIST images using RBFs and different exponentially-tilted priors, initialized from $\mathcal{N}(\bm{z};\;\bm{0},\bm{1})$, $\text{Lognormal}(\bm{z};\;\bm{0},\bm{1})$, and $\mathcal{U}(\bm{z};\;\bm{0},\bm{1})$. Lognormal is inverted for clarity.
Lognormal priors resulted in posterior mode collapse. T-KAM learned to generate an average of all MNIST or FMNIST classes, and failed to learn a meaningful latent prior. In contrast, Gaussian priors produced the sharpest images for FMNIST, and uniform yielded the sharpest MNIST digits.
In Fig. 2, we show how T-KAM allows visualization of its univariate latent distributions, which is an interpretability advantage that can help us explain the observed results. Unlike Gaussian and uniform reference distributions, which provide a more symmetric and balanced initial spread over the latent space, lognormal priors introduce a strong right-skew. This asymmetric initialization makes it difficult for T-KAM to learn flexible, separable representations, ultimately impairing its ability to produce binary output features.
In contrast, the symmetric nature of Gaussian and uniform priors allows the sculpting of latent distributions and their skews to support binary generation. When cross-referencing with the priors of other latent features plotted in Sec. A.8, it becomes evident that left-skewed distributions are more successful at MNIST and FMNIST generation. This explains why the standard lognormal prior, which is right-skewed, is poorly suited to these datasets.
The experiment also demonstrates the validity of importance sampling for posterior expectations when prior-posterior alignment can be improved by choosing an appropriate reference prior, since uniform and Gaussian priors produced realistic MNIST and FMNIST samples respectively.
3.1.2 2D Darcy flow pressures
The dataset used in this experiment was sourced from Li et al. (2021). The Darcy flow equation is a second-order elliptic partial differential equation (PDE) that models various physical phenomena, including fluid flow through porous media and heat conduction. At steady state, its two-dimensional form on the unit square with Dirichlet boundary conditions is given by:
$-\nabla\cdot\left(k(\bm{z}),\nabla u(\bm{z})\right)$ $=f(\bm{z}),$ $\bm{z}$ $\in[0,1]^{2},$
$u(\bm{z})$ $=0,$ $\bm{z}$ $\in\partial[0,1]^{2},$ (35)
where $k(\bm{z})$ is the 2D permeability field, $u(\bm{z})$ represents the 2D pressure field, and $f(\bm{z})$ is a forcing function evaluated at $\bm{z}$. Generating new flow-pressure samples is a challenging task in machine learning. One approach is to use input permeability samples while employing the Fourier Neural Operator (FNO) (Li et al., 2021), to learn the solution operator mapping between permeability and flow pressure. However, in this study, we use the purely generative task to illustrate how weak inductive biases can influence what T-KAM learns.
T-KAM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=50,N_{s}=50$. The training usespanned 600 epochs with 1,000 training examples, resulting in 12,000 parameter updates. All other hyperparameters are provided in Tab. 5. Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
(a) $\mathcal{U}(\bm{z};\;\bm{0},\bm{1})$
(b) $\text{Lognormal}(\bm{z};\;\bm{0},\bm{1})$
Figure 4: Generated Darcy flow pressures with exponentially-tilted priors initialized from standard uniform and lognormal distributions. RBF bases, are contrasted against FFT bases, where RBF is colored differently for clarity.
(a) RBF prior on $z_{1,1}$
(b) FFT prior on $z_{1,1}$
Figure 5: Comparison of exponentially-tilted lognormal priors, extracted after training on Darcy between RBF bases and FFT bases. RBF basis functions led to oversmoothing and posterior collapse, while FFT bases lacked smoothness, potentially compromising adherence to the continuous assumptions imposed by KART.
Although RBFs tended to oversmooth the solution space, FFTs were more effective at capturing large-scale spatial periodicity. In cases where the prior was misspecified with a lognormal reference, RBFs failed to produce coherent samples, while FFTs still yielded coherent outputs.
This result likely stems from their suitability in representing periodic spatial relationships, a property that aligns well with the structure of the Darcy flow, as previously showcased by the success of the FNO, which also uses Fourier methods. This is somewhat reflected by Fig. 5(b), which depicts some periodicity in the learned, exponentially-tilted distribution on $z_{1,1}$. However, FFT-parameterized energy functions were also found to be non-smooth and potentially discontinuous, which is also illustrated by the distribution plotted in Fig. 5(b). This might have compromised adherence to KART and future work might explore incorporating regularization to improve their smoothness.
This shows that the choice of basis functions serves as an additional mechanism for encoding inductive bias, complementing the role of the reference prior. Once again, importance sampling proved to be viable when appropriate inductive biases were encoded.
3.2 Comparing scaling strategies against CIFAR-10 & SVHN
T-KAM, when adherent to KART, can be interpreted as a dense or spline-based network. However, for spatial data, convolutional neural networks (CNNs) (LeCun et al., 2015), are more suitable as the generator, $\mathcal{G}$. In addition to their computational efficiency through weight sharing, CNNs offer a key advantage over dense networks due to their spatial inductive bias: sequences of convolutional layers implement shift-equivariant transformations that are well-suited for images.
Due to limited computational resources during this study, (detailed in Sec. A.2), training was shortened from 70,000 to 8,000 parameter updates. Given their parameter efficiency, Chebyshev bases (SS et al., 2024) were used for the prior energy functions. The deep prior architecture mirrored that of Pang et al. (2020), but it is important to note that dense KANs do not directly replicate the behavior of standard dense layers. Additionally, we used a single optimizer to train both models, rather than using separate optimizers for the prior and generator networks. We considered these discrepancies to be acceptable, given the comparative nature of this experiment.
Posterior sampling was conducted with ULA, and a Gaussian reference was used, $\pi_{0}=\mathcal{N}(\bm{z};\;\bm{0},\bm{1})$. T-KAM was trained for 20 epochs with a batch size of $N_{x}=100$ on 40,000 training examples. All remaining hyperparameters are provided in Tab. 6, and architectural details for the mixture and deep prior models, as well as the CNN generators, are included in Sec. A.5.1. Annealing was conducted with $N_{t}=5$ power posteriors and the cyclic temperature scheduling outlined in Sec. A.3.1.
For brevity, only the SVHN images are visualized in the main body. Samples of CIFAR-10 are provided under Sec. A.7.5
Table 1: Effectively unbiased FID and KID scores (lower is better) for SVHN.
Prior MLE SE
Mixture $\overline{\text{FID}}_{\infty}=81.02$ $\overline{\text{FID}}_{\infty}=82.13$
$\overline{\text{KID}}_{\infty}=0.0669$ $\overline{\text{KID}}_{\infty}=0.0673$
Deep $\overline{\text{FID}}_{\infty}=107.42$ $\overline{\text{FID}}_{\infty}=141.24$
$\overline{\text{KID}}_{\infty}=0.0970$ $\overline{\text{KID}}_{\infty}=0.1320$
Table 2: Effectively unbiased FID and KID scores (lower is better) for CIFAR-10.
Prior MLE SE
Mixture $\overline{\text{FID}}_{\infty}=159.11$ $\overline{\text{FID}}_{\infty}=180.00$
$\overline{\text{KID}}_{\infty}=0.1600$ $\overline{\text{KID}}_{\infty}=0.1858$
Deep $\overline{\text{FID}}_{\infty}=185.10$ $\overline{\text{FID}}_{\infty}=201.75$
$\overline{\text{KID}}_{\infty}=0.1780$ $\overline{\text{KID}}_{\infty}=0.1977$
(a) SVHN with mixture prior and MLE
(b) SVHN with mixture prior and SE
Figure 6: Generated samples from T-KAM trained on SVHN using a mixture prior.
(c) SVHN with deep prior and MLE
(d) SVHN with deep prior and SE
Figure 7: Generated samples from T-KAM trained on SVHN using a deep prior.
The mixture prior with ITS outperformed the deep prior with ULA, while also being more computationally efficient and enabling faster inference, (Tab. 9). This advantage is likely due to ITS providing exact samples from the prior, whereas ULA produces only approximate, biased samples.
SE did not outperform MLE and cannot be justified in this setting given its computational overhead. Alternative approaches, such as different annealing schedules or the use of autoMALA with bias correction and temperature-adaptive step size tuning, could improve performance, but they are not worth investigating until the hardware of Zetta (2024) becomes available, due to the added complexity, hyperparameter tuning, and extended training time.
The results confirm the superiority of mixture EBMs paired with ITS over deep EBMs sampled with ULA as a viable approach to scaling. However, the advantages of SE and population-based ULA remain inconclusive.
4 Conclusions
T-KAM provides a principled and scalable framework for structured and informed generative modeling. Our results demonstrate that incorporating inductive biases through both basis functions and reference priors can enable more efficient training strategies, such as importance sampling. We also show how the Kolmogorov–Arnold representation theorem can inform architectural design and introduce structural bias. Additionally, we find that the inverse transform method offers clear advantages over Langevin Monte Carlo sampling for energy-based priors. While mixture priors with inverse transform sampling consistently outperformed alternative approaches across both datasets, the benefits of annealing and population-based MCMC remain inconclusive, suggesting directions for further work. Future investigations into the practicality of our annealing methods depend on continued hardware advancements, such as those introduced by Zetta (2024).
4.1 Future work
4.1.1 Hardware acceleration
Computational benchmarks are provided in Sec. A.2. The univariate and localized nature of spline interpolations complicates their parallelization on GPUs, since they involve sparse or recursive computations. Additionally, the search operations in Eqs. 1323, and 70 were parallelized on the CPU across samples, despite the remainder of T-KAM being executed on the GPU. This design stems from the fact that GPUs are generally ill-suited for parallelized search algorithms, which often rely on conditional logic. Parallelizing conditionals can lead to thread divergence, making CPUs more efficient for these operations.
While ITS enables fast inference and scales when the prior is a mixture model, T-KAM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA. Nonetheless, the unique advantages of T-KAM justify continued research and development. Notably, Zetta (2024) introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control and high throughput.
The LFU presents a promising avenue for more efficient implementations of ULA, and improved posterior sampling, potentially including bias correction and adaptive step size tuning with autoMALA, (Sec. A.6.4). Even without annealing, replacing ULA with autoMALA may yield a more robust probabilistic model with stronger theoretical guarantees.
Overall, T-KAM may prove more scalable than conventional generative models, owing to its reliance on finite representations of continuous functions and its compatibility with the LFU.
4.1.2 Text-to-text, classification, and multi-modal generation
Kong et al. (2021) provided a robust framework for high-quality text-to-text generation, classification, and meta-learning, while Yuan et al. (2024) demonstrated how T-KAM might be adapted into a multi-modal generative model. This holds promise for future investigations and industry adoption, especially when the LFU becomes available. However, the symbol-vector coupling approach of Kong et al. (2021) must first be adapted to remain compatible with ITS.
4.1.3 Improved sampling for multimodal distributions
Despite its theoretical guarantees, annealing is difficult to justify given its computational cost and limited practical benefit. Future work could explore adaptations of the hierarchical EBMs proposed by Cui & Han (2024), without sacrificing the interpretability or efficiency of our model.
Appendix A Appendix
A.1 Codebase
The study was implemented at https://github.com/exa-laboratories/T-KAM.
A.2 Computational benchmarks
These benchmarks were collected for FP32 on a GPU (NVIDIA GeForce GTX 1650) with Julia’s BenchmarkTools package, (JuliaCI, 2024). The index search operations in Eqs. 2313, & 70 were parallelized across 12 threads on a CPU (i7-9750H @ 2.60GHz).
Figure 8: Benchmarks for varying latent dimensions, $n_{z}$, while sampling from independent EBM priors with ITS and estimating posterior expectations with IS. The prior had an input size of $n_{z}$ and an output size of $2n_{z}+1$. The generator received inputs of size $2n_{z}+1$, with a hidden size of $4n_{z}+2$ and an output size of $28\times 28$. All other hyperparameters were kept consistent with Tab. 4. They reflect the performance of a reverse autodifferentiation of the MLE criterion, Eq. 16.
Figure 9: Benchmarks for varying amounts of ULA iterations, $N_{prior}$, to sample from a deep EBM prior (Tab. 8). Posterior expectations were estimated with an Monte Carlo estimator of the MLE criterion, Eq. 16. All other hyperparameters were kept consistent with Tab. 6. They reflect the computational cost of generating samples from the prior with Eq. 33. The benchmarks for ULA appear relatively constant across different step counts due to Julia’s loop optimizations and efficient memory management. Nevertheless, the reference ITS sampler remains significantly faster in practice. Note: Since the reference ITS sampling executed very rapidly, its relatively high proportion of time attributed to garbage collection is not unexpected.
Figure 10: Benchmarks for varying amounts of temperatures, $N_{t}$, while using a mixture of EBM priors (Tab. 7) with ITS sampling. Posterior expectations were estimated with SE and population-based ULA. All other hyperparameters were kept consistent with Tab. 6. They reflect the computational cost of performing a single reverse autodifferentiation of the SE criterion, Eq. 31. Note: Since the reference MLE criterion is executed very rapidly, its relatively high proportion of time attributed to garbage collection is not unexpected.
A.3 Guidance on tempering
A.3.1 Cyclic annealing
Calderhead & Girolami (2009) demonstrated that uniformly scheduling the discretized temperatures in Eq. 30 results in unequal contributions of adjacent power posteriors to the Thermodynamic Integral. The success of TI and SE, relies on selecting each $t_{k}$ to ensure sufficient overlap of adjacent power posteriors. This overlap ensures similarity between adjacent power posteriors, as introduced in Sec. A.6.1, which influences the Monte Carlo estimator in Eq. 32, given that SE is an application of importance sampling, (Annis et al., 2019).
This overlap is quantified by the KL divergence term in Eq. 52. Maintaining small KL divergences between adjacent power posteriors can be achieved by clustering temperatures in regions where $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$ changes rapidly. To accomplish this, we follow Calderhead & Girolami (2009) in adopting a power-law relationship to schedule the temperatures of Eq. 30:
$t_{k}=\left(\frac{k}{N_{t}}\right)^{p},\quad k=0,1,\ldots,N_{t},$ (36)
The effect of varying $p$ on the schedule between 0 and 1 is visualized in Fig. 11. When $p>1$, spacing becomes denser near $t=0$, while $p<1$ creates denser spacing near $t=1$. The impact of varying $p$ on the discretized Thermodynamic Integral of Eq. 52 is illustrated in Fig. 12. The plots demonstrate that for $p>1$, the trapezoidal bins are clustered closer to $t=0$, concentrating the discrete approximation error towards $t=1$.
Figure 11: Power-law schedule.
(a) $p=0.35$ schedule has more bins clustered towards $t=1$.
(b) $p=1$ schedule is uniformly distributed between the bounds.
(c) $p=4$ schedule has more bins clustered towards $t=0$.
Figure 12: Effect of power-law schedule clusters on integral approximation error. Evaluation points can be skewed towards a particular bound by choosing pp. Increasing $p$ provides more tempering, concentrating more evaluation points on smoother likelihoods. Tempering helps to mitigate practical challenges associated with the non-smoothness of the univariate functions in Eq. 1.
The study by Calderhead & Girolami (2009) derives an analytic term for an optimal tempering distribution in the context of linear regression models, where the power posterior in Eq. 29 is used to evaluate the plausibility of model parameters in relation to the data, rather than reflecting a latent variable posterior. However, their findings may not hold in the context of T-KAM and generative modeling, making $p$ and the temperature schedule difficult variables to tune.
Instead, we schedule $p$ cyclically across all parameter updates of T-KAM’s training procedure, similar to the previous work of Fu et al. (2019). The initial and final $p$’s and the number of cycles are treated as tunable hyperparameters:
$\delta$ $=\text{linspace}\left(0,\ 2\pi(N_{\text{cycles}}+0.5);\ \text{len}=N_{\text{updates}}\right)$
$p(i)$ $=p_{\text{init}}+\frac{p_{\text{end}}-p_{\text{init}}}{2}\left(1-\cos\delta\right)$ (37)
where $N_{\text{updates}}$ denotes the total number of parameter updates, $i$ indexes the current training iteration, $N_{\text{cycles}}$ is the number of cycles, and $\delta$ is a vector of length $N_{\text{updates}}$.
Initializing $p>1$ focuses learning on smoother likelihoods during the early stages of the cycle, where global features emerge between adjacent power posteriors. As training progresses, T-KAM will have learned the global features sufficiently well, reducing the change of measure between adjacent power posteriors of a $p>1$ schedule. At this stage, $p(i)$ will have transitioned to $p\leq 1$, thereby shifting T-KAM’s focus towards capturing more detailed features of the posterior landscape, where training is expected to enhance its ability to generate finer details.
Guidance on selecting optimal values for the tempering hyperparameters is provided in Tab. 3.
A.3.2 Choosing hyperparameters
Table 3: Tempering hyperparameters.
Hyperparameter Description
${N_{t}}>1$ Specifies the number of power posteriors, enabling a smooth transition between the prior and posterior. Increasing this improves exploration but also increases cost due to the need for more MALA iterations. Typically, this is limited and therefore guided by hardware availability, such as $N_{t}=5$.
$N_{\text{local}}>1$ Determines the number of ULA or autoMALA iterations per power posterior. Although increasing this value improves the quality of samples at each stage, it also raises the computational cost, much like $N_{t}$. It can be kept at a moderate value, such as $N_{\text{local}}=20$, (based on the recommendations of Nijkamp et al. (2019b; a) regarding short-run Markov chain Monte Carlo).
$N_{\text{unadjusted}}>1$ Controls the number of autoMALA iterations performed without MH adjustments, reflecting the level of exploration at each power posterior. This can be set to 1, as guided by Biron-Lattes et al. (2024).
$p_{\text{init}}>1,$ $p_{\text{end}}\leq 1,$ $N_{\text{cycles}}$ Controls the evolution of tempering during training. If the data is expected to have smooth relationships, reduce both $p_{\text{init}}$ and $p_{\text{end}}$. For data with finer details, increase both values to allow for better focus on capturing intricate features. Recommended values are $p_{\text{init}}=4,$ and $p_{\text{end}}=1$, since our experience suggests that power posteriors change rapidly closer to the prior.
A.4 Derivations and other identities
A.4.1 Useful proof
The following short proof for a learning gradient with respect to an arbitrary set of parameters, $\bm{\theta}$, proves useful for other derivations in this study:
$\mathbb{E}_{P(\bm{x}\mid\bm{\theta})}\left[\nabla_{\bm{\theta}}\log P(\bm{x}\mid\bm{\theta})\right]$ $=\mathbb{E}_{P(\bm{x}\mid\bm{\theta})}\left[\frac{\nabla_{\bm{\theta}}P(\bm{x}\mid\bm{\theta})}{P(\bm{x}\mid\bm{\theta})}\right]=\int_{\mathcal{X}}\frac{\nabla_{\bm{\theta}}P(\bm{x}\mid\bm{\theta})}{P(\bm{x}\mid\bm{\theta})}\,\,\,dP(\bm{x}\mid\bm{\theta})$
$=\int_{\mathcal{X}}\nabla_{\bm{\theta}}P(\bm{x}\mid\bm{\theta})\,d\bm{x}=\nabla_{\bm{\theta}}\int_{\mathcal{X}}P(\bm{x}\mid\bm{\theta})\,d\bm{x}=\nabla_{\bm{\theta}}1=0.$ (38)
A.4.2 Marginal likelihood
$\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\left[\log\left(P(\bm{x}^{(b)},\bm{z}\mid\bm{f},\bm{\Phi})\right)\right]$
$=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\left[\log\left(P\left(\bm{z}\mid\bm{f}\right)\cdot P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)\right]$
$=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\left[\log\left(P\left(\bm{z}\mid\bm{f}\right)\right)+\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)\right].$ (39)
A.4.3 Maximum likelihood learning gradient
This proof demonstrates that the learning gradient with respect to an arbitrary set of parameters, $\theta$, is independent of sampling:
$\mathbb{E}_{P(\bm{z}\mid\bm{x})}\left[\nabla_{\bm{\theta}}\log P(\bm{x},\bm{z})\right]$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x})}\left[\nabla_{\bm{\theta}}\log P(\bm{z}\mid\bm{x})+\nabla_{\bm{\theta}}\log P(\bm{x})\right]$
$=\mathbb{E}_{P(\bm{z}\mid\bm{x})}\left[\nabla_{\bm{\theta}}\log P(\bm{z}\mid\bm{x})\right]+\mathbb{E}_{P(\bm{z}\mid\bm{x})}\left[\nabla_{\bm{\theta}}\log P(\bm{x})\right]$
$=0+\nabla_{\bm{\theta}}\log P(\bm{x}),$ (40)
where Eq. 38 has been used to zero the expected posterior. This allows gradient flow to be disregarded in sampling procedures as shown in Eq. 16.
A.4.4 Contrastive divergence learning gradient for log-prior
The learning gradient attributed to the prior can be derived into a contrastive divergence framework, which is more typical of energy-based modeling literature. The prior model is solely dependent on $\bm{f}$ in Eq. 4. We can rewrite Eq. 17 as:
$\nabla_{\bm{f}}\log\left(P\left(\bm{z}\mid\bm{f}\right)\right)$ $=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(\pi_{q,p}(z_{q,p})\right)=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)$ (41)
where $Z_{q,p}=\int_{\mathcal{Z}}\exp(f_{q,p}(z_{p}))\,d\pi_{0}(z_{p})$ is the normalization constant. We can simplify the constant using Eq. 38:
$\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg{[}\nabla_{\bm{f}}\log P\left(\bm{z}\mid\bm{f}\right)\Bigg{]}$ $=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg{[}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)\Bigg{]}=0$
$=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg{[}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg{]}-\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(Z_{q,p}\right)=0$
$\therefore\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(Z_{q,p}\right)$ $=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg{[}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg{]}$ (42)
Therefore the log-prior gradient of Eq. 41 can be evaluated without quadrature normalization as:
$\nabla_{\bm{f}}\log P\left(\bm{z}\mid\bm{f}\right)$ $=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})-\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg{[}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg{]},$ (43)
and the learning gradient attributed to the prior is finalized into a contrastive divergence format:
$\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg{[}\nabla_{\bm{f}}\left[\log P\left(\bm{z}\mid\bm{f}\right)\right]\Bigg{]}=\;$ $\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg{[}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg{]}$
-\; $\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg{[}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg{]}.$ (44)
A.4.5 Deriving the Thermodynamic Integral
Starting from the power posterior:
$P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)=\frac{P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\,\,P(\bm{z}\mid\bm{f})}{\bm{Z}_{t}}$ (45)
The partition function $\bm{Z}_{t}$ is given by:
$\bm{Z}_{t}=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]$ (46)
Now, differentiating the log-partition function with respect to $t$:
$\frac{\partial}{\partial t}\log(\bm{Z}_{t})$ $=\frac{1}{\bm{Z}_{t}}\frac{\partial}{\partial t}\bm{Z}_{t}=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[\frac{\partial}{\partial t}P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]$
$=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\cdot P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]\$
$=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]$ (47)
Integrating both sides from 0 to 1:
$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\,dt$ (48)
Using the fundamental theorem of calculus, (Larson & Edwards, 2008):
$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\,dt$ (49)
The limits of the integrals can be expressed as:
$\log(\bm{Z}_{t=0})$ $=\log\Big{(}\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{0}\right]\Big{)}=\log(1)=0$
$\log(\bm{Z}_{t=1})$ $=\log\Big{(}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\Big{)}=\log P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})$ (50)
Therefore, we obtain the log-marginal likelihood expressed as the Thermodynamic Integral:
$\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\log P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\,dt$ (51)
A.4.6 Discretizing the Thermodynamic Integral
Here, $t_{k}$ denotes the tempering at the $k$-th index of the schedule, and $N_{\text{t}}$ is the number of temperatures. As derived by Calderhead & Girolami (2009), Eq. 28 then be evaluated using:
$\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(P_{t_{k-1}}||P_{t_{k}})-D_{\text{KL}}(P_{t_{k}}||P_{t_{k-1}})$ (52)
Where:
$\Delta t_{k}$ $=t_{k}-t_{k-1},\quad\quad E_{k}=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\left[\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)\right],$
$D_{\text{KL}}(P_{t_{k-1}}||P_{t_{k}})$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}\Bigg{[}\log\Bigg{(}\frac{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Bigg{)}\Bigg{]}$ (53)
This reflects the trapezium rule for numerical integration, supported by bias correction to provide an estimate of the Riemann integral that is especially robust to error. The learning gradient derived from the discretized Thermodynamic Integral is unchanged from the MLE gradient in Eq. 16, as demonstrated under Sec. A.4.7.
A.4.7 Learning gradient for discretized Thermodynamic Integral and power posterior partition function
We can derive the learning gradient of the discretized Thermodynamic Integral in Eq. 52 by reducing it down into a sum of normalizing constants. Starting from the definitions attributed to the KL divergence terms:
$D_{\text{KL}}(P_{t_{k-1}}\|P_{t_{k}})$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}\Big{[}\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\Big{]}$
$\text{and}\quad D_{\text{KL}}(P_{t_{k}}\|P_{t_{k-1}})$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big{[}\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})\Big{]}$
$=-\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big{[}\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\Big{]}$ (54)
We can leverage Eq. 29 to quantify the log-difference in terms of the change in likelihood contribution plus some normalization constant:
$\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})$ $-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})=-\Delta t_{k}\cdot\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})+\log\Bigg{(}\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg{)}$ (55)
The definitions in Eq. 53 can be used to express the KL divergences in terms of the expected likelihoods:
$\therefore D_{\text{KL}}(P_{t_{k-1}}\|P_{t_{k}})$ $=-\Delta t_{k}\cdot\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}\Big{[}\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\Big{]}+\log\Bigg{(}\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg{)}$
$=-\Delta t_{k}\cdot E_{k-1}+\log\Bigg{(}\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg{)}$ (56)
$\text{and}\quad D_{\text{KL}}(P_{t_{k}}\|P_{t_{k-1}})$ $=\Delta t_{k}\cdot\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big{[}\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\Big{]}-\log\Bigg{(}\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg{)}$
$=\Delta t_{k}\cdot E_{k}-\log\Bigg{(}\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg{)}$ (57)
Substituting back into Eq. 52, we eliminate the expected likelihood terms:
$\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(P_{t_{k-1}}\|P_{t_{k}})$ $-D_{\text{KL}}(P_{t_{k}}\|P_{t_{k-1}})=-\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg{(}\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg{)}$
$\therefore\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)$ $=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(P_{t_{k-1}}||P_{t_{k}})-D_{\text{KL}}(P_{t_{k}}||P_{t_{k-1}})$
$=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-\Bigg{[}\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg{(}\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg{)}\Bigg{]}$
$=\sum_{k=1}^{N_{t}}\log\Bigg{(}\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg{)}=\sum_{k=1}^{N_{t}}\log(\bm{Z}_{t_{k}})-\log(\bm{Z}_{t_{k-1}})$ (58)
Following a similar method to Sec. A.4.4 we can use Eq. 38 to express the learning gradient attributed to each normalization constant in a simpler manner:
$\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big{[}\nabla_{\bm{f},\bm{\Phi}}\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\Big{]}$
$=\,$ $\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big{[}t_{k}\cdot\nabla_{\bm{\Phi}}\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})+\nabla_{\bm{f}}\log P(\bm{z}\mid\bm{f})\Big{]}-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=0$
$\therefore$ $\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big{[}t_{k}\cdot\nabla_{\bm{\Phi}}\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})+\nabla_{\bm{f}}\log P(\bm{z}\mid\bm{f})\Big{]}$ (59)
Telescoping the sum in Eq. 58, and noting that $k=1$ corresponds to the power posterior equaling the prior, and $k=N_{t}$ to the posterior, Eq. 58 yields T-KAM’s learning gradient:
$\nabla_{\bm{f},\bm{\Phi}}\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)$ $=\sum_{k=1}^{N_{t}}\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k-1}})$
$=\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{N_{t}}})-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{0}})$
$=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\Bigg{[}\nabla_{\bm{f}}\log P\left(\bm{z}\mid\bm{f}\right)+\nabla_{\bm{\Phi}}\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\Bigg{]}$
[…]
A.5 Hyperparameters and training details
Table 4: Hyperparameters for Sec. 3.1.1.
EBM Prior
layer_widths $p,q$ 40, 81
spline_type (-) variable
spline_activation (-) Silu
layer_norm (-) true
quadrature_method (-) gausslegendre
spline_degree $d$ 3
grid_size $N_{\text{grid}}$ RBF: 20
FFT: 50
grid_update_ratio $\alpha_{\text{grid}}$ 0.05
grid_range $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
Lognormal: $[0,4]$
Gaussian: [1.2,1.2][-1.2,1.2]
Spline basis $\varepsilon$ init. $\varepsilon$ 0.1
Spline basis $\mu$ scaling $\mu$ 1.0
Basis activation scaling $\sigma_{b}$ 1.0
Basis spline scaling $\sigma_{s}$ 1.0
Initial basis param. $\tau_{0}$ 1.0
Train basis param. $\tau_{\text{trainable}}$ true
GaussQuad_nodes $N_{\text{quad}}$ 100
Use mixture prior (-) false
KAN Likelihood
hidden_widths $q,2q+1$ 81, 162
All other KAN hparams Identical to EBM Prior
generator_variance $\sigma_{\epsilon}$ 1.0
output_activation $\sigma_{\text{act}}$ sigmoid
Importance sampling resampler resampler residual
ESS threshold $\gamma$ 0.5
Grid updating
update_prior_grid (-) true
update_llhood_grid (-) true
num_grid_updating_samples (-) 50
grid_update_frequency (-) 100
grid_update_decay (-) 0.999
Training
batch_size $N_{x}$ 100
importance_sample_size $N_{s}$ 100
Training examples $N_{\text{train}}$ 20,000
Testing examples $N_{\text{test}}$ 100
num_generated_samples (-) 3,000
Number of epochs $N_{\text{epochs}}$ 10
eps $\epsilon$ 0.001
Optimizer
type (-) adam
learning_rate $\eta_{LR}$ 0.001
betas (-) (0.9, 0.999)
Table 5: Hyperparameters for Sec. 3.1.2.
EBM Prior
layer_widths $p,q$ 40, 81
spline_type (-) variable
spline_activation (-) Silu
layer_norm (-) true
quadrature_method (-) gausslegendre
spline_degree $d$ 3
grid_size $N_{\text{grid}}$ RBF: 20
FFT: 50
grid_update_ratio $\alpha_{\text{grid}}$ 0.05
grid_range $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
Lognormal: $[0,4]$
Gaussian: [1.2,1.2][-1.2,1.2]
Spline basis $\varepsilon$ init. $\varepsilon$ 0.1
Spline basis $\mu$ scaling $\mu$ 1.0
Basis activation scaling $\sigma_{b}$ 1.0
Basis spline scaling $\sigma_{s}$ 1.0
Initial basis param. $\tau_{0}$ 1.0
Train basis param. $\tau_{\text{trainable}}$ true
GaussQuad_nodes $N_{\text{quad}}$ 100
Use mixture prior (-) false
KAN Likelihood
hidden_widths $q,2q+1$ 81, 162
All other KAN hparams Identical to EBM Prior
generator_variance $\sigma_{\epsilon}$ 0.1
output_activation $\sigma_{\text{act}}$ sigmoid
Importance sampling resampler resampler residual
ESS threshold $\gamma$ 0.5
Grid updating
update_prior_grid (-) true
update_llhood_grid (-) true
num_grid_updating_samples (-) 100
grid_update_frequency (-) 100
grid_update_decay (-) 0.999
Training
batch_size $N_{x}$ 50
importance_sample_size $N_{s}$ 50
Training examples $N_{\text{train}}$ 1,000
Testing examples $N_{\text{test}}$ 1,000
num_generated_samples (-) 3,000
Number of epochs $N_{\text{epochs}}$ 600
eps $\epsilon$ 0.001
Optimizer
type (-) adam
learning_rate $\eta_{LR}$ 0.001
betas (-) (0.9, 0.999)
Table 6: Hyperparameters for Sec. 3.2.
EBM Prior
layer_widths $p,q$ See Sec. A.5.1
spline_type (-) Cheby
spline_activation (-) Silu
layer_norm (-) false
quadrature_method (-) gausslegendre
spline_degree $d$ 3
grid_size $N_{\text{grid}}$ 0, (Chebyshev requires no grid)
grid_update_ratio $\alpha_{\text{grid}}$ N/A
grid_range $z_{\text{min}},z_{\text{max}}] [$1.2,1.2][-1.2,1.2]
Spline basis $\varepsilon$ init. $\varepsilon$ 0.1
Spline basis $\mu$ scaling $\mu$ 1.0
Basis activation scaling $\sigma_{b}$ 1.0
Basis spline scaling $\sigma_{s}$ 1.0
Initial basis param. $\tau_{0}$ 1.01
Train basis param. $\tau_{\text{trainable}}$ false
GaussQuad_nodes $N_{\text{quad}}$ 100
Use mixture prior (-) variable
KAN Likelihood
cnn_widths (-) See Sec. A.5.1
All other KAN hparams N/A
generator_variance $\sigma_{\epsilon}$ 0.3
output_activation $\sigma_{\text{act}}$ sigmoid
Importance sampling resampler resampler Not used
ESS threshold $\gamma$ N/A
Grid updating
update_prior_grid (-) false
update_llhood_grid (-) false
Training
batch_size $N_{x}$ 100
importance_sample_size $N_{s}$ N/A
Training examples $N_{\text{train}}$ 40,000
Testing examples $N_{\text{test}}$ 100
num_generated_samples (-) 3,000
Number of epochs $N_{\text{epochs}}$ 20
eps $\epsilon$ 0.0001
Optimizer
type (-) adam
learning_rate $\eta_{LR}$ 0.0005
betas (-) (0.9, 0.999)
PRIOR_LANGEVIN
iters $N_{\text{local}}$ 60
step_size $\eta$ 0.16
POSTERIOR_LANGEVIN
iters $N_{\text{local}}$ SVHN: 20
CIFAR10: 40
step_size $\eta$ 0.01
THERMODYNAMIC_INTEGRATION
iters $N_{\text{local}}$ 20
num_temps $N_{t}$ 5
Initial exponent in power law schedule $p_{\text{start}}$ 4
Final exponent in power law schedule $p_{\text{end}}$ 1
num_cycles $N_{\text{cycles}}$ 0
Global swap frequency, replica_exchange_frequency (-) 1
A.5.1 SVHN & CIFAR-10 Architectures
Table 7: Dense Chebyshev KAN architecture for mixture prior model
Layer Operation Output Dim.
Input Latent vector $\bm{z}$ $p=100$
100 Dense($100$) $q=201$
sizes=[100,201]
Table 8: Dense Chebyshev KAN architecture for deep prior model
Layer Operation Output Dim.
Input Latent vector $\bm{z}$ $p=100$
100 Dense($100$) $200$
200 Dense($200$) $200$
200 Dense($200$) $q=1$
sizes=[100,200,200,1]
Table 9: CNN Generator Architecture for SVHN
Layer Operation Output Shape Stride
Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
1 $\times4$ convT($512$), LeakyReLU $4\times 4\times 512$ 1
2 $\times4$ convT($256$), LeakyReLU $8\times 8\times 256$ 2
3 $\times4$ convT($128$), LeakyReLU $16\times 16\times 128$ 2
4 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
Config: kernel=4, strides=1/2/2/2, paddings=0/1/1/1, activation=LeakyReLU, batchnorm=False
Table 10: CNN Generator Architecture for CIFAR-10
Layer Operation Output Shape Stride
Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
1 $\times8$ convT($1024$), LeakyReLU $8\times 8\times 1024$ 1
2 $\times4$ convT($512$), LeakyReLU $16\times 16\times 512$ 2
3 $\times4$ convT($256$), LeakyReLU $32\times 32\times 256$ 2
4 $\times3$ convT(3), Sigmoid $32\times 32\times 3$ 1
Config: kernel=8/4/4/3, strides=1/2/2/1, paddings=0/1/1/1, activation=LeakyReLU, batchnorm=False
A.6 Sampling theory
This section provides more detail regarding T-KAM’s sampling procedures.
A.6.1 Importance Sampling
Importance Sampling (IS) is a Monte Carlo method rooted in the measure-theoretic formulation of probability. It enables the estimation of expectations with respect to one probability measure using samples drawn from another. For a detailed exposition, see Kloek & van Dijk (1978).
The Radon-Nikodym derivative
IS involves rewriting the expectation of an arbitrary function, $\rho(\bm{z})$, under the target measure, $P(\bm{z})$, in terms of a more tractable proposal measure, $Q(\bm{z})$. Specifically, we use the Radon-Nikodym derivative $\frac{dP}{dQ}$ to rewrite the expectation of an arbitrary function of $\bm{z}$, denoted as $\rho(\bm{z})$:
$\mathbb{E}_{P(\bm{z})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\,dP(\bm{z})=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\cdot\frac{dP}{dQ}\,dQ(\bm{z}).$ (61)
Here, $\frac{dP}{dQ}$ acts as a weight that adjusts the contribution of each sample to account for the difference between the target and proposal measures. This requires that the proposal measure $Q$ is absolutely continuous with respect to the target measure $P$, i.e. $Q(A)=0$ implies $P(A)=0$, for every measurable subset $A\in\mathcal{Z}$.
In the context of T-KAM, the target distribution is the posterior, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is proportional to the product of the likelihood and the prior:
$P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})\propto P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\cdot P(\bm{z}\mid\bm{f}).$ (62)
The exact normalization constant, (the marginal likelihood $P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})$), is intractable, but Importance Sampling circumvents the need to compute it explicitly. Instead, the expectation of $\rho(\bm{z})$ under the posterior can be expressed as:
$\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})w(\bm{z})\,dQ(\bm{z}),$ (63)
where the importance weights $w(\bm{z})$ are given by:
$w(\bm{z})=\frac{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}{Q(\bm{z})}=\frac{P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\cdot P(\bm{z}\mid\bm{f})}{Q(\bm{z})}.$ (64)
Proposal distribution and importance weights
A practical choice for the proposal distribution $Q(\bm{z})$ is the prior, $P(\bm{z}\mid\bm{f})$, given the availability of draws from the prior using the method outlined in Sec. 2.4.3. This simplifies the importance weights, as the prior has a tractable form and covers the support of the posterior. This yields:
$w(\bm{z}^{(s)})=P(\bm{x}^{(b)}\mid\bm{z}^{(s)},\bm{\Phi})\propto\exp(\log(P(\bm{x}^{(b)}\mid\bm{z}^{(s)},\bm{\Phi}))).$ (65)
This reflects that the likelihood now directly informs how much weight to assign to each sample $\bm{z}$. Intuitively, samples that better explain the observed data $\bm{x}^{(b)}$, (as measured by the likelihood), are given higher importance. Using the definition of softmax, the normalized weights are given by:
$\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big{(}\log(P(\bm{x}^{(b)}\mid\bm{z}^{(s)},\bm{\Phi}))\Big{)}.$ (66)
Resampling
The suitability of these weights depends on how well $Q(\bm{z})$ matches the posterior. In T-KAM, there is likely to be significant mismatch due to the complexity of the likelihood models described in Eq. 20. Consequently, the weights will exhibit high variance, resulting in many samples contributing minimally. This reduces the Effective Sample Size (ESS) and introduces bias into the estimates.
To address this, the weights are resampled before proceeding with the posterior expectation. This is accomplished by using one of the methods outlined by Douc et al. (2005), whenever the ESS falls below a certain threshold determined by Eq. 26. Resampling redistributes the sample weights, creating a new population with uniformly distributed weights while preserving the statistical properties of the original distribution. As training progresses and the prior becomes a better match for the posterior, the need for resampling decreases.
Importance Sampling estimator
To estimate the expectation, we draw $N_{s}$ independent samples $\{\bm{z}^{(s)}\}_{s=1}^{N_{s}}\sim Q(\bm{z})$ using the procedure outlined in Sec. 2.4.3 and compute a weighted sum:
$\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}[\rho(\bm{z})]$ $\approx{\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})\cdot w_{\text{norm, resampled}}(\bm{\bar{z}_{\text{resampled}}}^{(s)})}.$ (67)
The estimator is unbiased, implying that its expectation matches the true posterior expectation. As $N_{s}\to\infty$, the approximation converges almost surely to the true value, provided that $Q(\bm{z})$ sufficiently covers the posterior to ensure a well-conditioned Radon-Nikodym derivative. This estimator is used for the log-marginal likelihood and its gradient in Eqs. 15 &  16, since it provides a means of estimating expectations with respect to the posterior distribution.
A.6.2 Residual resampling
In this study, we adopted residual resampling to redraw weights. This redistributes the population of latent samples according to the normalized weights of Sec. A.6.1. The procedure is only conducted when Eq. 26 regarding ESS is satisfied:
  1. 1.
    Integer replication: The sample, $s$, is replicated $r_{s}$ times, where:
    $r_{s}=\left\lfloor N_{s}\cdot w_{\text{norm}}(\bm{z}^{(s)})\right\rfloor$ (68)
  2. 2.
    Residual weights: The total number of replicated samples after the previous stage will be $\sum_{s=1}^{N_{s}}r_{s}$. Therefore, $N_{s}-\sum_{s=1}^{N_{s}}r_{s}$ remain, which must be resampled using the residuals:
    $w_{\text{residual}}(\bm{z}^{(s)})$ $=w_{\text{norm}}(\bm{z}^{(s)})\cdot(N_{s}-r_{s})$
    $w_{\text{norm},\ \text{residual}}(\bm{z}^{(s)})$ $=\frac{w_{\text{residual}}(\bm{z}^{(s)})}{\sum_{s=1}^{N_{s}}w_{\text{residual}}(\bm{z}^{(s)})}$ (69)
    If a sample has been replicated at the previous stage, then its corresponding $w_{\text{norm},\,\text{residual}}$ will be reduced, leading to a lower probability of it being resampled again.
  3. 3.
    Resample: The remaining samples are drawn with multinomial resampling, based on the cumulative distribution function of the residuals:
    $k^{*}=\min\left\{j\mid\sum^{j}_{s=1}w_{\text{norm},\,\text{residual}}(\bm{z}^{(s)})\geq u_{i}\right\}\quad\text{where}\quad u_{i}\sim\mathcal{U}([0,1]).$ (70)
    where $k^{*}$ represents the index of the sample to keep.
A.6.3 Metropolis-adjusted Langevin algorithm (MALA)
The Metropolis-Adjusted Langevin Algorithm (MALA) (Rossky et al., 1978a; Zhu & Mumford, 1998), is a Markov Chain Monte Carlo (MCMC) method designed to sample from complex distributions by leveraging gradient information. It constructs a proposal mechanism that efficiently explores the target measure by combining stochastic perturbations with information from the target’s local geometry to propose local moves, thereby refining the standard Metropolis-Hastings (MH) algorithm (Hastings, 1970).
MALA operates on the target (power) posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$. The algorithm constructs a sequence of random variables $\{\bm{z}^{(i,\;t_{k})}\}_{i=1}^{\infty}$ that forms a Markov chain converging in distribution to the target measure. The target posterior measure, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$, is:
$\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})=\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})^{t_{k}}+\log P(\bm{z}\mid\bm{f})-\text{const},$ (71)
MALA
Starting with the initial state of the first chain, $\bm{z}^{(1,\;{t}_{1})}\sim Q(\bm{z})=P(\bm{z}\mid\bm{f})$, the local state within a temperature is updated as follows:
  1. 1.
    Langevin diffusion: A new state $\bm{z}^{\prime\;(i,\;t_{k})}$ is proposed for a local chain, operating with a specific $t_{k}$, using a transition kernel inspired by (overdamped) Langevin dynamics (Roberts & Stramer, 2002; Brooks et al., 2011):
    Target: $\;\,\log\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)\propto\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})^{t_{k}}+\log P(\bm{z}\mid\bm{f})$
    Proposal: $\;\,\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})}\sim q\left(\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})}\right)$
    Kernel: $\;\,q(\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})})=\mathcal{N}\Bigg{(}\bm{z}^{\prime\;(i,\;t_{k})};\;\bm{z}^{(i,\;t_{k})}+\frac{\eta}{2}\bm{C}\cdot\nabla_{\bm{z}}\log\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right),\;\left(\sqrt{\eta}\cdot\bm{C}^{\frac{1}{2}}\right)^{2}\Bigg{)},$ (72)
    where $\eta>0$ is the step size, $\nabla_{\bm{z}}\log\gamma$ is the log-posterior gradient with respect to $\bm{z}^{(i,\;t_{k})}$, and $\bm{C}$ is a pre-conditioning, positive-definite matrix, (for example, the identity matrix, $\bm{I}$).
  2. 2.
    MH criterion: Once $N_{\text{unadjusted}}$ iterations have elapsed, Metropolis-Hastings adjustments, (MH) (Metropolis & Ulam, 1949), are introduced. The criterion for the local proposal is:
    $r_{\text{local}}=\frac{\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)\,q\left(\bm{z}^{(i,\;t_{k})}\mid\bm{z}^{\prime\;(i,\;t_{k})}\right)}{\gamma\left(\bm{z}^{(i,\;t_{k})}\right)\,q\left(\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})}\right)},$ (73)
  3. 3.
    Acceptance: The proposal is accepted with probability $\min(1,r_{\text{local}})$. If accepted, $\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{\prime\;(i,\;t_{k})}$; otherwise, the current state is retained: $\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{(i,\;t_{k})}$.
  4. 4.
    Global Swaps: Global swaps are proposed and accepted subject to the criterion outlined in Eq. 34
Convergence
Under mild regularity conditions, the local Markov chain $\{\bm{z}^{(i,\;t_{k})}\}$ generated by MALA converges in distribution to $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$ as $N_{\text{local}}\to\infty$. The proposal mechanism in Eq. 72 ensures that the chain mixes efficiently, particularly in high-dimensional spaces, provided that the step size $\eta$ is tuned appropriately. The algorithm provides theoretical guarantees of the target, due to its satisfaction of detailed balance under Eq. 73:
$\gamma\left(\bm{z}^{(i,\;t_{k})}\right)\,q\left(\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})}\right)$ $=\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)\,q\left(\bm{z}^{(i,\;t_{k})}\mid\bm{z}^{\prime\;(i,\;t_{k})}\right),$ (74)
ensuring that the stationary distribution of the local chain matches the local target.
A.6.4 autoMALA
More details on the Metropolis-adjusted Langevin Algorithm (MALA) are provided in Sec. A.6.3.
Selecting an appropriate step size for unaltered MALA presents a challenge, particularly for SE, as each power posterior exhibits distinct local curvature. To address this, we adaptively tune the step size with autoMALA (Biron-Lattes et al., 2024), which reinterprets MALA as Hamiltonian Monte Carlo (HMC) with a single leapfrog step (Duane et al., 1987).
The algorithm operates locally on a target power posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$. It constructs a sequence of random variables per temperature, $\{\bm{z}^{(i,\;t_{k})}\}_{i=1}^{\infty}$, that forms a Markov chain converging in distribution to the local target:
$\log\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)=\log P(\bm{z}^{(i,\;t_{k})}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\propto\log P(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})},\bm{\Phi})^{t_{k}}+\log P(\bm{z}^{(i,\;t_{k})}\mid\bm{f}).$
Starting with the initial state of the first chain, $\bm{z}^{(1,\;t_{1})}\sim Q(\bm{z})=P(\bm{z}\mid\bm{f})$, the local state within a temperature is updated as follows:
  1. 1.
    Acceptance thresholds: Two bounding acceptance thresholds are sampled uniformly:
    $a,b$ $\sim\mathcal{U}\Big{(}\left(a,b\right);\bm{0},\bm{1}\Big{)},\quad(a,b)\in[0,1]^{2},\quad b>a$ (75)
  2. 2.
    Mass matrix: Random pre-conditioning matrices are initialized per latent dimension:
    $\varepsilon^{(i,\;t_{k})}$ $\sim\text{Beta}_{\;01}\left(1,1;\left[\frac{1}{2},\frac{2}{3}\right]\right),$
    $M_{p,p}^{1/2\;^{(i,\;t_{k},\;q)}}$ $=\varepsilon^{(i,\;t_{k})}\cdot\Sigma_{p,p}^{-1/2\;{(i,\;t_{k},\;q)}}+(1-\varepsilon^{(i,\;t_{k})}),$ (76)
    where $\Sigma_{p,p}^{(i,\;t_{k},\;q)}=\text{VAR}_{\,s}\left[\bar{z}_{q,p}^{(i,\;t_{k},\,s)}\right]$. Here, $\text{Beta}_{\;01}\left[a,b;c,d\right]$ denotes a zero-one-inflated Beta distribution. The parameters $a,b>0$ represent the shape parameters of the Beta distribution, while $c,d\in[0,1]^{2}$ specify the probabilities governing the mixture. The positive definite mass matrix, $\bm{M}$, can be related to Eq. 72 as $\bm{M}=\bm{C}^{-1}$, and is efficient to invert as a diagonal matrix.
  3. 3.
    Leapfrog proposal: The following transition is proposed for a local chain, $t_{k}$, with an adaptive step size, $\eta^{(i,t_{k})}$:
    $\bm{p}^{(i,\;t_{k})}$ $\sim\mathcal{N}\left(\bm{p};\;\bm{0},\bm{M}^{(i,\;t_{k})}\right)$
    $\bm{p}^{\prime\;(i,\;t_{k})}_{1/2}$ $=\bm{p}^{(i,\;t_{k})}+\frac{\eta^{(i,t_{k})}}{2}\nabla_{\bm{z}}\log\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)$
    $\bm{z}^{\prime\;(i,\;t_{k})}$ $=\bm{z}^{(i,\;t_{k})}+\eta^{(i,t_{k})}\left(\bm{M}^{-1}\bm{p}^{\prime}_{1/2}\right)^{(i,\;t_{k})}$
    $\bm{\hat{p}}^{(i,\;t_{k})}$ $=\bm{p}^{\prime\;(i,\;t_{k})}_{1/2}+\frac{\eta^{(i,t_{k})}}{2}\nabla_{\bm{z}}\log\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)$
    $\bm{p}^{\prime\;(i,\;t_{k})}$ $=-\bm{\hat{p}}^{(i,\;t_{k})},$ (77)
  4. 4.
    MH criterion: Once $N_{\text{unadjusted}}$ iterations have elapsed, Metropolis-Hastings (MH) adjustments are introduced. The MH acceptance criterion for the local proposal is:
    $r_{\text{local}}=\frac{\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)\,\mathcal{N}\left(\bm{p}^{\prime\;(i,\;t_{k})};\;\bm{0},\bm{M}^{(i,\;t_{k})}\right)}{\gamma\left(\bm{z}^{(i,\;t_{k})}\right)\,\mathcal{N}\left(\bm{p}^{(i,\;t_{k})};\;\bm{0},\bm{M}^{(i,\;t_{k})}\right)},$ (78)
  5. 5.
    Step size adaptation: Starting with an initial estimate, $\eta^{(i,t_{k})}_{\text{init}}$, set as the average accepted step size from the previous training iteration, (used as a simple alternative to the round-based tuning algorithm proposed by Biron-Lattes et al. (2024)), the step size is adjusted as follows:
    • If $a<r_{\text{local}}<b$, then $\eta^{(i,t_{k})}=\eta^{(i,t_{k})}_{\text{init}}$
    • If $r_{\text{local}}\leq a$, then $\eta^{\prime\;(i,t_{k})}=\eta^{\;(i,t_{k})}_{\text{init}}/\Delta\eta^{k}$, where $k\leftarrow k+1$ is chosen recursively until $r_{\text{local}}>a$, and the step size is accepted, $\eta^{(i,t_{k})}=\eta^{\prime\;(i,t_{k})}$
    • If $r_{\text{local}}\geq b$, then $\eta^{\prime\;(i,t_{k})}=\eta^{(i,t_{k})}_{\text{init}}\cdot\Delta\eta^{k}$, where $k\leftarrow k+1$ is chosen recursively until $r_{\text{local}}<b$, and the step size is accepted, $\eta^{(i,t_{k})}=\eta^{\prime\;(i,t_{k})}/\Delta\eta$
    Here, $\Delta\eta$ is a tunable hyperparameter that controls the rate of step size adaptation. To improve efficiency in our implementation, tuning is prematurely terminated when $\eta^{(i,t_{k})}$ falls out of a user-defined range, $\eta_{\text{min}},\eta_{\text{max}}]$. Excessively large or small step sizes are not beneficial, so preventing the algorithm from searching extreme values saves time without compromising performance.
  6. 6.
    Reversibility check: If the step size was modified, the reversibility of the update in Eq. 77 is verified before proceeding with MH acceptance. If $\eta^{(i,t_{k})}_{\text{init}}$ cannot be recovered from the reversed step-size adjustment process with the proposed state, $\left\{\bm{z}^{\prime\;(i,\;t_{k})},\;{p}^{\prime\;(i,\;t_{k})},\eta^{(i,t_{k})}\right\}$, the proposal is rejected regardless of the outcome of the MH adjustment.
  7. 7.
    Acceptance: The proposal is accepted with probability $\min(1,r_{\text{local}})$. If accepted, $\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{\prime\;(i,\;t_{k})}$; otherwise, the current state is retained: $\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{(i,\;t_{k})}$.
  8. 8.
    Global Swaps: Global swaps are proposed and accepted subject to Eq. 34
A.7 $20\times 20$ image grids
A.7.1 MNIST
Figure 13: Generated MNIST, (Deng, 2012), after 2,000 parameter updates using MLE / IS adhering to KART’s structure. Uniform, lognormal, and Gaussian priors are contrasted using Radial Basis Functions, (Li, 2024). Lognormal is inverted for clarity.
Figure 14: Generated MNIST, (Deng, 2012), after 2,000 parameter updates using MLE / IS adhering to KART’s structure. Uniform, lognormal, and Gaussian priors are contrasted using Fourier Bases, (Xu et al., 2024). Lognormal is inverted for clarity.
A.7.2 FMNIST
Figure 15: Generated FMNIST, (Xiao et al., 2017), after 2,000 parameter updates using MLE / IS adhering to KART’s structure. Uniform, lognormal, and Gaussian priors are contrasted using Radial Basis Functions, (Li, 2024). Lognormal is inverted for clarity.
Figure 16: Generated FMNIST, (Xiao et al., 2017), after 2,000 parameter updates using MLE / IS adhering to KART’s structure. Uniform, lognormal, and Gaussian priors are contrasted using Fourier Bases, (Xu et al., 2024). Lognormal is inverted for clarity.
A.7.3 Darcy flow
Figure 17: Generated 2D Darcy flow pressures, (Li et al., 2021), after 12,000 parameter updates using MLE / IS while adhering to KART. Radial Basis Functions, (RBFs) (Li, 2024), are contrasted against Fourier bases, (FFTs) (Xu et al., 2024), using Gaussian latent priors. RBF is colored differently for clarity.
Figure 18: Generated 2D Darcy flow pressures, (Li et al., 2021), after 12,000 parameter updates using MLE / IS while adhering to KART. Radial Basis Functions, (RBFs) (Li, 2024), are contrasted against Fourier bases, (FFTs) (Xu et al., 2024), using lognormal latent priors. RBF is colored differently for clarity.
Figure 19: Generated 2D Darcy flow pressures, (Li et al., 2021), after 12,000 parameter updates using MLE / IS while adhering to KART. Radial Basis Functions, (RBFs) (Li, 2024), are contrasted against Fourier bases, (FFTs) (Xu et al., 2024), using uniform latent priors. RBF is colored differently for clarity.
A.7.4 SVHN
(a) SVHN with mixture prior and MLE
(b) SVHN with mixture prior and SE
Figure 20: Generated samples from T-KAM trained on SVHN using a mixture Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ITS prior and ULA posterior sampling.
(a) SVHN with deep prior and MLE
(b) SVHN with deep prior and SE
Figure 21: Generated samples from T-KAM trained on SVHN using a deep Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ULA prior and ULA posterior sampling.
A.7.5 CIFAR-10
(a) CIFAR-10 with mixture prior and MLE
(b) CIFAR-10 with mixture prior and SE
Figure 22: Generated samples from T-KAM trained on CIFAR-10 using a mixture Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ITS prior and ULA posterior sampling.
(a) CIFAR-10 with deep prior and MLE
(b) CIFAR-10 with deep prior and SE
Figure 23: Samples from T-KAM trained on CIFAR-10 using a deep Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ULA prior and ULA posterior sampling.
A.8 Top 3 extracted priors
The following priors were extracted from trained T-KAM models, each corresponding to the first three components of prior $q=1$.
A.8.1 MNIST RBF
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 24: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with RBF bases and Gaussian constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 25: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with RBF bases and lognormal constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 26: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with RBF bases and uniform constrained priors.
A.8.2 MNIST FFT
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 27: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with FFT bases and Gaussian constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 28: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with FFT bases and lognormal constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 29: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with FFT bases and uniform constrained priors.
A.8.3 FMNIST RBF
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 30: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with RBF bases and Gaussian constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 31: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with RBF bases and lognormal constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 32: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with RBF bases and uniform constrained priors.
A.8.4 FMNIST FFT
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 33: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with FFT bases and Gaussian constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 34: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with FFT bases and lognormal constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 35: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with FFT bases and uniform constrained priors.
A.8.5 Darcy RBF
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 36: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with RBF bases and Gaussian constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 37: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with RBF bases and lognormal constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 38: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with RBF bases and uniform constrained priors.
A.8.6 Darcy FFT
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 39: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with FFT bases and Gaussian constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 40: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with FFT bases and lognormal constrained priors.
(a) $q=1,p=1$
(b) $q=1,p=2$
(c) $q=1,p=3$
Figure 41: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with FFT bases and uniform constrained priors.

### v4

@@
 Abstract
-Learning an energy-based model (EBM) in the latent space of a top-down generative model offers an expressive and interpretable framework for text and image generation.
-However, it remains unclear how this interpretability can be used to guide model design, improve generative quality, and reduce training time.
+Learning an energy-based model (EBM) in the latent space of a top-down generative model offers a versatile framework for generation across multiple data modalities.
+However, it remains unclear how its interpretability can be used to guide model design, improve generative quality, and reduce training time.
 Moreover, the reliance on Langevin Monte Carlo (LMC) sampling presents challenges in efficiency and sampling multimodal latent distributions.
-In this work, we propose a novel adaptation of the Kolmogorov–Arnold representation theorem for generative modeling and introduce the Thermodynamic Kolmogorov–Arnold Model (T-KAM) to take advantage of structural and inductive biases.
+In this work, we propose a novel adaptation of the Kolmogorov-Arnold representation theorem for generative modeling and introduce the Thermodynamic Kolmogorov-Arnold Model (T-KAM) to take advantage of structural and inductive biases.
 By constraining the prior to univariate relationships, T-KAM enables fast and exact inference via the inverse transform method.
-With the low dimensionality of the latent space and suitable inductive biases encoded, we demonstrate that importance sampling becomes a viable, unbiased, and highly efficient training strategy.
-We also introduce a training criterion using population-based LMC, which decomposes posterior sampling into a sequence of annealed distributions to improve multimodal exploration.
-T-KAM elegantly balances common trade-offs in generative modeling, offering fast inference, high sample quality, and stable training, while being naturally suited to upcoming Zettascale Computing Co.
-hardware and extendable to other high-impact research directions in generative intelligence.
+With the low dimensionality of the latent space and suitable inductive biases encoded, we demonstrate that importance sampling (IS) becomes a viable, unbiased, and highly efficient posterior sampler.
+For situations where IS fails, we investigate a novel strategy using population-based LMC, which decomposes posterior sampling into a sequence of annealed distributions to improve multimodal sampling.
+T-KAM elegantly balances common trade-offs in generative modeling, offering fast inference, interpretability, and stable training, while being naturally suited to upcoming Zettascale Computing Corp.
+hardware.
 1 Introduction
@@
 Despite being extendable to several high-impact research directions, such as multimodal generative modeling (Yuan et al., 2024), and interpretable classification and meta-learning (Kong et al., 2021), (which involves rapid adaptability and few-shot generalization to out-of-distribution tasks), latent EBMs face critical limitations due to their reliance on MCMC sampling.
-Firstly, MCMC is inherently iterative, and algorithms such as ULA require repeated gradient evaluations through the model, introducing significant computational overhead during both training and inference.
-Secondly, ULA is best suited for sampling from smooth, unimodal distributions and tends to struggle when exploring the multimodal latent distributions that are typical in latent variable generative modeling.
-Moreover, separate ULA samplers are required for prior and posterior sampling, which is problematic given that ULA is sensitive to hyperparameter tuning: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias due to its use of the Euler–Maruyama method to evolve the underlying Langevin equation.
+Firstly, MCMC is inherently iterative, and algorithms such as ULA require repeated gradient passes through the model, which introduces significant computational overhead during both training and inference.
+Secondly, ULA is best suited for sampling from smooth, unimodal distributions and can struggle to explore the multimodal latent distributions that are typical in latent variable generative modeling.
+Moreover, the framework presented by Pang et al.
+(2020) requires separate ULA samplers for the prior and posterior.
+This introduces additional experimental complexity since ULA is sensitive to hyperparameter tuning: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias due to its use of the Euler-Maruyama method (Wilkie, 2004) to evolve the underlying Langevin equation.
 Additionally, it is not evident how the unique separation between the prior and generator networks can be leveraged for interpretability or training efficiency.
@@
 Although ideas from mechanistic interpretability, as reviewed by Bereska & Gavves (2024), could potentially be applied, there is currently no established method to leverage the unique strengths of the EBM prior framework for this purpose.
-1.1 Inductive biases and importance sampling
-To address these limitations, we reimagine the EBM prior framework into a novel univariate or element-wise model, drawing inspiration from Kolmogorov–Arnold Networks (KANs) (Liu et al., 2024) and deep spline networks Bohra et al.
+1.1 Inductive biases and fast sampling
+To address these limitations, we reimagine the EBM prior into a novel univariate or element-wise model, drawing inspiration from Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024) and deep spline networks Bohra et al.
 (2020), which replace the sigmoidal activations used in multilayer perceptrons Murtagh (1991) with spline interpolations.
-By representing the latent space with univariate functions and distributions, we improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning.
+By representing the latent space with univariate functions and distributions, we enable fast and exact prior sampling with the inverse transform method (Devroye, 1986) and improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning.
 This allows the EBM prior to be aligned with the posterior before training begins.
-We show that improved prior–posterior alignment enables efficient training with importance sampling (IS) (Goertzel, 1949), a technique for estimating posterior expectations that is otherwise avoided in the literature due to its degeneracy under severe prior–posterior mismatch and the curse of dimensionality, (Dupuis & Wang, 2004).
-Moreover, the trained prior can be visualized or recovered, offering new opportunities for transfer learning and scientific discovery.
-This is tested in Sec.
-3.1.
-Multiple KAN bases and reference priors are used to verify the effect of inductive bias.
-However, Gaussian Radial Basis Functions (RBF KANs) (Li, 2024), were used instead of the cubic B-spline bases proposed by Liu et al.
+In Sec.
+3.1, we show that adequately chosen priors can enable efficient training with importance sampling (IS) (Goertzel, 1949).
+This is a technique for estimating posterior expectations that is otherwise avoided in the literature due to its high variance under severe prior-posterior mismatch and the curse of dimensionality, (Dupuis & Wang, 2004), which can be avoided in our model due to the low dimensionality of the latent space.
+Moreover, we show that the trained priors are easy to visualize and recover, offering new opportunities for scientific discovery and knowledge transfer to models with different latent structures.
+Multiple KAN bases and reference priors were used to verify the effect of inductive bias on training with IS.
+However, Gaussian Radial Basis Functions (RBF KANs) (Li, 2024), were adopted instead of the cubic B-spline bases proposed by Liu et al.
 (2024), as RBFs are more efficient for GPU implementation, and Poggio & Girosi (1990) previously demonstrated that they provide an optimal solution to the regularized approximation/interpolation problem.
-To derive our new model, which we call the Thermodynamic Kolmogorov-Arnold Model (T-KAM), we first introduce a novel adaptation of the Kolmogorov–Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces.
-Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on heuristics and automated tuning.
-This reinterpretation allows us to frame inverse transform sampling (ITS) (Devroye, 1986) as a measurable transformation applied to uniform noise, thus preserving the deterministic structure of KART while admitting a probabilistic interpretation that aligns with MLE.
+To derive our new model, which we call the Thermodynamic Kolmogorov-Arnold Model (T-KAM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces.
+Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on intuition and automated tuning.
+This reinterpretation allows us to frame inverse transform sampling (ITS) as a measurable transformation applied to uniform noise, thus preserving the deterministic structure of KART while admitting a probabilistic interpretation that aligns with MLE.
 To address the potential inflexibility of strict adherence to KART and the limitations of using independent energy-based priors, we introduce a scalable strategy based on mixture distributions.
@@
 1.2 Population-based Langevin sampling and Thermodynamic training
-However, while ITS enables fast inference, posterior sampling during training must still be conducted using ULA when the IS estimator is insufficient.
+However, while ITS enables fast inference, posterior sampling must still be conducted using ULA when the IS estimator is insufficient.
 Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by Cui & Han (2024) and the diffusion-assisted EBM by Zhang et al.
@@
 To maintain prior interpretability and fast inference post-training, we instead introduce posterior annealing as a novel and alternative strategy.
-This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and sampling with population-based MCMC methods, specifically Parallel Tempering (PT) (Swendsen & Wang, 1986) (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996).
-Although this may reduce expressivity compared to hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training.
-Our training criterion is then derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al.
+This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and recasting ULA with population-based MCMC methods (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996).
+The resulting training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al.
 (2019) to ensure the impact of annealing on learning dynamics and parameter updates.
+Although this reduces expressivity compared to hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, annealing is embarrassingly parallel.
 While ULA is computationally efficient and enables more flexible chain dynamics, it does not preserve detailed balance or reversibility.
 This can be recovered by incorporating Metropolis-Hastings (MH) adjustments (Metropolis & Ulam, 1949) via the Metropolis-adjusted Langevin Algorithm (MALA) (Rossky et al., 1978a).
-The codebase in Sec.
+The code in Sec.
 A.1 optionally supports MALA with locally adaptive step-size tuning, (autoMALA) (Biron-Lattes et al., 2024), to assist when a single step size fails to accommodate the varying geometries of power posteriors, and when the MH criterion deteriorates under the curse of dimensionality, (Hairer et al., 2014).
-MALA and autoMALA are discussed under Secs.
+This is discussed in Secs.
 A.6.3 and A.6.4.
-However, ULA remains the preferred method in our work, motivated by its demonstrated efficiency and effectiveness by Pang et al.
-(2020), and further supported by the findings of Nijkamp et al.
+However, ULA remains the preferred method in our work, motivated by its efficiency and effectiveness, demonstrated by Pang et al.
+(2020) and further supported by the findings of Nijkamp et al.
 (2019b; a), with regards to how short-run, biased MCMC methods can still be highly effective for sampling from EBMs.
-These scaling strategies are compared in Sec.
+MLE and SE are compared in Sec.
 3.2.
-Despite our optimizations, the element-wise and localized nature of KAN functions complicate their parallelization on GPUs, which are best suited to matrix arithmetic.
+Despite our optimizations, the element-wise and localized nature of KANs complicate their parallelization on GPUs, which are best suited to matrix arithmetic.
 As a result, forward passes through T-KAM and the gradient evaluations required for ULA were found to be slow, necessitating reduced experimental complexity.
-We ultimately found that our mixture prior model outperformed a deep KAN prior architecture closely aligned with the model used in the study by Pang et al.
-(2020), while also offering greater computational efficiency.
-In contrast, the use of annealing could not be justified due to its significant computational overhead and limited gains in quality.
-However, this assessment may change as a greater number of power posteriors and more sophisticated posterior sampling algorithms become feasible with the emerging hardware advances proposed by Zetta (2024).
+The experiments in Sec.
+3.2 demonstrate that our mixture prior model is able to generate coherent, (although low-quality), samples after a few training epochs on RGB images, while still supporting efficient, exact sampling via component-wise ITS.
+In contrast, annealing incurred substantial computational overhead, but produced little to no improvements in sample quality, rendering its use difficult to justify.
+However, this result may change as a larger number of power posteriors, longer training durations, and more advanced posterior sampling algorithms become practical with the hardware advances proposed by Zetta (2024).
 Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting T-KAM at scale, while improving the potential of univariate priors with its function fusion features.
@@
 This section introduces the model and its associated sampling procedures.
-We begin by presenting a novel interpretation of the Kolmogorov–Arnold Representation Theorem (KART) to define the model’s baseline architecture.
+We begin by presenting a novel interpretation of the Kolmogorov-Arnold Representation Theorem (KART) to define the model’s baseline architecture.
 Our key practical contribution is the introduction of a univariate energy-based prior, detailed in Sec.
 2.1.
-The top-down generator can be instantiated in various forms, including Kolmogorov–Arnold Networks (KANs), which enables strict adherence to KART.
+The top-down generator can be instantiated in various forms, including Kolmogorov-Arnold Networks (KANs), which enables strict adherence to KART.
 We then propose scalable extensions to the model, including mixture distributions in Sec.
@@
 2.5 to improve multimodal sampling during training.
-The Kolmogorov–Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of T-KAM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
+The Kolmogorov-Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of T-KAM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
 The Kolmgorov-Arnold Representation theorem 1
@@
 To ensure adherence to it, any transformations must be derived to be continuous, measurable, and deterministic.
-For readers less familiar with the technical background, the two most practically significant equations of Sec..
+For readers less familiar with the technical background, the two most practically significant equations of Sec.
 2.1 are Eqs.
@@
 Here, $\delta_{\psi_{q,p}(u_{p})}$ denotes the Dirac measure centred at $\psi_{q,p}(u_{p})$, ensuring the kernel is deterministic: i.e., once $u_{p}$ is sampled, the outcome in $\mathcal{Z}$ is fixed.
-This measurability requirement guarantees that the pushforward from $([0,1],\mathcal{B}([0,1]),\mu)$ to $(\mathcal{Z},\mathcal{B}(\mathcal{Z}),\pi)$ is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$.
+This guarantees that the pushforward from $([0,1],\mathcal{B}([0,1]),\mu)$ to $(\mathcal{Z},\mathcal{B}(\mathcal{Z}),\pi)$ is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$.
 Here, $\mathcal{B}(\mathcal{Z})$ is the Borel $\sigma$-algebra on the latent space, $\mathcal{Z}$, and $\pi$ is the latent measure.
@@
 $Z_{q,p}=\int_{\mathcal{Z}}\exp(f_{q,p}(z))\,d\pi_{0}(z)<\infty.$ (5)
-In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov–Arnold Networks (KANs) (Liu et al., 2024).
+In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024).
 For instance, smoothing splines, such as those used by KANs or deep spline networks, (Bohra et al., 2020).
@@
 The partition function $Z_{q,p}$ is often intractable in energy-based models.
-However, in this univariate setting, where the energy function is smooth and defined on a bounded domain $\left[z_{\text{min}},z_{\text{max}}\right]$, it can be accurately approximated using Gauss–Kronrod quadrature, (Laurie, 1997):
+However, in this univariate setting, where the energy function is smooth and defined on a bounded domain $\left[z_{\text{min}},z_{\text{max}}\right]$, it can be accurately approximated using Gauss-Kronrod quadrature, (Laurie, 1997):
 $Z_{q,p}$ $=\int_{z_{\text{min}}}^{z_{\text{max}}}\exp(f_{q,p}(z))\pi_{0}(z)dz\approx\sum_{i=1}^{N_{\text{quad}}}w_{i},G(z_{i}^{\text{node}}),$ (6)
@@
 Strict adherence to KART can be achieved for hidden layer $h_{l_{k},l_{k-1}}$ by applying either a sigmoid or affine transformation after layer $h_{l_{k-2},l_{k-3}}$ to shift its range back to $0,1]^{L_{k-1}}$.
-However, using sigmoid may lead to vanishing gradients and we contend that function domains are less critical to KART than its structural properties and should not be prioritized over maintaining gradient flow and avoiding saturation.
+However, using sigmoid may lead to vanishing gradients and we suggest that function domains are less critical to KART than its structural properties and should not be prioritized over maintaining gradient flow and avoiding saturation.
 Adopting a KAN-based generator also allows for training with L1 regularization, followed by pruning with the method proposed by Liu et al.
@@
 However, this violates adherence to KART, since $F$ now corresponds to a discrete rather than continuous CDF.
+The mixture proportions are learned by replacing the log-prior in Eq.
+18 with the logarithm of Eq.12.
 To encourage balanced weighting and reduce model complexity, L1 regularization can be applied to $\alpha_{q,p}$ and small values can be pruned.
-The mixture proportions are learned by adding $\log\alpha_{q,p}$ to the gradients of the prior in Sec.
-2.4.1.
 Component selection follows a categorical distribution over $p=1,\dots,n_{z}$, determined by $\alpha_{q,p}$.
@@
 2.3.3 Deep KAN prior
-If required, we may further deepen the energy functions $f_{q,p}$ using the KAN approach.
+If required, the energy functions $f_{q,p}$ can be deepened using the KAN approach.
 In fact, each $f_{q,p}$ can itself be a realization of Eq.
@@
 2.5.2.
+This is not investigated in the experiments, since it is simply equivalent to the model of Pang et al.
+(2020), reformulated as a KAN.
 2.4 Maximum likelihood training
@@
 The learning gradient attributed to the log-likelihood can also be evaluated using autodifferentiation, after first approximating the posterior expectation with IS or MC estimators.
+The code in Sec.
+A.1 also provides an option to augment the first likelihood model with an additional perceptual loss term (Johnson et al., 2016) using features extracted from the pretrained VGG-16 CNN (Simonyan & Zisserman, 2015).
+This could have yielded higher-fidelity images for Sec.
+3.2, but it was not adopted in this study due to the computational overhead of Langevin sampling.
 2.4.3 Prior inverse transform sampling
@@
 High-variance weights with a low ESS indicate that many samples contribute minimally, resulting in inefficient sampling.
-Resampling focuses learning on samples that are more representative of the posterior distribution, thereby maximizing the usefulness of each parameter update.
+Resampling focuses learning on samples that are more representative of the posterior distribution, which maximizes the usefulness of each parameter update.
   4. 4.
@@
     $\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}[\rho(\bm{z})]\approx{\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})\cdot w_{\text{norm}}(\bm{\bar{z}_{\text{resampled}}}^{(s)})}.$ (27)
-This estimator is used for computing the log-marginal likelihood and its gradient in Eqs.
+This is used to estimate the log-marginal likelihood and its gradient in Eqs.
 15 and 16.
@@
 2.5.1 Thermodynamic Training Criterion
-We inform our derivations using the Thermodynamic Integral. In place of the log-marginal likelihood outlined in Eq. 15, this exploits the following representation derived in Sec. A.4.5:
+We inform our derivations using the Thermodynamic Integral. In place of the log-marginal likelihood outlined in Eq. 15, this uses the following representation derived in Sec. A.4.5:
 $\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)\right]\,dt$ (28)
@@
 28 is feasible given that the bounds are $t=0$, corresponding to the expectation being taken with respect to the prior, and $t=1$, where it is taken with respect to the posterior.
-However, the integral can also be expressed exactly without collapsing the temperatures, offering greater flexibility in how the integral evolves through the temperature schedule, $\{t_{k}\}_{k=0}^{N_{t}}$.
+However, the integral can also be expressed exactly without collapsing the temperatures.
+This allows greater flexibility in how the integral evolves through the temperature schedule, $\{t_{k}\}_{k=0}^{N_{t}}$.
 In particular, the temperature schedule is discretized as:
@@
 (2020) in using the Unadjusted Langevin Algorithm (ULA) (Rossky et al., 1978b; Brooks et al., 2011) without MH correction.
-Specifically, we update samples using the forward Euler–Maruyama discretization of the Langevin SDE, as detailed by Roberts & Stramer (2002), with the transition kernel described in Sec.
+Specifically, we update samples using the forward Euler-Maruyama discretization of the Langevin SDE, as detailed by Roberts & Stramer (2002), with the transition kernel described in Sec.
 A.6.3:
@@
 $\log\gamma\left(\bm{z}^{(i,\;t_{k})}\right)=\log P(\bm{z}^{(i,\;t_{k})}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\propto t_{k}\log P(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})},\bm{\Phi})+\log P(\bm{z}^{(i,\;t_{k})}\mid\bm{f}).$
-ULA is best suited for smooth, unimodal distributions, but annealing with power posteriors might improve sampling from multimodal distributions.
+ULA is best suited for smooth, unimodal distributions, but annealing with power posteriors can improve sampling from multimodal distributions.
 Mixing can be further improved by periodically swapping populations between randomly selected adjacent temperatures, subject to:
@@
 In Sec.
-3.2, we compare the mixture EBM prior sampled using ITS against a deep EBM prior sampled using ULA.
-The latter shares an architectural design similar to that used by Pang et al.
-(2020), but with Chebyshev KAN (SS et al., 2024) energy functions.
-We further compare T-KAM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion.
-These comparisons serve to evaluate the practical strategies proposed for scaling and training T-KAM, as well as to compare ITS and our mixture EBM against ULA and deep EBMs, previously used by Pang et al.
+3.2, we evaluate the mixture EBM prior, sampled via component-wise ITS and paired with CNN-decoder architectures, which are smaller variants of those used by Pang et al.
 (2020).
+We additionally compare T-KAM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion.
+These comparisons assess the practicality of using mixture distributions for scaling T-KAM and using annealing to improve training for high-dimensional RGB images.
 To quantify the fidelity of generated data in a more reliable manner than previous studies in the literature, we used effectively unbiased metrics based on the principles outlined by Chong & Forsyth (2020).
@@
 Furthermore, no metrics were applied in Secs.
-3.1.1 &  3.1.1, as vision metrics rely on the Inception V3 model trained on ImageNet (Szegedy et al., 2015), potentially introducing bias against other datasets.
-More image samples are provided in Sec. A.7.
-3.1 Inductive biases enable Importance Sampling
+3.1.1 &  3.1.1, as vision metrics rely on the Inception V3 model trained on ImageNet (Szegedy et al., 2015), potentially introducing bias against other datasets, especially non-RGB images.
+3.1 Importance Sampling
 3.1.1 MNIST & FMNIST
@@
 4.
-T-KAM was trained for 10 epochs with 20,000 training examples, an importance sample size of $N_{s}=100$, and a batch size of $N_{x}=100$, (corresponding to 2,000 parameter updates) Prior sampling was conducted with ITS and posterior expectations were estimated with IS.
+T-KAM was trained for 10 epochs with 50,000 training examples, an importance sample size of $N_{s}=100$, and a batch size of $N_{x}=100$, (corresponding to 5,000 parameter updates).
+Prior sampling was conducted with ITS and posterior expectations were estimated with IS.
 The architecture was guided by KART and independent priors were used.
-(a) Uniform
-(b) Lognormal
-(c) Gaussian
-Figure 2: Exponentially-tilted priors, extracted after training on MNIST with RBFs.
-These plots illustrate how T-KAM can be visualized and interpreted by how its priors are adjusted relative to their reference distributions.
-The lognormal prior remains untrained due to posterior collapse, while the uniform prior has been skewed towards $z_{1,2}=1$.
-Gaussian is only slightly tilted.
-(a) MNIST
-(b) FMNIST
-Figure 3: Generated NIST images using RBFs and different exponentially-tilted priors, initialized from $\mathcal{N}(\bm{z};\;\bm{0},\bm{1})$, $\text{Lognormal}(\bm{z};\;\bm{0},\bm{1})$, and $\mathcal{U}(\bm{z};\;\bm{0},\bm{1})$.
-Lognormal is inverted for clarity.
-Lognormal priors resulted in posterior mode collapse.
-T-KAM learned to generate an average of all MNIST or FMNIST classes, and failed to learn a meaningful latent prior.
-In contrast, Gaussian priors produced the sharpest images for FMNIST, and uniform yielded the sharpest MNIST digits.
-In Fig.
-2, we show how T-KAM allows visualization of its univariate latent distributions, which is an interpretability advantage that can help us explain the observed results.
-Unlike Gaussian and uniform reference distributions, which provide a more symmetric and balanced initial spread over the latent space, lognormal priors introduce a strong right-skew.
-This asymmetric initialization makes it difficult for T-KAM to learn flexible, separable representations, ultimately impairing its ability to produce binary output features.
-In contrast, the symmetric nature of Gaussian and uniform priors allows the sculpting of latent distributions and their skews to support binary generation.
-When cross-referencing with the priors of other latent features plotted in Sec.
-A.8, it becomes evident that left-skewed distributions are more successful at MNIST and FMNIST generation.
-This explains why the standard lognormal prior, which is right-skewed, is poorly suited to these datasets.
-The experiment also demonstrates the validity of importance sampling for posterior expectations when prior-posterior alignment can be improved by choosing an appropriate reference prior, since uniform and Gaussian priors produced realistic MNIST and FMNIST samples respectively.
+(a) Real samples
+(b) No ref. prior, FFT basis
+(c) Gaussian ref. prior, RBF basis
+Figure 2: Generated MNIST examples using different bases and exponentially-tilted priors
+(a) Real samples
+(b) No ref. prior, RBF basis
+(c) Gaussian ref. prior, RBF basis
+Figure 3: Generated FMNIST examples using different bases and exponentially-tilted priors
+T-KAM generated a diverse set of samples when a non-restrictive reference prior was used, (Gaussian or no reference), thus validating importance sampling and independent energy-based priors for simple datasets.
+Subjectively, the best MNIST samples arose with no reference prior and RBF bases, whereas the best FMNIST samples came from a Gaussian reference prior with RBF bases, potentially because FMNIST has smoother features than MNIST.
+The additional samples plotted in Sec.
+A.7.1 illustrate that the choice and combination of bases and reference priors are important.
+Poorly chosen combinations led to posterior collapse and low-quality samples, thereby showcasing the effect of inductive bias in guiding sample quality.
+In Sec.
+A.8, examples of trained univariate latent distributions have been plotted to demonstrate T-KAM’s interpretability.
+Learning can be understood by how T-KAM reshaped its latent representation relative to the initial reference prior.
+For instance, lognormal priors showed little change, indicating posterior collapse, and FFTs caused much sharper relationships than the smoother RBFs.
 3.1.2 2D Darcy flow pressures
@@
 However, in this study, we use the purely generative task to illustrate how weak inductive biases can influence what T-KAM learns.
-T-KAM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=50,N_{s}=50$.
-The training usespanned 600 epochs with 1,000 training examples, resulting in 12,000 parameter updates.
+T-KAM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=100,N_{s}=100$, and the first likelihood model in Eq.
+20.
+Training spanned 1,000 epochs with 1,000 training examples, resulting in 10,000 parameter updates.
 All other hyperparameters are provided in Tab.
@@
 The architecture was guided by KART and independent priors were used.
-(a) $\mathcal{U}(\bm{z};\;\bm{0},\bm{1})$
-(b) $\text{Lognormal}(\bm{z};\;\bm{0},\bm{1})$
-Figure 4: Generated Darcy flow pressures with exponentially-tilted priors initialized from standard uniform and lognormal distributions.
-RBF bases, are contrasted against FFT bases, where RBF is colored differently for clarity.
-(a) RBF prior on $z_{1,1}$
-(b) FFT prior on $z_{1,1}$
-Figure 5: Comparison of exponentially-tilted lognormal priors, extracted after training on Darcy between RBF bases and FFT bases.
-RBF basis functions led to oversmoothing and posterior collapse, while FFT bases lacked smoothness, potentially compromising adherence to the continuous assumptions imposed by KART.
-Although RBFs tended to oversmooth the solution space, FFTs were more effective at capturing large-scale spatial periodicity.
-In cases where the prior was misspecified with a lognormal reference, RBFs failed to produce coherent samples, while FFTs still yielded coherent outputs.
-This result likely stems from their suitability in representing periodic spatial relationships, a property that aligns well with the structure of the Darcy flow, as previously showcased by the success of the FNO, which also uses Fourier methods.
+(a) Real samples
+(b) Gaussian ref. prior, FFT basis
+(c) Uniform ref. prior, RBF basis
+Figure 4: Generated Darcy flow examples using different bases and exponentially-tilted priors
+RBFs tended to oversmooth the solution space while FFTs were more effective at capturing large-scale spatial periodicity.
+This likely stems from their suitability in representing periodic spatial relationships, a property that aligns well with the structure of the Darcy flow, as previously showcased by the success of the FNO, which also uses Fourier methods.
 This is somewhat reflected by Fig.
-5(b), which depicts some periodicity in the learned, exponentially-tilted distribution on $z_{1,1}$.
-However, FFT-parameterized energy functions were also found to be non-smooth and potentially discontinuous, which is also illustrated by the distribution plotted in Fig.
-5(b).
+32, which depicts some periodicity in the learned, exponentially-tilted distributions on $z_{q,p}$.
+However, FFT-parameterized energy functions were also found to be non-smooth and potentially discontinuous, which is also illustrated by the distributions plotted in Fig.
+32.
 This might have compromised adherence to KART and future work might explore incorporating regularization to improve their smoothness.
-This shows that the choice of basis functions serves as an additional mechanism for encoding inductive bias, complementing the role of the reference prior.
-Once again, importance sampling proved to be viable when appropriate inductive biases were encoded.
-3.2 Comparing scaling strategies against CIFAR-10 & SVHN
+Overall, the results show that the choice of basis functions serves as an additional mechanism for encoding inductive bias, complementing the role of the reference prior.
+Once again, importance sampling only proved to be viable when appropriate inductive biases were encoded, and some combinations of bases and reference priors failed to learn meaningful features, (demonstrated by the extra samples of Sec.
+A.7.3).
+3.2 Langevin Monte Carlo and Steppingstone estimator
 T-KAM, when adherent to KART, can be interpreted as a dense or spline-based network.
@@
 In addition to their computational efficiency through weight sharing, CNNs offer a key advantage over dense networks due to their spatial inductive bias: sequences of convolutional layers implement shift-equivariant transformations that are well-suited for images.
+We evaluated T-KAM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015).
+The first likelihood model in Eq.
+20 was adopted for both datasets.
 Due to limited computational resources during this study, (detailed in Sec.
-A.2), training was shortened from 70,000 to 8,000 parameter updates.
-Given their parameter efficiency, Chebyshev bases (SS et al., 2024) were used for the prior energy functions.
-The deep prior architecture mirrored that of Pang et al.
-(2020), but it is important to note that dense KANs do not directly replicate the behavior of standard dense layers.
-Additionally, we used a single optimizer to train both models, rather than using separate optimizers for the prior and generator networks.
-We considered these discrepancies to be acceptable, given the comparative nature of this experiment.
-Posterior sampling was conducted with ULA, and a Gaussian reference was used, $\pi_{0}=\mathcal{N}(\bm{z};\;\bm{0},\bm{1})$.
-T-KAM was trained for 20 epochs with a batch size of $N_{x}=100$ on 40,000 training examples.
+A.2), training was shortened to 8,000 parameter updates, (70,000 were used in the study by Pang et al.
+(2020)), and smaller versions of the CNN decoders used by Pang et al.
+(2020) were adopted as the generator architectures.
+Additionally, a single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator networks.
+Chebyshev KAN bases (SS et al., 2024) were used for the prior energy functions, and full architectural details are provided under Sec.
+A.5.1.
+Posterior sampling was conducted with ULA, and the standard normal reference prior was used, $\pi_{0}=\mathcal{N}\left(\bm{z};\;\bm{0},\bm{1}\right)$.
+T-KAM was trained for 40 epochs with a batch size of $N_{x}=100$ on 20,000 training examples.
 All remaining hyperparameters are provided in Tab.
-6, and architectural details for the mixture and deep prior models, as well as the CNN generators, are included in Sec.
-A.5.1.
-Annealing was conducted with $N_{t}=5$ power posteriors and the cyclic temperature scheduling outlined in Sec.
+6.
+Annealing was conducted with $N_{t}=10$ power posteriors and the cyclic temperature scheduling outlined in Sec.
 A.3.1.
-For brevity, only the SVHN images are visualized in the main body. Samples of CIFAR-10 are provided under Sec. A.7.5
-Table 1: Effectively unbiased FID and KID scores (lower is better) for SVHN.
-Prior MLE SE
-Mixture $\overline{\text{FID}}_{\infty}=81.02$ $\overline{\text{FID}}_{\infty}=82.13$
-$\overline{\text{KID}}_{\infty}=0.0669$ $\overline{\text{KID}}_{\infty}=0.0673$
-Deep $\overline{\text{FID}}_{\infty}=107.42$ $\overline{\text{FID}}_{\infty}=141.24$
-$\overline{\text{KID}}_{\infty}=0.0970$ $\overline{\text{KID}}_{\infty}=0.1320$
-Table 2: Effectively unbiased FID and KID scores (lower is better) for CIFAR-10.
-Prior MLE SE
-Mixture $\overline{\text{FID}}_{\infty}=159.11$ $\overline{\text{FID}}_{\infty}=180.00$
-$\overline{\text{KID}}_{\infty}=0.1600$ $\overline{\text{KID}}_{\infty}=0.1858$
-Deep $\overline{\text{FID}}_{\infty}=185.10$ $\overline{\text{FID}}_{\infty}=201.75$
-$\overline{\text{KID}}_{\infty}=0.1780$ $\overline{\text{KID}}_{\infty}=0.1977$
-(a) SVHN with mixture prior and MLE
-(b) SVHN with mixture prior and SE
-Figure 6: Generated samples from T-KAM trained on SVHN using a mixture prior.
-(c) SVHN with deep prior and MLE
-(d) SVHN with deep prior and SE
-Figure 7: Generated samples from T-KAM trained on SVHN using a deep prior.
-The mixture prior with ITS outperformed the deep prior with ULA, while also being more computationally efficient and enabling faster inference, (Tab.
-9).
-This advantage is likely due to ITS providing exact samples from the prior, whereas ULA produces only approximate, biased samples.
-SE did not outperform MLE and cannot be justified in this setting given its computational overhead.
+Table 1: Effectively unbiased FID and KID scores (lower is better) for SVHN ($32\times 32$).
+MLE SE
+$\overline{\text{FID}}_{\infty}=97.90$ $\overline{\text{FID}}_{\infty}=145.64$
+$\overline{\text{KID}}_{\infty}=0.0827$ $\overline{\text{KID}}_{\infty}=0.1407$
+Table 2: Effectively unbiased FID and KID scores (lower is better) for CelebA ($64\times 64$).
+MLE SE
+$\overline{\text{FID}}_{\infty}=226.88$ $\overline{\text{FID}}_{\infty}=209.30$
+$\overline{\text{KID}}_{\infty}=0.2590$ $\overline{\text{KID}}_{\infty}=0.2138$
+(a) Real samples
+(b) Maximum Likelihood
+(c) Steppingstone Estimator
+Figure 5: Generated SVHN ($32\times 32$) examples using mixture prior architecture with inverse transform sampling from the prior and Langevin sampling from the posterior
+(a) Real samples
+(b) Maximum Likelihood
+(c) Steppingstone Estimator
+Figure 6: Generated CelebA ($64\times 64$) examples using mixture prior architecture with inverse transform sampling from the prior and Langevin sampling from the posterior
+Despite being trained with reduced complexity for a small iteration count, our mixture prior model was able to generate low quality samples of SVHN and CelebA.
+This demonstrates the potential of component-wise ITS for exact and highly efficient sampling in replacement to the approach of Pang et al.
+(2020), which relied on ULA to sample from the prior, thus introducing bias and inefficiency.
+SE only resulted in better image quality for CelebA and cannot be justified in this setting given its computational overhead.
 Alternative approaches, such as different annealing schedules or the use of autoMALA with bias correction and temperature-adaptive step size tuning, could improve performance, but they are not worth investigating until the hardware of Zetta (2024) becomes available, due to the added complexity, hyperparameter tuning, and extended training time.
-The results confirm the superiority of mixture EBMs paired with ITS over deep EBMs sampled with ULA as a viable approach to scaling.
-However, the advantages of SE and population-based ULA remain inconclusive.
 4 Conclusions
 T-KAM provides a principled and scalable framework for structured and informed generative modeling.
-Our results demonstrate that incorporating inductive biases through both basis functions and reference priors can enable more efficient training strategies, such as importance sampling.
-We also show how the Kolmogorov–Arnold representation theorem can inform architectural design and introduce structural bias.
-Additionally, we find that the inverse transform method offers clear advantages over Langevin Monte Carlo sampling for energy-based priors.
-While mixture priors with inverse transform sampling consistently outperformed alternative approaches across both datasets, the benefits of annealing and population-based MCMC remain inconclusive, suggesting directions for further work.
+Our results demonstrate that correctly chosen basis functions and reference priors can enable more efficient training on simple or well-understood datasets through importance sampling.
+We also show how the Kolmogorov-Arnold representation theorem can inform architectural design and introduce structural bias, allowing T-KAM’s architecture to be guided entirely by the size of its latent space.
+Moreover, we find that the inverse transform method offers clear advantages over Langevin Monte Carlo sampling for energy-based priors in terms of speed and exact sampling.
+However, the benefits of annealing and population-based MCMC remain inconclusive, suggesting directions for further work.
 Future investigations into the practicality of our annealing methods depend on continued hardware advancements, such as those introduced by Zetta (2024).
@@
 4.1.1 Hardware acceleration
-Computational benchmarks are provided in Sec.
+Benchmarks are provided in Sec.
 A.2.
-The univariate and localized nature of spline interpolations complicates their parallelization on GPUs, since they involve sparse or recursive computations.
-Additionally, the search operations in Eqs.
-1323, and 70 were parallelized on the CPU across samples, despite the remainder of T-KAM being executed on the GPU.
-This design stems from the fact that GPUs are generally ill-suited for parallelized search algorithms, which often rely on conditional logic.
-Parallelizing conditionals can lead to thread divergence, making CPUs more efficient for these operations.
-While ITS enables fast inference and scales when the prior is a mixture model, T-KAM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA.
+The univariate, localized nature of spline interpolations complicates their parallelization on GPUs, since they involve sparse and recursive computations.
+The search operations in Eqs.
+1323, and 69 were also difficult to parallelize on GPUs, which are generally ill-suited for search algorithms, as they often rely on conditional logic.
+Parallelizing conditionals can lead to thread divergence, making GPUs inefficient for these operations.
+Additionally, while ITS enables fast inference and scales when the prior is a mixture model, T-KAM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA.
+GPUs are not well suited to Langevin Monte Carlo methods due to the irregular memory access patterns and bandwidth pressure caused by differentiating with respect to inputs, which requires storing the outputs of each layer within the generator.
 Nonetheless, the unique advantages of T-KAM justify continued research and development.
-Notably, Zetta (2024) introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control and high throughput.
+Moreover, Zetta (2024) introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control that can mitigate the challenges of sparsity, recursion, and irregular memory access patterns.
 The LFU presents a promising avenue for more efficient implementations of ULA, and improved posterior sampling, potentially including bias correction and adaptive step size tuning with autoMALA, (Sec.
@@
 Overall, T-KAM may prove more scalable than conventional generative models, owing to its reliance on finite representations of continuous functions and its compatibility with the LFU.
-4.1.2 Text-to-text, classification, and multi-modal generation
-Kong et al.
-(2021) provided a robust framework for high-quality text-to-text generation, classification, and meta-learning, while Yuan et al.
-(2024) demonstrated how T-KAM might be adapted into a multi-modal generative model.
-This holds promise for future investigations and industry adoption, especially when the LFU becomes available.
-However, the symbol-vector coupling approach of Kong et al.
-(2021) must first be adapted to remain compatible with ITS.
-4.1.3 Improved sampling for multimodal distributions
-Despite its theoretical guarantees, annealing is difficult to justify given its computational cost and limited practical benefit.
-Future work could explore adaptations of the hierarchical EBMs proposed by Cui & Han (2024), without sacrificing the interpretability or efficiency of our model.
+4.1.2 More opportunities for importance sampling
+For SVHN and CelebA, IS proved infeasible as a posterior sampler, so ULA was used for Sec.
+3.2.
+This may be attributed to prior–posterior mismatch, which rendered the IS estimator ineffective.
+Moreover, relying on IS restricts the prior to its initial domain, since proposals are generated from the prior using ITS and finite-domain KAN energy functions.
+In contrast, ULA can explore values of $\bm{z}$ beyond this domain, enabling adaptation of the prior domain within the grid-updating scheme for KANs proposed by Liu et al.
+(2024).
+A possible remedy for the curse of dimensionality is to learn in a reduced-order space using Principal Component Analysis (PCA) (Jolliffe & Cadima, 2016), similar to the approach of Kingma et al.
+(2014), who also applied PCA to train VAEs.
+To address prior–posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins.
+Additionally, efforts should be directed toward selecting an appropriate initial domain or adapting the domain dynamically during training with IS.
+4.1.3 Expressivity
+The requirements for ITS outlined in Sec.
+2.1 are quite restrictive, which may limit expressivity at scale.
+Additionally, it is often desirable to enrich the latent space with more expressive distributions.
+For example, Normalizing Flows (Papamakarios et al., 2021) achieve this by applying a sequence of reversible transformations that progressively map a simple base distribution into a complex one.
+This might be achievable in T-KAM by introducing more latent distributions, or revising Eq.
+12 into a Mixture of Experts framework (Cai et al., 2024) where $\alpha_{q,p}$ could be realized as a more flexible kernel to enable different inductive biases.
+However, this might be expensive for the model by Pang et al.
+(2020) since its use of Langevin sampling is computationally intense.
+For example, the previous work of Cui & Han (2024), which divides their latent EBM space using the diffusion strategy by Ho et al.
+(2020), states that sampling is expensive due to their use of MCMC methods.
+T-KAM could potentially circumvent the additional cost by using IS as a fast sampler, though this would require addressing its current limitations.
 Appendix A Appendix
 A.1 Codebase
-The study was implemented at https://github.com/exa-laboratories/T-KAM.
+The study was implemented at https://github.com/thezettascale/T-KAM.
 A.2 Computational benchmarks
-These benchmarks were collected for FP32 on a GPU (NVIDIA GeForce GTX 1650) with Julia’s BenchmarkTools package, (JuliaCI, 2024).
+These benchmarks were collected for FP32 on a GPU (NVIDIA GeForce RTX 4060) with Julia’s BenchmarkTools package, (JuliaCI, 2024).
 The index search operations in Eqs.
-2313, & 70 were parallelized across 12 threads on a CPU (i7-9750H @ 2.60GHz).
+2313, & 69 were parallelized across 20 threads on a CPU (i9-13900H @ 5.40GHz).
 Figure 7: Benchmarks for varying latent dimensions, $n_{z}$, while sampling from independent EBM priors with ITS and estimating posterior expectations with IS.
@@
 The generator received inputs of size $2n_{z}+1$, with a hidden size of $4n_{z}+2$ and an output size of $28\times 28$.
-All other hyperparameters were kept consistent with Tab.
-4.
+All other hyperparameters were consistent with Tab.
+4 (RBF bases).
 They reflect the performance of a reverse autodifferentiation of the MLE criterion, Eq.
 16.
-Figure 9: Benchmarks for varying amounts of ULA iterations, $N_{prior}$, to sample from a deep EBM prior (Tab.
-8).
-Posterior expectations were estimated with an Monte Carlo estimator of the MLE criterion, Eq.
-16.
-All other hyperparameters were kept consistent with Tab.
-6.
-They reflect the computational cost of generating samples from the prior with Eq.
-33.
-The benchmarks for ULA appear relatively constant across different step counts due to Julia’s loop optimizations and efficient memory management.
-Nevertheless, the reference ITS sampler remains significantly faster in practice.
-Note: Since the reference ITS sampling executed very rapidly, its relatively high proportion of time attributed to garbage collection is not unexpected.
-Figure 10: Benchmarks for varying amounts of temperatures, $N_{t}$, while using a mixture of EBM priors (Tab.
+Figure 8: Benchmarks for varying amounts of temperatures, $N_{t}$, using a mixture of EBM priors (Tab.
 7) with ITS sampling.
@@
 All other hyperparameters were kept consistent with Tab.
-6.
+6 (Celeb-A).
 They reflect the computational cost of performing a single reverse autodifferentiation of the SE criterion, Eq.
 31.
-Note: Since the reference MLE criterion is executed very rapidly, its relatively high proportion of time attributed to garbage collection is not unexpected.
 A.3 Guidance on tempering
@@
 This overlap is quantified by the KL divergence term in Eq.
-52.
+51.
 Maintaining small KL divergences between adjacent power posteriors can be achieved by clustering temperatures in regions where $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$ changes rapidly.
@@
 The effect of varying $p$ on the schedule between 0 and 1 is visualized in Fig.
-11.
+9.
 When $p>1$, spacing becomes denser near $t=0$, while $p<1$ creates denser spacing near $t=1$.
@@
 51 is illustrated in Fig.
-12.
+10.
 The plots demonstrate that for $p>1$, the trapezoidal bins are clustered closer to $t=0$, concentrating the discrete approximation error towards $t=1$.
@@
 However, their findings may not hold in the context of T-KAM and generative modeling, making $p$ and the temperature schedule difficult variables to tune.
-Instead, we schedule $p$ cyclically across all parameter updates of T-KAM’s training procedure, similar to the previous work of Fu et al.
-(2019).
-The initial and final $p$’s and the number of cycles are treated as tunable hyperparameters:
-$\delta$ $=\text{linspace}\left(0,\ 2\pi(N_{\text{cycles}}+0.5);\ \text{len}=N_{\text{updates}}\right)$
-$p(i)$ $=p_{\text{init}}+\frac{p_{\text{end}}-p_{\text{init}}}{2}\left(1-\cos\delta\right)$ (37)
-where $N_{\text{updates}}$ denotes the total number of parameter updates, $i$ indexes the current training iteration, $N_{\text{cycles}}$ is the number of cycles, and $\delta$ is a vector of length $N_{\text{updates}}$.
+Instead, we schedule $p$ sinusoidally across all training parameter updates, starting with a higher value of $p$ and ending with a lower value, treated as tunable hyperparameters.
+Figure 11: Cosine $p$ schedule, decreasing from 10 to 1
 Initializing $p>1$ focuses learning on smoother likelihoods during the early stages of the cycle, where global features emerge between adjacent power posteriors.
@@
 Increasing this improves exploration but also increases cost due to the need for more MALA iterations.
-Typically, this is limited and therefore guided by hardware availability, such as $N_{t}=5$.
+Typically, this is memory-bound and guided by hardware availability, such as $N_{t}=5$.
 $N_{\text{local}}>1$ Determines the number of ULA or autoMALA iterations per power posterior.
 Although increasing this value improves the quality of samples at each stage, it also raises the computational cost, much like $N_{t}$.
-It can be kept at a moderate value, such as $N_{\text{local}}=20$, (based on the recommendations of Nijkamp et al.
+It can be kept at a small value, such as $N_{\text{local}}=40$, (based on the recommendations of Nijkamp et al.
 (2019b; a) regarding short-run Markov chain Monte Carlo).
@@
 (2024).
-$p_{\text{init}}>1,$ $p_{\text{end}}\leq 1,$ $N_{\text{cycles}}$ Controls the evolution of tempering during training.
+$p_{\text{init}}>1,$ $p_{\text{end}}\leq 1,$ $N_{\text{cycles}}>0$ Controls the evolution of tempering during training.
 If the data is expected to have smooth relationships, reduce both $p_{\text{init}}$ and $p_{\text{end}}$.
 For data with finer details, increase both values to allow for better focus on capturing intricate features.
-Recommended values are $p_{\text{init}}=4,$ and $p_{\text{end}}=1$, since our experience suggests that power posteriors change rapidly closer to the prior.
+Recommended values are $p_{\text{init}}=10,$ and $p_{\text{end}}=3$, since our experience suggests that power posteriors change rapidly closer to the prior.
+The number of cycles, $N_{\text{cycles}}$, controls the periodicity of cosine scheduling similar to cyclic annealing (Fu et al., 2019).
 A.4 Derivations and other identities
@@
 spline_activation (-) Silu
-layer_norm (-) true
+layer_norm (-) false
 quadrature_method (-) gausslegendre
@@
 All other KAN hparams Identical to EBM Prior
-generator_variance $\sigma_{\epsilon}$ 1.0
+generator_noise $\sigma_{\epsilon}$ 0.1
+generator_variance $\sigma_{\text{llhood}}$ 0.1
 output_activation $\sigma_{\text{act}}$ sigmoid
@@
 update_llhood_grid (-) true
-num_grid_updating_samples (-) 50
+num_grid_updating_samples (-) 100
 grid_update_frequency (-) 100
@@
 importance_sample_size $N_{s}$ 100
-Training examples $N_{\text{train}}$ 20,000
+Training examples $N_{\text{train}}$ 50,000
 Testing examples $N_{\text{test}}$ 100
@@
 spline_activation (-) Silu
-layer_norm (-) true
+layer_norm (-) false
 quadrature_method (-) gausslegendre
@@
 All other KAN hparams Identical to EBM Prior
-generator_variance $\sigma_{\epsilon}$ 0.1
+generator_noise $\sigma_{\epsilon}$ 0.1
+generator_variance $\sigma_{\text{llhood}}$ 0.1
 output_activation $\sigma_{\text{act}}$ sigmoid
@@
 Training
-batch_size $N_{x}$ 50
-importance_sample_size $N_{s}$ 50
+batch_size $N_{x}$ 100
+importance_sample_size $N_{s}$ 100
 Training examples $N_{\text{train}}$ 1,000
@@
 num_generated_samples (-) 3,000
-Number of epochs $N_{\text{epochs}}$ 600
+Number of epochs $N_{\text{epochs}}$ 1,000
 eps $\epsilon$ 0.001
@@
 spline_type (-) Cheby
-spline_activation (-) Silu
+spline_activation (-) None
 layer_norm (-) false
 quadrature_method (-) gausslegendre
-spline_degree $d$ 3
-grid_size $N_{\text{grid}}$ 0, (Chebyshev requires no grid)
+spline_degree $d$ 5
+grid_size $N_{\text{grid}}$ 1
 grid_update_ratio $\alpha_{\text{grid}}$ N/A
-grid_range $z_{\text{min}},z_{\text{max}}] [$1.2,1.2][-1.2,1.2]
+grid_range $z_{\text{min}},z_{\text{max}}] [$1,1][-1,1]
 Spline basis $\varepsilon$ init. $\varepsilon$ 0.1
@@
 Basis spline scaling $\sigma_{s}$ 1.0
-Initial basis param. $\tau_{0}$ 1.01
+Initial basis param. $\tau_{0}$ 1.001
 Train basis param. $\tau_{\text{trainable}}$ false
-GaussQuad_nodes $N_{\text{quad}}$ 100
-Use mixture prior (-) variable
+GaussQuad_nodes $N_{\text{quad}}$ 200
+Use mixture prior (-) true
 KAN Likelihood
 cnn_widths (-) See Sec. A.5.1
-All other KAN hparams N/A
-generator_variance $\sigma_{\epsilon}$ 0.3
+All other KAN hparams N/A, an MLP CNN was used
+generator_noise $\sigma_{\epsilon}$ 0.01
+generator_variance $\sigma_{\epsilon}$ 0.25
 output_activation $\sigma_{\text{act}}$ sigmoid
@@
 ESS threshold $\gamma$ N/A
+VGG16 perceptual loss scale, perceptual_scale (-) N/A
 Grid updating
-update_prior_grid (-) false
+update_prior_grid (-) true
 update_llhood_grid (-) false
@@
 importance_sample_size $N_{s}$ N/A
-Training examples $N_{\text{train}}$ 40,000
+Training examples $N_{\text{train}}$ 20,000
 Testing examples $N_{\text{test}}$ 100
 num_generated_samples (-) 3,000
-Number of epochs $N_{\text{epochs}}$ 20
+Number of epochs $N_{\text{epochs}}$ 40
 eps $\epsilon$ 0.0001
+Use VGG16 perceptual loss, use_perceptual_loss (-) false
 Optimizer
 type (-) adam
-learning_rate $\eta_{LR}$ 0.0005
+learning_rate $\eta_{LR}$ 0.0002
 betas (-) (0.9, 0.999)
-PRIOR_LANGEVIN
-iters $N_{\text{local}}$ 60
-step_size $\eta$ 0.16
 POSTERIOR_LANGEVIN
-iters $N_{\text{local}}$ SVHN: 20
-CIFAR10: 40
-step_size $\eta$ 0.01
+iters $N_{\text{local}}$ 40
+step_size $\eta$ 0.002
 THERMODYNAMIC_INTEGRATION
-iters $N_{\text{local}}$ 20
-num_temps $N_{t}$ 5
-Initial exponent in power law schedule $p_{\text{start}}$ 4
-Final exponent in power law schedule $p_{\text{end}}$ 1
-num_cycles $N_{\text{cycles}}$ 0
+iters $N_{\text{local}}$ (see posterior langevin)
+num_temps $N_{t}$ 10
+Initial exponent in power law schedule $p_{\text{start}}$ 10
+Final exponent in power law schedule $p_{\text{end}}$ 3
+Number of cosine cycles $N_{\text{cycles}}$ 0
 Global swap frequency, replica_exchange_frequency (-) 1
-A.5.1 SVHN & CIFAR-10 Architectures
-Table 7: Dense Chebyshev KAN architecture for mixture prior model
+A.5.1 SVHN & CelebA Architectures
+Table 7: The dense Chebyshev KAN network used as the mixture prior model
 Layer Operation Output Dim.
-Input Latent vector $\bm{z}$ $p=100$
-100 Dense($100$) $q=201$
+Input Sample $\bm{z}$ $p=20$
+Energy functions Dense(2020) $q=41$
 sizes=[20,41]
-Table 8: Dense Chebyshev KAN architecture for deep prior model
-Layer Operation Output Dim.
-Input Latent vector $\bm{z}$ $p=100$
-100 Dense($100$) $200$
-200 Dense($200$) $200$
-200 Dense($200$) $q=1$
-sizes=[100,200,200,1]
-Table 9: CNN Generator Architecture for SVHN
+Table 8: CNN Generator Architecture for SVHN ($32\times 32$)
+Layer Operation Output Shape Stride
+Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
+1 $\times4$ convT($256$), LeakyReLU $4\times 4\times 256$ 1
+2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
+3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
+4 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
+Config: kernel=4,4,4,4 strides=1/2/2/2, paddings=0/1/1/1, activation=LeakyReLU, batchnorm=False
+Table 9: CNN Generator Architecture for CelebA ($64\times 64$)
 Layer Operation Output Shape Stride
@@
 3 $\times4$ convT($128$), LeakyReLU $16\times 16\times 128$ 2
-4 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
-Config: kernel=4, strides=1/2/2/2, paddings=0/1/1/1, activation=LeakyReLU, batchnorm=False
-Table 10: CNN Generator Architecture for CIFAR-10
-Layer Operation Output Shape Stride
-Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
-1 $\times8$ convT($1024$), LeakyReLU $8\times 8\times 1024$ 1
-2 $\times4$ convT($512$), LeakyReLU $16\times 16\times 512$ 2
-3 $\times4$ convT($256$), LeakyReLU $32\times 32\times 256$ 2
-4 $\times3$ convT(3), Sigmoid $32\times 32\times 3$ 1
-Config: kernel=8/4/4/3, strides=1/2/2/1, paddings=0/1/1/1, activation=LeakyReLU, batchnorm=False
+4 $\times4$ convT(64), LeakyReLU $32\times 32\times 64$ 2
+5 $\times4$ convT(3), Sigmoid $64\times 64\times 3$ 2
+Config: kernel=4,4,4,4,4 strides=1/2/2/2/2, paddings=0/1/1/1/1, activation=LeakyReLU, batchnorm=False
 A.6 Sampling theory
-This section provides more detail regarding T-KAM’s sampling procedures.
+This section provides more theoretical detail regarding T-KAM’s sampling procedures.
 A.6.1 Importance Sampling
@@
 The algorithm provides theoretical guarantees of the target, due to its satisfaction of detailed balance under Eq.
-73:
+72:
 $\gamma\left(\bm{z}^{(i,\;t_{k})}\right)\,q\left(\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})}\right)$ $=\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)\,q\left(\bm{z}^{(i,\;t_{k})}\mid\bm{z}^{\prime\;(i,\;t_{k})}\right),$ (73)
@@
     Global Swaps: Global swaps are proposed and accepted subject to Eq. 34
-A.7 $20\times 20$ image grids
+A.7 Extra image samples
 A.7.1 MNIST
-Figure 13: Generated MNIST, (Deng, 2012), after 2,000 parameter updates using MLE / IS adhering to KART’s structure.
-Uniform, lognormal, and Gaussian priors are contrasted using Radial Basis Functions, (Li, 2024).
-Lognormal is inverted for clarity.
-Figure 14: Generated MNIST, (Deng, 2012), after 2,000 parameter updates using MLE / IS adhering to KART’s structure.
-Uniform, lognormal, and Gaussian priors are contrasted using Fourier Bases, (Xu et al., 2024).
-Lognormal is inverted for clarity.
+(a) Real samples
+(b) No ref. prior, RBF basis
+(c) No ref. prior, FFT basis
+(d) Gaussian ref. prior, RBF basis
+(e) Gaussian ref. prior, FFT basis
+(f) Uniform ref. prior, RBF basis
+(g) Uniform ref. prior, $F$ basis
+(h) Lognormal prior, RBF basis
+(i) Lognormal ref. prior, FFT basis
+Figure 12: Generated MNIST examples using different bases and exponentially-tilted priors (Sec. 3.1.1).
 A.7.2 FMNIST
-Figure 15: Generated FMNIST, (Xiao et al., 2017), after 2,000 parameter updates using MLE / IS adhering to KART’s structure.
-Uniform, lognormal, and Gaussian priors are contrasted using Radial Basis Functions, (Li, 2024).
-Lognormal is inverted for clarity.
-Figure 16: Generated FMNIST, (Xiao et al., 2017), after 2,000 parameter updates using MLE / IS adhering to KART’s structure.
-Uniform, lognormal, and Gaussian priors are contrasted using Fourier Bases, (Xu et al., 2024).
-Lognormal is inverted for clarity.
+(a) Real samples
+(b) No ref. prior, RBF basis
+(c) No ref. prior, FFT basis
+(d) Gaussian ref. prior, RBF basis
+(e) Gaussian ref. prior, FFT basis
+(f) Uniform ref. prior, RBF basis
+(g) Uniform ref. prior, $F$ basis
+(h) Lognormal prior, RBF basis
+(i) Lognormal ref. prior, FFT basis
+Figure 13: Generated FMNIST examples using different bases and exponentially-tilted priors (Sec. 3.1.1).
 A.7.3 Darcy flow
-Figure 17: Generated 2D Darcy flow pressures, (Li et al., 2021), after 12,000 parameter updates using MLE / IS while adhering to KART.
-Radial Basis Functions, (RBFs) (Li, 2024), are contrasted against Fourier bases, (FFTs) (Xu et al., 2024), using Gaussian latent priors.
-RBF is colored differently for clarity.
-Figure 18: Generated 2D Darcy flow pressures, (Li et al., 2021), after 12,000 parameter updates using MLE / IS while adhering to KART.
-Radial Basis Functions, (RBFs) (Li, 2024), are contrasted against Fourier bases, (FFTs) (Xu et al., 2024), using lognormal latent priors.
-RBF is colored differently for clarity.
-Figure 19: Generated 2D Darcy flow pressures, (Li et al., 2021), after 12,000 parameter updates using MLE / IS while adhering to KART.
-Radial Basis Functions, (RBFs) (Li, 2024), are contrasted against Fourier bases, (FFTs) (Xu et al., 2024), using uniform latent priors.
-RBF is colored differently for clarity.
+(a) Real samples
+(b) No ref. prior, RBF basis
+(c) No ref. prior, FFT basis
+(d) Gaussian ref. prior, RBF basis
+(e) Gaussian ref. prior, FFT basis
+(f) Uniform ref. prior, RBF basis
+(g) Uniform ref. prior, $F$ basis
+(h) Lognormal prior, RBF basis
+(i) Lognormal ref. prior, FFT basis
+Figure 14: Generated Darcy flow examples using different bases and exponentially-tilted priors (Sec. 3.1.2).
 A.7.4 SVHN
-(a) SVHN with mixture prior and MLE
-(b) SVHN with mixture prior and SE
-Figure 20: Generated samples from T-KAM trained on SVHN using a mixture Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ITS prior and ULA posterior sampling.
-(a) SVHN with deep prior and MLE
-(b) SVHN with deep prior and SE
-Figure 21: Generated samples from T-KAM trained on SVHN using a deep Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ULA prior and ULA posterior sampling.
-A.7.5 CIFAR-10
-(a) CIFAR-10 with mixture prior and MLE
-(b) CIFAR-10 with mixture prior and SE
-Figure 22: Generated samples from T-KAM trained on CIFAR-10 using a mixture Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ITS prior and ULA posterior sampling.
-(a) CIFAR-10 with deep prior and MLE
-(b) CIFAR-10 with deep prior and SE
-Figure 23: Samples from T-KAM trained on CIFAR-10 using a deep Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ULA prior and ULA posterior sampling.
-A.8 Top 3 extracted priors
+(a) Maximum Likelihood
+(b) Steppingstone estimator
+Figure 15: Generated SVHN examples by T-KAM trained with MLE and SE.
+A.7.5 CelebA
+(a) Maximum Likelihood
+(b) Steppingstone estimator
+Figure 16: Generated CelebA examples by T-KAM trained with MLE and SE.
+A.8 Trained latent priors
 The following priors were extracted from trained T-KAM models, each corresponding to the first three components of prior $q=1$.
@@
 (c) $q=1,p=3$
-Figure 24: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with RBF bases and Gaussian constrained priors.
+Figure 17: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 25: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with RBF bases and lognormal constrained priors.
+Figure 18: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 26: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with RBF bases and uniform constrained priors.
+Figure 19: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
 A.8.2 MNIST FFT
@@
 (c) $q=1,p=3$
-Figure 27: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with FFT bases and Gaussian constrained priors.
+Figure 20: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 28: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with FFT bases and lognormal constrained priors.
+Figure 21: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 29: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with FFT bases and uniform constrained priors.
+Figure 22: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
 A.8.3 FMNIST RBF
@@
 (c) $q=1,p=3$
-Figure 30: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with RBF bases and Gaussian constrained priors.
+Figure 23: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 31: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with RBF bases and lognormal constrained priors.
+Figure 24: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 32: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with RBF bases and uniform constrained priors.
+Figure 25: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
 A.8.4 FMNIST FFT
@@
 (c) $q=1,p=3$
-Figure 33: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with FFT bases and Gaussian constrained priors.
+Figure 26: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 34: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with FFT bases and lognormal constrained priors.
+Figure 27: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 35: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with FFT bases and uniform constrained priors.
+Figure 28: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
 A.8.5 Darcy RBF
@@
 (c) $q=1,p=3$
-Figure 36: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with RBF bases and Gaussian constrained priors.
+Figure 29: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 37: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with RBF bases and lognormal constrained priors.
+Figure 30: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 38: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with RBF bases and uniform constrained priors.
+Figure 31: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and uniform constrained priors.
 A.8.6 Darcy FFT
@@
 (c) $q=1,p=3$
-Figure 39: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with FFT bases and Gaussian constrained priors.
+Figure 32: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 40: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with FFT bases and lognormal constrained priors.
+Figure 33: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 41: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with FFT bases and uniform constrained priors.
+Figure 34: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and uniform constrained priors.

### v5

@@
 (2020), which replace the sigmoidal activations used in multilayer perceptrons Murtagh (1991) with spline interpolations.
-By representing the latent space with univariate functions and distributions, we enable fast and exact prior sampling with the inverse transform method (Devroye, 1986) and improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning.
+By representing the latent space with univariate functions and distributions, we enable fast, (single forward pass), and exact prior sampling with the inverse transform method (Devroye, 1986) and improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning.
 This allows the EBM prior to be aligned with the posterior before training begins.
@@
 This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and recasting ULA with population-based MCMC methods (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996).
-The resulting training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al.
+The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al.
 (2019) to ensure the impact of annealing on learning dynamics and parameter updates.
-Although this reduces expressivity compared to hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, annealing is embarrassingly parallel.
+Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, annealing is embarrassingly parallel.
 While ULA is computationally efficient and enables more flexible chain dynamics, it does not preserve detailed balance or reversibility.
@@
 2.1 are Eqs.
-46.
+57.
 2.1 Univariate energy-based prior
@@
 Here, $\delta_{\psi_{q,p}(u_{p})}$ denotes the Dirac measure centred at $\psi_{q,p}(u_{p})$, ensuring the kernel is deterministic: i.e., once $u_{p}$ is sampled, the outcome in $\mathcal{Z}$ is fixed.
-This guarantees that the pushforward from $([0,1],\mathcal{B}([0,1]),\mu)$ to $(\mathcal{Z},\mathcal{B}(\mathcal{Z}),\pi)$ is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$.
+This ensures the pushforward,
+$\left([0,1],\mathcal{B}\left([0,1]\right),\mu\right)\;\to\;\left(\mathcal{Z},\mathcal{B}\left(\mathcal{Z}\right),\pi\right),$ (3)
+is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$.
 Here, $\mathcal{B}(\mathcal{Z})$ is the Borel $\sigma$-algebra on the latent space, $\mathcal{Z}$, and $\pi$ is the latent measure.
@@
 $\pi_{q,p}(z)=\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),$ (5)
-where $f_{q,p}(z)$ is a learned energy function and $Z_{q,p}$ is the normalizing constant (partition function) that ensures $\pi_{q,p}$ defines a valid probability.
+where $f_{q,p}(z)$ is a learned energy function and $Z_{q,p}$ is the finite normalizing constant (partition function) that ensures $\pi_{q,p}$ defines a valid probability.
 The inspiration for this exponential tilting form can be traced to the energy-based prior introduced by Pang et al.
@@
 For $\pi_{q,p}$ to define a valid probability distribution, it must satisfy countable additivity while being non-negative and normalized.
-This is guaranteed since $f_{q,p}:\mathcal{Z}\to\mathbb{R}$ is Borel-measurable and integrable with respect to the reference density $\pi_{0}$, ensuring the finiteness of the partition function:
-$Z_{q,p}=\int_{\mathcal{Z}}\exp(f_{q,p}(z))\,d\pi_{0}(z)<\infty.$ (5)
+Assuming $f_{q,p}:\mathcal{Z}\to\mathbb{R}$ is $\mathcal{B}(\mathcal{Z})$-measurable with respect to the reference density $\pi_{0}$, this requires exponential integrability, i.e.,
+$\exp(f_{q,p}(z))$ $\in L^{1}\left(\pi_{0}\right)$
+$\text{such that}\quad Z_{q,p}$ $=\int_{\mathcal{Z}}\exp(f_{q,p}(z))\,d\pi_{0}(z)<\infty.$ (6)
 In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024).
@@
 To address this, particularly for more complex generative tasks such as those in Sec.
-3.2, we relax the strict adherence to KART by reformulating the prior as a mixture distribution.
+3.2, we relax the strict adherence to KART by reformulating the prior as a mixture distribution, similar to the approach described by Rivera (2024).
 This involves moving the summation over $p$ out of the generator and absorbing it into the prior, yielding a mixture model:
 $\sum_{p=1}^{n_{z}}\pi_{q,p}(z)$ $=\sum_{p=1}^{n_{z}}\alpha_{q,p}\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),\quad\text{where}\quad\sum_{p=1}^{n_{z}}\alpha_{q,p}=1,\quad\alpha_{q,p}\geq 0.$ (13)
-Here, $\alpha_{q,p}$ denote mixture weights, practically enforced via a softmax function.
+Here, $\alpha_{q,p}$ represent the mixture proportions, which must satisfy non-negativity and unit-sum constraints.
+In this study, they were fixed uniformly as $1/n_{z}$, guided by Rivera (2024).
+They could alternatively be learned by passing trainable parameters through a softmax function, (to enforce the constraints), however this approach was not used for the experiments in Sec.
+3.
 The resulting distribution defines a mixture of energy-based priors.
@@
 2.3.3 Deep KAN prior
-If required, the energy functions $f_{q,p}$ can be deepened using the KAN approach.
+If required, the energy functions $f_{q,p}$ can be deepened using the KAN approach instead of the mixture approach in Sec.
+2.3.2.
 In fact, each $f_{q,p}$ can itself be a realization of Eq.
 1 when its domain is restricted to $[0,1]$.
-This extension introduces new challenges:
+This would introduce new challenges:
   1. 1.
@@
 Tractable normalization is available through Eq.
-6, however it may be more efficient to discard normalization in the learning gradient.
+7, however it may be more efficient to discard normalization in the learning gradient.
 In Sec.
@@
 In Eqs.
-8 &  19, $z_{q,p}$ is drawn from $\pi_{q,p}$ with inverse transform sampling, (ITS) (Devroye, 1986).
+9 &  20, $z_{q,p}$ is drawn from $\pi_{q,p}$ with inverse transform sampling, (ITS) (Devroye, 1986).
 ITS can be used to sample from a component of the latent distribution, $\pi_{q,p}$, made possible given the unique characteristics of T-KAM.
@@
 The log-marginal likelihood and its learning gradient presented in Eqs.
-15 &  16, involve expectations with respect to the posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is potentially multimodal and complex.
+16 &  17, involve expectations with respect to the posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is potentially multimodal and complex.
 We propose Importance Sampling (IS) (Goertzel, 1949) as an efficient method for estimating its expectation when prior task knowledge can be incorporated into the model to improve prior-posterior alignment and the quality of the IS estimator.
@@
 This is used to estimate the log-marginal likelihood and its gradient in Eqs.
-15 and 16.
+16 and 17.
 However, unlike ULA (discussed in Sec.
@@
 Given the form of Eq.
-29, setting $t=0$ results in the power posterior distribution assuming the form of the prior distribution.
+30, setting $t=0$ results in the power posterior distribution assuming the form of the prior distribution.
 As $t$ increments, the posterior gradually adopts a shape more closely resembling the true Bayesian posterior measure.
@@
 This establishes a connection between the two, which we use to derive the learning gradient and inform our approach to the clustering the temperatures in Eq.
-30:
+31:
 […]
@@
 3.1, we demonstrate the effects of introducing different inductive biases through the choice of reference distributions in Eq.
-4, as well as through the KAN basis functions used for $f_{q,p}$ and $\Phi_{o,q}$ with comparisons between Radial Basis Functions (RBFs) (Li, 2024) and Fourier bases (FFTs) (Xu et al., 2024).
+5, as well as through the KAN basis functions used for $f_{q,p}$ and $\Phi_{o,q}$ with comparisons between Radial Basis Functions (RBFs) (Li, 2024) and Fourier bases (FFTs) (Xu et al., 2024).
 We show that, when appropriate biases are encoded, importance sampling becomes a viable and highly efficient training strategy.
@@
 The MNIST (Deng, 2012) and FMNIST (Xiao et al., 2017) datasets are grayscale datasets that can be modeled with the first likelihood model from Eq.
-20.
+21.
 The first collection of benchmarks presented in Sec.
@@
 T-KAM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=100,N_{s}=100$, and the first likelihood model in Eq.
-20.
+21.
 Training spanned 1,000 epochs with 1,000 training examples, resulting in 10,000 parameter updates.
@@
 Due to limited computational resources during this study, (detailed in Sec.
-A.2), training was shortened to 8,000 parameter updates, (70,000 were used in the study by Pang et al.
+A.2), training was shortened to 4,000 parameter updates, (70,000 were used in the study by Pang et al.
 (2020)), and smaller versions of the CNN decoders used by Pang et al.
 (2020) were adopted as the generator architectures.
-Additionally, a single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator networks.
-Chebyshev KAN bases (SS et al., 2024) were used for the prior energy functions, and full architectural details are provided under Sec.
+A single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator.
+FFT bases were used for the prior energy functions, and full architectural details are provided under Sec.
 A.5.1.
-Posterior sampling was conducted with ULA, and the standard normal reference prior was used, $\pi_{0}=\mathcal{N}\left(\bm{z};\;\bm{0},\bm{1}\right)$.
-T-KAM was trained for 40 epochs with a batch size of $N_{x}=100$ on 20,000 training examples.
-All remaining hyperparameters are provided in Tab.
+Posterior sampling was conducted with 20 steps of ULA, and no reference prior was used, $\pi_{0}=1$.
+T-KAM was trained for 20 epochs with a batch size of $N_{x}=100$ on 20,000 training examples.
+All remaining hyperparameters are in Tab.
 6.
-Annealing was conducted with $N_{t}=10$ power posteriors and the cyclic temperature scheduling outlined in Sec.
+Annealing was conducted with $N_{t}=10$ power posteriors and the $p$ scheduling outlined in Sec.
 A.3.1.
@@
 MLE SE
-$\overline{\text{FID}}_{\infty}=97.90$ $\overline{\text{FID}}_{\infty}=145.64$
-$\overline{\text{KID}}_{\infty}=0.0827$ $\overline{\text{KID}}_{\infty}=0.1407$
+$\overline{\text{FID}}_{\infty}=120.53$ $\overline{\text{FID}}_{\infty}=167.79$
+$\overline{\text{KID}}_{\infty}=0.1103$ $\overline{\text{KID}}_{\infty}=0.1644$
 Table 2: Effectively unbiased FID and KID scores (lower is better) for CelebA ($64\times 64$).
 MLE SE
-$\overline{\text{FID}}_{\infty}=226.88$ $\overline{\text{FID}}_{\infty}=209.30$
-$\overline{\text{KID}}_{\infty}=0.2590$ $\overline{\text{KID}}_{\infty}=0.2138$
+$\overline{\text{FID}}_{\infty}=174.69$ $\overline{\text{FID}}_{\infty}=173.78$
+$\overline{\text{KID}}_{\infty}=0.1917$ $\overline{\text{KID}}_{\infty}=0.1851$
 (a) Real samples
@@
 (2020), which relied on ULA to sample from the prior, thus introducing bias and inefficiency.
-SE only resulted in better image quality for CelebA and cannot be justified in this setting given its computational overhead.
-Alternative approaches, such as different annealing schedules or the use of autoMALA with bias correction and temperature-adaptive step size tuning, could improve performance, but they are not worth investigating until the hardware of Zetta (2024) becomes available, due to the added complexity, hyperparameter tuning, and extended training time.
+Our proposed SE training strategy resulted in slightly better image metrics for CelebA, but significantly worsened them with SVHN.
+However, from our subjective perspective, the generated images appeared more well-defined, yet less diverse under SE than under MLE.
+Ultimately, we cannot justify the Thermodynamic training strategy due to its substantial computational overhead.
+With more available VRAM than in this study, it would be possible to parallelize a greater number of temperatures than $N_{t}=10$, which could potentially improve quality, as suggested by the results of Calderhead & Girolami (2009).
+More posterior sampling options, such as modified annealing schedules, a larger number of ULA steps, or the use of autoMALA with bias correction and temperature-adaptive step size tuning may further improve results.
+Nevertheless, exploring these options is not worthwhile until the hardware described by Zetta (2024) becomes available, given the added algorithmic complexity, hyperparameter sensitivity, and extended training times involved.
 4 Conclusions
@@
 The search operations in Eqs.
-1323, and 69 were also difficult to parallelize on GPUs, which are generally ill-suited for search algorithms, as they often rely on conditional logic.
+1424, and 70 were also difficult to parallelize on GPUs, which are generally ill-suited for search algorithms, as they often rely on conditional logic.
 Parallelizing conditionals can lead to thread divergence, making GPUs inefficient for these operations.
@@
 To address prior–posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins.
-Additionally, efforts should be directed toward selecting an appropriate initial domain or adapting the domain dynamically during training with IS.
+Additionally, efforts should be directed toward selecting appropriate initial domains or adapting the domains dynamically.
 4.1.3 Expressivity
@@
 The index search operations in Eqs.
-2313, & 69 were parallelized across 20 threads on a CPU (i9-13900H @ 5.40GHz).
+2414, & 70 were parallelized across 20 threads on a CPU (i9-13900H @ 5.40GHz).
 Figure 7: Benchmarks for varying latent dimensions, $n_{z}$, while sampling from independent EBM priors with ITS and estimating posterior expectations with IS.
@@
 They reflect the performance of a reverse autodifferentiation of the MLE criterion, Eq.
-16.
+17.
 Figure 8: Benchmarks for varying amounts of temperatures, $N_{t}$, using a mixture of EBM priors (Tab.
@@
 They reflect the computational cost of performing a single reverse autodifferentiation of the SE criterion, Eq.
-31.
+32.
 A.3 Guidance on tempering
@@
 A.6.1, which influences the Monte Carlo estimator in Eq.
-32, given that SE is an application of importance sampling, (Annis et al., 2019).
+33, given that SE is an application of importance sampling, (Annis et al., 2019).
 This overlap is quantified by the KL divergence term in Eq.
-51.
+52.
 Maintaining small KL divergences between adjacent power posteriors can be achieved by clustering temperatures in regions where $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$ changes rapidly.
 To accomplish this, we follow Calderhead & Girolami (2009) in adopting a power-law relationship to schedule the temperatures of Eq.
-30:
+31:
 $t_{k}=\left(\frac{k}{N_{t}}\right)^{p},\quad k=0,1,\ldots,N_{t},$ (37)
@@
 However, their findings may not hold in the context of T-KAM and generative modeling, making $p$ and the temperature schedule difficult variables to tune.
-Instead, we schedule $p$ sinusoidally across all training parameter updates, starting with a higher value of $p$ and ending with a lower value, treated as tunable hyperparameters.
-Figure 11: Cosine $p$ schedule, decreasing from 10 to 1
-Initializing $p>1$ focuses learning on smoother likelihoods during the early stages of the cycle, where global features emerge between adjacent power posteriors.
-As training progresses, T-KAM will have learned the global features sufficiently well, reducing the change of measure between adjacent power posteriors of a $p>1$ schedule.
-At this stage, $p(i)$ will have transitioned to $p\leq 1$, thereby shifting T-KAM’s focus towards capturing more detailed features of the posterior landscape, where training is expected to enhance its ability to generate finer details.
+Instead, we schedule $p$ sinusoidally across all training parameter updates, starting with a lower value of $p$ and ending with a higher value, treated as tunable hyperparameters.
+Figure 11: Cosine $p$ schedule, increasing from 1 to 8
+Initializing $p\leq 1$ focuses learning on high-variance, data-informed likelihoods during the early stages of the cycle.
+Here, features emerge closer to the true posterior, and learning requires exploration and higher variance gradient estimates.
+As training progresses, T-KAM will have learned global features sufficiently well, reducing the meaningful change of measure between adjacent power posteriors of a $p\leq 1$ schedule.
+At this stage, $p(i)$ will have transitioned to $p>1$, thereby shifting focus towards exploitation and sculpting the prior with lower variance gradient estimates.
 Guidance on selecting optimal values for the tempering hyperparameters is provided in Tab. 3.
@@
 (2024).
-$p_{\text{init}}>1,$ $p_{\text{end}}\leq 1,$ $N_{\text{cycles}}>0$ Controls the evolution of tempering during training.
+$p_{\text{init}}\leq 1,$ $p_{\text{end}}>1,$ $N_{\text{cycles}}\geq 0$ Controls the evolution of tempering during training.
 If the data is expected to have smooth relationships, reduce both $p_{\text{init}}$ and $p_{\text{end}}$.
 For data with finer details, increase both values to allow for better focus on capturing intricate features.
-Recommended values are $p_{\text{init}}=10,$ and $p_{\text{end}}=3$, since our experience suggests that power posteriors change rapidly closer to the prior.
+The values used for this study were $p_{\text{init}}=1$, and $p_{\text{end}}=8$, since our experience suggests that power posteriors change rapidly closer to the prior.
 The number of cycles, $N_{\text{cycles}}$, controls the periodicity of cosine scheduling similar to cyclic annealing (Fu et al., 2019).
@@
 The prior model is solely dependent on $\bm{f}$ in Eq.
-4.
+5.
 We can rewrite Eq.
@@
 The learning gradient derived from the discretized Thermodynamic Integral is unchanged from the MLE gradient in Eq.
-16, as demonstrated under Sec.
+17, as demonstrated under Sec.
 A.4.7.
@@
 grid_update_ratio $\alpha_{\text{grid}}$ 0.05
-grid_range $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
+Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
 Lognormal: $[0,4]$
@@
 grid_update_ratio $\alpha_{\text{grid}}$ 0.05
-grid_range $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
+Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
 Lognormal: $[0,4]$
@@
 quadrature_method (-) gausslegendre
-spline_degree $d$ 5
+spline_degree $d$ 3
 grid_size $N_{\text{grid}}$ 1
 grid_update_ratio $\alpha_{\text{grid}}$ N/A
-grid_range $z_{\text{min}},z_{\text{max}}] [$1,1][-1,1]
+Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}] [$1.5,1.5][-1.5,1.5]
 Spline basis $\varepsilon$ init. $\varepsilon$ 0.1
@@
 Train basis param. $\tau_{\text{trainable}}$ false
-GaussQuad_nodes $N_{\text{quad}}$ 200
+GaussQuad_nodes $N_{\text{quad}}$ 400
 Use mixture prior (-) true
@@
 generator_noise $\sigma_{\epsilon}$ 0.01
-generator_variance $\sigma_{\epsilon}$ 0.25
+generator_variance $\sigma_{\epsilon}$ 0.3
 output_activation $\sigma_{\text{act}}$ sigmoid
@@
 num_generated_samples (-) 3,000
-Number of epochs $N_{\text{epochs}}$ 40
+Number of epochs $N_{\text{epochs}}$ 20
 eps $\epsilon$ 0.0001
@@
 type (-) adam
-learning_rate $\eta_{LR}$ 0.0002
+learning_rate $\eta_{LR}$ 0.0001
 betas (-) (0.9, 0.999)
 POSTERIOR_LANGEVIN
-iters $N_{\text{local}}$ 40
-step_size $\eta$ 0.002
+iters $N_{\text{local}}$ 20
+step_size $\eta$ 0.01
 THERMODYNAMIC_INTEGRATION
@@
 num_temps $N_{t}$ 10
-Initial exponent in power law schedule $p_{\text{start}}$ 10
-Final exponent in power law schedule $p_{\text{end}}$ 3
+Initial exponent in power law schedule $p_{\text{start}}$ 1
+Final exponent in power law schedule $p_{\text{end}}$ 8
 Number of cosine cycles $N_{\text{cycles}}$ 0
@@
 Layer Operation Output Dim.
-Input Sample $\bm{z}$ $p=20$
-Energy functions Dense(2020) $q=41$
+Input Sample $\bm{z}$ $p=40$
+Energy functions Dense($40$) $q=81$
 sizes=[40,81]
@@
 In T-KAM, there is likely to be significant mismatch due to the complexity of the likelihood models described in Eq.
-20.
+21.
 Consequently, the weights will exhibit high variance, resulting in many samples contributing minimally.
@@
 (2005), whenever the ESS falls below a certain threshold determined by Eq.
-26.
+27.
 Resampling redistributes the sample weights, creating a new population with uniformly distributed weights while preserving the statistical properties of the original distribution.
@@
 This estimator is used for the log-marginal likelihood and its gradient in Eqs.
-15 &  16, since it provides a means of estimating expectations with respect to the posterior distribution.
+16 &  17, since it provides a means of estimating expectations with respect to the posterior distribution.
 A.6.2 Residual resampling
@@
 The algorithm provides theoretical guarantees of the target, due to its satisfaction of detailed balance under Eq.
-72:
+73:
 $\gamma\left(\bm{z}^{(i,\;t_{k})}\right)\,q\left(\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})}\right)$ $=\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)\,q\left(\bm{z}^{(i,\;t_{k})}\mid\bm{z}^{\prime\;(i,\;t_{k})}\right),$ (74)

### v6

@@
-Structured Generative Modeling with The Thermodynamic Kolmogorov-Arnold Model
+Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling
 Abstract
-Learning an energy-based model (EBM) in the latent space of a top-down generative model offers a versatile framework for generation across multiple data modalities.
+Learning an energy-based model (EBM) in the latent space of a top-down generative model offers a powerful framework for generation across many data modalities.
 However, it remains unclear how its interpretability can be used to guide model design, improve generative quality, and reduce training time.
 Moreover, the reliance on Langevin Monte Carlo (LMC) sampling presents challenges in efficiency and sampling multimodal latent distributions.
-In this work, we propose a novel adaptation of the Kolmogorov-Arnold representation theorem for generative modeling and introduce the Thermodynamic Kolmogorov-Arnold Model (T-KAM) to take advantage of structural and inductive biases.
-By constraining the prior to univariate relationships, T-KAM enables fast and exact inference via the inverse transform method.
+We propose a novel adaptation of the Kolmogorov-Arnold representation theorem for generative modeling and introduce the Kolmogorov-Arnold Energy Model (KAEM) to take advantage of structural and inductive biases.
+By constraining the prior to univariate relationships, KAEM enables fast and exact inference via the inverse transform method.
 With the low dimensionality of the latent space and suitable inductive biases encoded, we demonstrate that importance sampling (IS) becomes a viable, unbiased, and highly efficient posterior sampler.
-For situations where IS fails, we investigate a novel strategy using population-based LMC, which decomposes posterior sampling into a sequence of annealed distributions to improve multimodal sampling.
-T-KAM elegantly balances common trade-offs in generative modeling, offering fast inference, interpretability, and stable training, while being naturally suited to upcoming Zettascale Computing Corp.
-hardware.
+For domains where IS fails, we introduce a strategy based on population-based LMC, decomposing the posterior into a sequence of annealed distributions to improve LMC mixing.
+KAEM balances common generative modeling trade-offs, offering fast inference, interpretability, and stable training, while being naturally suited to Zettascale Computing hardware.
 1 Introduction
@@
 (2024), as RBFs are more efficient for GPU implementation, and Poggio & Girosi (1990) previously demonstrated that they provide an optimal solution to the regularized approximation/interpolation problem.
-To derive our new model, which we call the Thermodynamic Kolmogorov-Arnold Model (T-KAM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces.
+To derive our new model, which we call the Kolmogorov-Arnold Energy Model (KAEM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces.
 Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on intuition and automated tuning.
@@
 This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and recasting ULA with population-based MCMC methods (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996).
-The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al.
+can The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al.
 (2019) to ensure the impact of annealing on learning dynamics and parameter updates.
-Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, annealing is embarrassingly parallel.
+Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, which scales sequentially, annealing can scale by adding more temperature in parallel.
 While ULA is computationally efficient and enables more flexible chain dynamics, it does not preserve detailed balance or reversibility.
@@
 Despite our optimizations, the element-wise and localized nature of KANs complicate their parallelization on GPUs, which are best suited to matrix arithmetic.
-As a result, forward passes through T-KAM and the gradient evaluations required for ULA were found to be slow, necessitating reduced experimental complexity.
+As a result, forward passes through KAEM and the gradient evaluations required for ULA were found to be slow, necessitating reduced experimental complexity.
 The experiments in Sec.
@@
 However, this result may change as a larger number of power posteriors, longer training durations, and more advanced posterior sampling algorithms become practical with the hardware advances proposed by Zetta (2024).
-Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting T-KAM at scale, while improving the potential of univariate priors with its function fusion features.
-We present T-KAM as an initial step to inspire further adaptations of KART in machine learning.
+Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting KAEM at scale, while improving the potential of univariate priors with its function fusion features.
+We present KAEM as an initial step to inspire further adaptations of KART in machine learning.
 This work aims to lay the groundwork for more exploration amongst the research community, with the far-reaching hope of eventually advancing towards a broader conceptualization: “The Kolmogorov-Arnold Representation Theorem Is All You Need.“
-2 The Thermodynamic Kolmogorov-Arnold Model
+2 The Kolmogorov-Arnold Energy Model
 This section introduces the model and its associated sampling procedures.
@@
 2.5 to improve multimodal sampling during training.
-The Kolmogorov-Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of T-KAM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
+The Kolmogorov-Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of KAEM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
 The Kolmgorov-Arnold Representation theorem 1
@@
 Eq.
-1 can provide structural bias and guide the design of T-KAM’s architecture.
+1 can provide structural bias and guide the design of KAEM’s architecture.
 To ensure adherence to it, any transformations must be derived to be continuous, measurable, and deterministic.
@@
 To ensure that the transformation $u_{p}\mapsto\psi_{q,p}(u_{p})$ is well defined in a probabilistic setting, we require it to be a Markov kernel.
-For all Borel measurable sets $A\subseteq\mathcal{Z}$, where $\mathcal{Z}\subset\mathbb{R}$ is introduced as the latent space of T-KAM, the map
+For all Borel measurable sets $A\subseteq\mathcal{Z}$, where $\mathcal{Z}\subset\mathbb{R}$ is introduced as the latent space of KAEM, the map
 $u_{p}\ \mapsto\ K(u_{p},A)=\delta_{\psi_{q,p}(u_{p})}(A)$ (2)
@@
 The smoothness of $G$, together with its bounded support, makes this approximation highly accurate.
-With an valid form for the normalized CDF, inverse transform sampling (ITS) can be used to draw samples from $\pi_{q,p}$ as deterministic transformations of uniform noise.
+With a valid form for the normalized CDF, inverse transform sampling (ITS) can be used to draw samples from $\pi_{q,p}$ as deterministic transformations of uniform noise.
 This preserves the deterministic structure of KART while enabling flexible, learnable latent distributions.
@@
 The functions, $\Phi_{o,q}:\mathbb{R}\to\mathbb{R}$, are realizable as basis or KAN functions similar to $f_{q,p}:\mathcal{Z}\to\mathbb{R}$.
-2.3 Scaling T-KAM
+2.3 Scaling KAEM
 The trainable components of the prior and inner functions are shared across the features of each input $\bm{x}^{(b)}$, while the outer generator functions remain separate for each output dimension.
@@
 Deepening this generator, as guided by Liu et al.
-(2024) for the Kolmogorov-Arnold Network (KAN), does not replace KART as the innermost foundation of T-KAM:
+(2024) for the Kolmogorov-Arnold Network (KAN), does not replace KART as the innermost foundation of KAEM:
 $x_{o}^{(s)}=$ $\sum_{l_{n}=1}^{L_{n}}h_{o,l_{n}}\Bigg(\dots h_{l_{3},l_{2}}\Bigg(\sum_{l_{1}=1}^{L_{1}}h_{l_{2},l_{1}}\Bigg(\sum_{q=1}^{2n_{z}+1}\Phi_{l_{1},q}\Bigg(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\Bigg)\Bigg)\Bigg)\Bigg)+\epsilon_{o}^{(s)}.$ (11)
@@
 $\min_{\bm{f},\bm{\Phi}}\mathcal{L}(\bm{x}^{(1)},\dots,\bm{x}^{(N_{x})})=\min_{\bm{f},\bm{\Phi}}\left[-\sum_{b=1}^{N_{x}}\log\left(P(\bm{x}^{(b)}\mid\bm{f},\mathbf{\Phi})\right)\right].$ (16)
-The conditional dependency on the collection, $\bm{f},\mathbf{\Phi}$, represents all learnable elements of T-KAM.
+The conditional dependency on the collection, $\bm{f},\mathbf{\Phi}$, represents all learnable elements of KAEM.
 For each sample of $\bm{x}^{(b)}$, the log-marginal likelihood in Eq.
@@
 In Sec.
-A.4.4, we derive the contrastive divergence learning gradient for T-KAM, which faster to train and more typical of energy-based modeling:
+A.4.4, we derive the contrastive divergence learning gradient for KAEM, which faster to train and more typical of energy-based modeling:
 $\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\left[\log P\left(\bm{z}\mid\bm{f}\right)\right]\Bigg]=\;$ $\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]$
@@
 9 &  20, $z_{q,p}$ is drawn from $\pi_{q,p}$ with inverse transform sampling, (ITS) (Devroye, 1986).
-ITS can be used to sample from a component of the latent distribution, $\pi_{q,p}$, made possible given the unique characteristics of T-KAM.
+ITS can be used to sample from a component of the latent distribution, $\pi_{q,p}$, made possible given the unique characteristics of KAEM.
 Specifically, the univariate, finite-grid structure of the constituent functions enables quadrature/interpolating integral approximations.
@@
 3 Experiments
-Our experiments are designed to highlight the unique advantages of T-KAM and to compare the proposed strategies for scaling and training.
+Our experiments are designed to highlight the unique advantages of KAEM and to compare the proposed strategies for scaling and training.
 In Sec.
@@
 (2020).
-We additionally compare T-KAM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion.
-These comparisons assess the practicality of using mixture distributions for scaling T-KAM and using annealing to improve training for high-dimensional RGB images.
+We additionally compare KAEM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion.
+These comparisons assess the practicality of using mixture distributions for scaling KAEM and using annealing to improve training for high-dimensional RGB images.
 To quantify the fidelity of generated data in a more reliable manner than previous studies in the literature, we used effectively unbiased metrics based on the principles outlined by Chong & Forsyth (2020).
@@
 4.
-T-KAM was trained for 10 epochs with 50,000 training examples, an importance sample size of $N_{s}=100$, and a batch size of $N_{x}=100$, (corresponding to 5,000 parameter updates).
+KAEM was trained for 10 epochs with 50,000 training examples, an importance sample size of $N_{s}=100$, and a batch size of $N_{x}=100$, (corresponding to 5,000 parameter updates).
 Prior sampling was conducted with ITS and posterior expectations were estimated with IS.
@@
 Figure 3: Generated FMNIST examples using different bases and exponentially-tilted priors
-T-KAM generated a diverse set of samples when a non-restrictive reference prior was used, (Gaussian or no reference), thus validating importance sampling and independent energy-based priors for simple datasets.
+KAEM generated a diverse set of samples when a non-restrictive reference prior was used, (Gaussian or no reference), thus validating importance sampling and independent energy-based priors for simple datasets.
 Subjectively, the best MNIST samples arose with no reference prior and RBF bases, whereas the best FMNIST samples came from a Gaussian reference prior with RBF bases, potentially because FMNIST has smoother features than MNIST.
@@
 In Sec.
-A.8, examples of trained univariate latent distributions have been plotted to demonstrate T-KAM’s interpretability.
-Learning can be understood by how T-KAM reshaped its latent representation relative to the initial reference prior.
+A.8, examples of trained univariate latent distributions have been plotted to demonstrate KAEM’s interpretability.
+Learning can be understood by how KAEM reshaped its latent representation relative to the initial reference prior.
 For instance, lognormal priors showed little change, indicating posterior collapse, and FFTs caused much sharper relationships than the smoother RBFs.
@@
 One approach is to use input permeability samples while employing the Fourier Neural Operator (FNO) (Li et al., 2021), to learn the solution operator mapping between permeability and flow pressure.
-However, in this study, we use the purely generative task to illustrate how weak inductive biases can influence what T-KAM learns.
-T-KAM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=100,N_{s}=100$, and the first likelihood model in Eq.
+However, in this study, we use the purely generative task to illustrate how weak inductive biases can influence what KAEM learns.
+KAEM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=100,N_{s}=100$, and the first likelihood model in Eq.
 21.
@@
 3.2 Langevin Monte Carlo and Steppingstone estimator
-T-KAM, when adherent to KART, can be interpreted as a dense or spline-based network.
+KAEM, when adherent to KART, can be interpreted as a dense or spline-based network.
 However, for spatial data, convolutional neural networks (CNNs) (LeCun et al., 2015), are more suitable as the generator, $\mathcal{G}$.
 In addition to their computational efficiency through weight sharing, CNNs offer a key advantage over dense networks due to their spatial inductive bias: sequences of convolutional layers implement shift-equivariant transformations that are well-suited for images.
-We evaluated T-KAM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015).
+We evaluated KAEM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015).
 The first likelihood model in Eq.
@@
 Due to limited computational resources during this study, (detailed in Sec.
-A.2), training was shortened to 4,000 parameter updates, (70,000 were used in the study by Pang et al.
-(2020)), and smaller versions of the CNN decoders used by Pang et al.
+A.2), training was shortened to 8,000 parameter updates, (70,000 were used in the study by Pang et al.
+(2020)), and significantly smaller versions of the CNN decoders used by Pang et al.
 (2020) were adopted as the generator architectures.
@@
 A.5.1.
-Posterior sampling was conducted with 20 steps of ULA, and no reference prior was used, $\pi_{0}=1$.
-T-KAM was trained for 20 epochs with a batch size of $N_{x}=100$ on 20,000 training examples.
+Posterior sampling was conducted with 40 steps of ULA, and no reference prior was used, $\pi_{0}=1$.
+KAEM was trained for 40 epochs with a batch size of $N_{x}=100$ on 20,000 training examples.
 All remaining hyperparameters are in Tab.
 6.
-Annealing was conducted with $N_{t}=10$ power posteriors and the $p$ scheduling outlined in Sec.
+Annealing was conducted with $N_{t}=5$ power posteriors and the $p$ scheduling outlined in Sec.
 A.3.1.
@@
 MLE SE
-$\overline{\text{FID}}_{\infty}=120.53$ $\overline{\text{FID}}_{\infty}=167.79$
-$\overline{\text{KID}}_{\infty}=0.1103$ $\overline{\text{KID}}_{\infty}=0.1644$
+$\overline{\text{FID}}_{\infty}=108.40$ $\overline{\text{FID}}_{\infty}=131.85$
+$\overline{\text{KID}}_{\infty}=0.0955$ $\overline{\text{KID}}_{\infty}=0.1188$
 Table 2: Effectively unbiased FID and KID scores (lower is better) for CelebA ($64\times 64$).
 MLE SE
-$\overline{\text{FID}}_{\infty}=174.69$ $\overline{\text{FID}}_{\infty}=173.78$
-$\overline{\text{KID}}_{\infty}=0.1917$ $\overline{\text{KID}}_{\infty}=0.1851$
+$\overline{\text{FID}}_{\infty}=122.18$ $\overline{\text{FID}}_{\infty}=127.54$
+$\overline{\text{KID}}_{\infty}=0.1349$ $\overline{\text{KID}}_{\infty}=0.1420$
 (a) Real samples
@@
 (2020), which relied on ULA to sample from the prior, thus introducing bias and inefficiency.
-Our proposed SE training strategy resulted in slightly better image metrics for CelebA, but significantly worsened them with SVHN.
+Our proposed SE training strategy resulted in worse image metrics for both datasets.
 However, from our subjective perspective, the generated images appeared more well-defined, yet less diverse under SE than under MLE.
 Ultimately, we cannot justify the Thermodynamic training strategy due to its substantial computational overhead.
-With more available VRAM than in this study, it would be possible to parallelize a greater number of temperatures than $N_{t}=10$, which could potentially improve quality, as suggested by the results of Calderhead & Girolami (2009).
+With more VRAM than available in this study, it would be possible to parallelize a greater number of temperatures than $N_{t}=5$, which could potentially improve quality, as suggested by the results of Calderhead & Girolami (2009).
 More posterior sampling options, such as modified annealing schedules, a larger number of ULA steps, or the use of autoMALA with bias correction and temperature-adaptive step size tuning may further improve results.
@@
 4 Conclusions
-T-KAM provides a principled and scalable framework for structured and informed generative modeling.
+KAEM provides a principled and scalable framework for structured and informed generative modeling.
 Our results demonstrate that correctly chosen basis functions and reference priors can enable more efficient training on simple or well-understood datasets through importance sampling.
-We also show how the Kolmogorov-Arnold representation theorem can inform architectural design and introduce structural bias, allowing T-KAM’s architecture to be guided entirely by the size of its latent space.
+We also show how the Kolmogorov-Arnold representation theorem can inform architectural design and introduce structural bias, allowing KAEM’s architecture to be guided entirely by the size of its latent space.
 Moreover, we find that the inverse transform method offers clear advantages over Langevin Monte Carlo sampling for energy-based priors in terms of speed and exact sampling.
@@
 Parallelizing conditionals can lead to thread divergence, making GPUs inefficient for these operations.
-Additionally, while ITS enables fast inference and scales when the prior is a mixture model, T-KAM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA.
+Additionally, while ITS enables fast inference and scales when the prior is a mixture model, KAEM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA.
 GPUs are not well suited to Langevin Monte Carlo methods due to the irregular memory access patterns and bandwidth pressure caused by differentiating with respect to inputs, which requires storing the outputs of each layer within the generator.
-Nonetheless, the unique advantages of T-KAM justify continued research and development.
+Nonetheless, the unique advantages of KAEM justify continued research and development.
 Moreover, Zetta (2024) introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control that can mitigate the challenges of sparsity, recursion, and irregular memory access patterns.
@@
 Even without annealing, replacing ULA with autoMALA may yield a more robust probabilistic model with stronger theoretical guarantees.
-Overall, T-KAM may prove more scalable than conventional generative models, owing to its reliance on finite representations of continuous functions and its compatibility with the LFU.
+Overall, KAEM may prove more scalable than conventional generative models, owing to its reliance on finite representations of continuous functions and its compatibility with the LFU.
 4.1.2 More opportunities for importance sampling
@@
 3.2.
-This may be attributed to prior–posterior mismatch, which rendered the IS estimator ineffective.
+This may be attributed to prior-posterior mismatch, which rendered the IS estimator ineffective.
 Moreover, relying on IS restricts the prior to its initial domain, since proposals are generated from the prior using ITS and finite-domain KAN energy functions.
@@
 (2014), who also applied PCA to train VAEs.
-To address prior–posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins.
+To address prior-posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins.
 Additionally, efforts should be directed toward selecting appropriate initial domains or adapting the domains dynamically.
@@
 For example, Normalizing Flows (Papamakarios et al., 2021) achieve this by applying a sequence of reversible transformations that progressively map a simple base distribution into a complex one.
-This might be achievable in T-KAM by introducing more latent distributions, or revising Eq.
+This might be achievable in KAEM by introducing more latent distributions, or revising Eq.
 13 into a Mixture of Experts framework (Cai et al., 2024) where $\alpha_{q,p}$ could be realized as a more flexible kernel to enable different inductive biases.
@@
 (2020), states that sampling is expensive due to their use of MCMC methods.
-T-KAM could potentially circumvent the additional cost by using IS as a fast sampler, though this would require addressing its current limitations.
+KAEM could potentially circumvent the additional cost by using IS as a fast sampler, though this would require addressing its current limitations.
 Appendix A Appendix
 A.1 Codebase
-The study was implemented at https://github.com/thezettascale/T-KAM.
+The study was implemented at https://github.com/thezettascale/KAEM.
 A.2 Computational benchmarks
@@
 30 is used to evaluate the plausibility of model parameters in relation to the data, rather than reflecting a latent variable posterior.
-However, their findings may not hold in the context of T-KAM and generative modeling, making $p$ and the temperature schedule difficult variables to tune.
+However, their findings may not hold in the context of KAEM and generative modeling, making $p$ and the temperature schedule difficult variables to tune.
 Instead, we schedule $p$ sinusoidally across all training parameter updates, starting with a lower value of $p$ and ending with a higher value, treated as tunable hyperparameters.
@@
 Here, features emerge closer to the true posterior, and learning requires exploration and higher variance gradient estimates.
-As training progresses, T-KAM will have learned global features sufficiently well, reducing the meaningful change of measure between adjacent power posteriors of a $p\leq 1$ schedule.
+As training progresses, KAEM will have learned global features sufficiently well, reducing the meaningful change of measure between adjacent power posteriors of a $p\leq 1$ schedule.
 At this stage, $p(i)$ will have transitioned to $p>1$, thereby shifting focus towards exploitation and sculpting the prior with lower variance gradient estimates.
@@
 $\therefore$ $\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})+\nabla_{\bm{f}}\log P(\bm{z}\mid\bm{f})\Big]$ (59)
-Telescoping the sum in Eq. 58, and noting that $k=1$ corresponds to the power posterior equaling the prior, and $k=N_{t}$ to the posterior, Eq. 58 yields T-KAM’s learning gradient:
+Telescoping the sum in Eq. 58, and noting that $k=1$ corresponds to the power posterior equaling the prior, and $k=N_{t}$ to the posterior, Eq. 58 yields KAEM’s learning gradient:
 $\nabla_{\bm{f},\bm{\Phi}}\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)$ $=\sum_{k=1}^{N_{t}}\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k-1}})$
@@
 layer_widths $p,q$ See Sec. A.5.1
-spline_type (-) Cheby
-spline_activation (-) None
+spline_type (-) FFT
+spline_activation (-) LeakyReLU
 layer_norm (-) false
@@
 Basis spline scaling $\sigma_{s}$ 1.0
-Initial basis param. $\tau_{0}$ 1.001
-Train basis param. $\tau_{\text{trainable}}$ false
-GaussQuad_nodes $N_{\text{quad}}$ 400
+Initial basis param. $\tau_{0}$ 1
+Train basis param. $\tau_{\text{trainable}}$ true
+GaussQuad_nodes $N_{\text{quad}}$ 200
 Use mixture prior (-) true
@@
 update_llhood_grid (-) false
+num_grid_updating_samples (-) 100
+grid_update_frequency (-) 100
+grid_update_decay (-) 0.9999
 Training
@@
 num_generated_samples (-) 3,000
-Number of epochs $N_{\text{epochs}}$ 20
+Number of epochs $N_{\text{epochs}}$ 40
 eps $\epsilon$ 0.0001
@@
 type (-) adam
-learning_rate $\eta_{LR}$ 0.0001
+learning_rate $\eta_{LR}$ 0.0002
 betas (-) (0.9, 0.999)
 POSTERIOR_LANGEVIN
-iters $N_{\text{local}}$ 20
-step_size $\eta$ 0.01
+iters $N_{\text{local}}$ 40
+step_size $\eta$ 0.005
 THERMODYNAMIC_INTEGRATION
 iters $N_{\text{local}}$ (see posterior langevin)
-num_temps $N_{t}$ 10
+num_temps $N_{t}$ 5
 Initial exponent in power law schedule $p_{\text{start}}$ 1
@@
 Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
-1 $\times4$ convT($256$), LeakyReLU $4\times 4\times 256$ 1
-2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
-3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
+1 $\times4$ convT($128$), LeakyReLU $4\times 4\times 128$ 1
+2 $\times4$ convT($64$), LeakyReLU $8\times 8\times 64$ 2
+3 $\times4$ convT($32$), LeakyReLU $16\times 16\times 32$ 2
 4 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
@@
 Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
-1 $\times4$ convT($512$), LeakyReLU $4\times 4\times 512$ 1
-2 $\times4$ convT($256$), LeakyReLU $8\times 8\times 256$ 2
-3 $\times4$ convT($128$), LeakyReLU $16\times 16\times 128$ 2
-4 $\times4$ convT(64), LeakyReLU $32\times 32\times 64$ 2
+1 $\times4$ convT($256$), LeakyReLU $4\times 4\times 256$ 1
+2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
+3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
+4 $\times4$ convT(32), LeakyReLU $32\times 32\times 32$ 2
 5 $\times4$ convT(3), Sigmoid $64\times 64\times 3$ 2
@@
 A.6 Sampling theory
-This section provides more theoretical detail regarding T-KAM’s sampling procedures.
+This section provides more theoretical detail regarding KAEM’s sampling procedures.
 A.6.1 Importance Sampling
@@
 $Q(A)=0$ implies $P(A)=0$, for every measurable subset $A\in\mathcal{Z}$.
-In the context of T-KAM, the target distribution is the posterior, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is proportional to the product of the likelihood and the prior:
+In the context of KAEM, the target distribution is the posterior, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is proportional to the product of the likelihood and the prior:
 $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})\propto P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\cdot P(\bm{z}\mid\bm{f}).$ (62)
@@
 The suitability of these weights depends on how well $Q(\bm{z})$ matches the posterior.
-In T-KAM, there is likely to be significant mismatch due to the complexity of the likelihood models described in Eq.
+In KAEM, there is likely to be significant mismatch due to the complexity of the likelihood models described in Eq.
 21.
@@
 (b) Steppingstone estimator
-Figure 15: Generated SVHN examples by T-KAM trained with MLE and SE.
+Figure 15: Generated SVHN examples by KAEM trained with MLE and SE.
 A.7.5 CelebA
@@
 (b) Steppingstone estimator
-Figure 16: Generated CelebA examples by T-KAM trained with MLE and SE.
+Figure 16: Generated CelebA examples by KAEM trained with MLE and SE.
 A.8 Trained latent priors
-The following priors were extracted from trained T-KAM models, each corresponding to the first three components of prior $q=1$.
+The following priors were extracted from trained KAEM models, each corresponding to the first three components of prior $q=1$.
 A.8.1 MNIST RBF
@@
 (c) $q=1,p=3$
-Figure 17: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
+Figure 17: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 18: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
+Figure 18: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 19: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
+Figure 19: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
 A.8.2 MNIST FFT
@@
 (c) $q=1,p=3$
-Figure 20: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
+Figure 20: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 21: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
+Figure 21: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 22: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
+Figure 22: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
 A.8.3 FMNIST RBF
@@
 (c) $q=1,p=3$
-Figure 23: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
+Figure 23: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 24: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
+Figure 24: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 25: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
+Figure 25: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
 A.8.4 FMNIST FFT
@@
 (c) $q=1,p=3$
-Figure 26: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
+Figure 26: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 27: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
+Figure 27: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 28: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
+Figure 28: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
 A.8.5 Darcy RBF
@@
 (c) $q=1,p=3$
-Figure 29: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and Gaussian constrained priors.
+Figure 29: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 30: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and lognormal constrained priors.
+Figure 30: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 31: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and uniform constrained priors.
+Figure 31: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and uniform constrained priors.
 A.8.6 Darcy FFT
@@
 (c) $q=1,p=3$
-Figure 32: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and Gaussian constrained priors.
+Figure 32: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and Gaussian constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 33: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and lognormal constrained priors.
+Figure 33: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and lognormal constrained priors.
 (a) $q=1,p=1$
@@
 (c) $q=1,p=3$
-Figure 34: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and uniform constrained priors.
+Figure 34: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and uniform constrained priors.

### v7

@@
 Deep generative modeling is a core subfield of machine learning that enables systems to learn complex data distributions, generate novel data, and support a wide range of tasks including simulation, compression, and representation learning.
-Two prominent and widely studied approaches are Variational Autoencoders (VAEs) (Kingma & Welling, 2022) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014).
+Two prominent and widely studied approaches are Variational Autoencoders (VAEs) (kingma2022autoencodingvariationalbayes) and Generative Adversarial Networks (GANs) (goodfellow2014generativeadversarialnetworks).
 Both frameworks assume that each data point is generated from a low-dimensional latent vector via a top-down neural network, where the latent vector is typically drawn from a simple, non-informative prior, such as a uniform or isotropic Gaussian distribution.
-Latent energy-based prior models (EBMs) (Pang et al., 2020) extend this framework by learning an expressive, data-dependent prior directly in the latent space.
+Latent energy-based prior models (EBMs) (pang2020learninglatentspaceenergybased) extend this framework by learning an expressive, data-dependent prior directly in the latent space.
 This prior can be trained jointly with the top-down generator using maximum likelihood estimation (MLE), where the EBM learns by the statistical discrepancy between samples from the prior and posterior distributions, following a paradigm similar to Empirical Bayes.
-Each learning iteration requires Markov Chain Monte Carlo (MCMC) sampling from both the prior and posterior, specifically the Unadjusted Langevin Algorithm (ULA) (Brooks et al., 2011; Rossky et al., 1978b; Roberts & Stramer, 2002), which offers greater modeling flexibility compared to the amortized inference methods of VAEs.
-Despite being extendable to several high-impact research directions, such as multimodal generative modeling (Yuan et al., 2024), and interpretable classification and meta-learning (Kong et al., 2021), (which involves rapid adaptability and few-shot generalization to out-of-distribution tasks), latent EBMs face critical limitations due to their reliance on MCMC sampling.
+Each learning iteration requires Markov Chain Monte Carlo (MCMC) sampling from both the prior and posterior, specifically the Unadjusted Langevin Algorithm (ULA) (Brooks2011; Rossky1978; Roberts2002), which often offers greater modeling flexibility compared to the amortized inference methods of VAEs.
+Despite being extendable to several high-impact research directions, such as multimodal generative modeling (yuan2024learningmultimodallatentgenerative), and interpretable classification, text-to-text, and meta-learning (pang2021latentspaceenergybasedmodeltext), (which involves rapid adaptability and few-shot generalization to out-of-distribution tasks), latent EBMs face critical limitations due to their reliance on MCMC sampling.
 Firstly, MCMC is inherently iterative, and algorithms such as ULA require repeated gradient passes through the model, which introduces significant computational overhead during both training and inference.
 Secondly, ULA is best suited for sampling from smooth, unimodal distributions and can struggle to explore the multimodal latent distributions that are typical in latent variable generative modeling.
-Moreover, the framework presented by Pang et al.
-(2020) requires separate ULA samplers for the prior and posterior.
-This introduces additional experimental complexity since ULA is sensitive to hyperparameter tuning: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias due to its use of the Euler-Maruyama method (Wilkie, 2004) to evolve the underlying Langevin equation.
-Additionally, it is not evident how the unique separation between the prior and generator networks can be leveraged for interpretability or training efficiency.
-Unlike the framework introduced by Karniadakis et al.
-(2021), which describes how scientific inductive and learning biases can be embedded into neural networks to reduce their reliance on data and improve training efficiency, generalization, and explainability, (all of which are critical for ensuring trust in Artificial Intelligence), it remains unclear how the EBM prior can be effectively aligned with the posterior to achieve similar effects.
-Although ideas from mechanistic interpretability, as reviewed by Bereska & Gavves (2024), could potentially be applied, there is currently no established method to leverage the unique strengths of the EBM prior framework for this purpose.
+Moreover, the framework presented by pang2020learninglatentspaceenergybased requires separate ULA samplers for the prior and posterior.
+This introduces additional experimental complexity since ULA is sensitive to hyperparameter tuning: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias due to its use of the Euler-Maruyama method (Wilkie_2004) to evolve the underlying Langevin equation.
+Additionally, it is unclear how the unique separation between the prior and generator networks can be leveraged for interpretability or training efficiency.
+Unlike the framework introduced by Karniadakis2021, which describes how scientific inductive and learning biases can be embedded into neural networks to reduce their reliance on data and improve training efficiency, generalization, and explainability, (all of which are critical for ensuring trust in Artificial Intelligence), it remains unclear how the EBM prior can be effectively aligned with the posterior to achieve similar effects.
+Although ideas from mechanistic interpretability, as reviewed by bereska2024mechanisticinterpretabilityaisafety, could potentially be applied, there is currently no established method to leverage the unique strengths of the EBM prior framework for this purpose.
 1.1 Inductive biases and fast sampling
-To address these limitations, we reimagine the EBM prior into a novel univariate or element-wise model, drawing inspiration from Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024) and deep spline networks Bohra et al.
-(2020), which replace the sigmoidal activations used in multilayer perceptrons Murtagh (1991) with spline interpolations.
-By representing the latent space with univariate functions and distributions, we enable fast, (single forward pass), and exact prior sampling with the inverse transform method (Devroye, 1986) and improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning.
-This allows the EBM prior to be aligned with the posterior before training begins.
+To address these limitations, we reimagine the EBM prior into a novel univariate or element-wise model, drawing inspiration from Kolmogorov-Arnold Networks (KANs) (liu2024kankolmogorovarnoldnetworks) and deep spline networks 9264754, which replace the sigmoidal activations used in multilayer perceptrons MURTAGH1991183 with spline interpolations.
+By representing the latent space with univariate functions and distributions, we enable fast, (single forward pass), and exact prior sampling with the inverse transform method (devroye1986nonuniform) and improve interpretability for domain experts who may not be deeply familiar with machine learning.
+This allows the EBM prior to be aligned with the posterior before training begins, and for latent distributions to be plotted, viewed, and potentially understood with the symbolic regression approach proposed by liu2024kankolmogorovarnoldnetworks.
 In Sec.
-3.1, we show that adequately chosen priors can enable efficient training with importance sampling (IS) (Goertzel, 1949).
-This is a technique for estimating posterior expectations that is otherwise avoided in the literature due to its high variance under severe prior-posterior mismatch and the curse of dimensionality, (Dupuis & Wang, 2004), which can be avoided in our model due to the low dimensionality of the latent space.
-Moreover, we show that the trained priors are easy to visualize and recover, offering new opportunities for scientific discovery and knowledge transfer to models with different latent structures.
-Multiple KAN bases and reference priors were used to verify the effect of inductive bias on training with IS.
-However, Gaussian Radial Basis Functions (RBF KANs) (Li, 2024), were adopted instead of the cubic B-spline bases proposed by Liu et al.
-(2024), as RBFs are more efficient for GPU implementation, and Poggio & Girosi (1990) previously demonstrated that they provide an optimal solution to the regularized approximation/interpolation problem.
-To derive our new model, which we call the Kolmogorov-Arnold Energy Model (KAEM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces.
+3.1, we show that simple datasets can allow for efficient training with importance sampling (IS) (goertzel1949quota).
+This is a technique for estimating posterior expectations that is otherwise avoided in the literature due to its high variance under severe prior-posterior mismatch and the curse of dimensionality, (dupuis2004importance), which can be avoided in our model due to the low dimensionality of the latent space.
+Moreover, we show that the trained priors are easy to visualize and recover, offering new opportunities for latent discovery and knowledge transfer to models with different latent structures.
+To derive our new model, which we call the Kolmogorov-Arnold Energy Model (KAEM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Arnold2009) by reinterpreting its inner functions as Markov kernels between probability spaces.
 Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on intuition and automated tuning.
@@
 1.2 Population-based Langevin sampling and Thermodynamic training
-However, while ITS enables fast inference, posterior sampling must still be conducted using ULA when the IS estimator is insufficient.
-Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by Cui & Han (2024) and the diffusion-assisted EBM by Zhang et al.
-(2023).
+However, while ITS enables fast inference, posterior sampling must still be conducted using ULA when the IS estimator is rendered ineffective due to variance.
+Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by cui2024learninglatentspacehierarchical and the diffusion-assisted EBM by zhang2023persistentlytraineddiffusionassistedenergybased.
 To maintain prior interpretability and fast inference post-training, we instead introduce posterior annealing as a novel and alternative strategy.
-This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and recasting ULA with population-based MCMC methods (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996).
-can The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al.
-(2019) to ensure the impact of annealing on learning dynamics and parameter updates.
-Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, which scales sequentially, annealing can scale by adding more temperature in parallel.
+This involves decomposing the posterior into a sequence of power posteriors, presented by 2d186535-310b-3428-8c6f-e4ef298483d7, and recasting ULA with population-based MCMC methods (PT; Marinari1992; Hukushima1996).
+The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by CALDERHEAD20094028; Annis2019 to ensure the impact of annealing on learning dynamics and parameter updates.
+Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, which scales sequentially, annealing can scale by adding more annealed distributions in parallel.
 While ULA is computationally efficient and enables more flexible chain dynamics, it does not preserve detailed balance or reversibility.
-This can be recovered by incorporating Metropolis-Hastings (MH) adjustments (Metropolis & Ulam, 1949) via the Metropolis-adjusted Langevin Algorithm (MALA) (Rossky et al., 1978a).
-The code in Sec.
-A.1 optionally supports MALA with locally adaptive step-size tuning, (autoMALA) (Biron-Lattes et al., 2024), to assist when a single step size fails to accommodate the varying geometries of power posteriors, and when the MH criterion deteriorates under the curse of dimensionality, (Hairer et al., 2014).
-This is discussed in Secs.
-A.6.3 and A.6.4.
-However, ULA remains the preferred method in our work, motivated by its efficiency and effectiveness, demonstrated by Pang et al.
-(2020) and further supported by the findings of Nijkamp et al.
-(2019b; a), with regards to how short-run, biased MCMC methods can still be highly effective for sampling from EBMs.
+This can be recovered by incorporating Metropolis-Hastings (MH) adjustments (13ab5b5e-0237-33fb-a7a8-6f6e4e0d4e0f) via the Metropolis-adjusted Langevin Algorithm (MALA) (10.1111/j.1467-9868.2010.00765.x).
+However, ULA remains the preferred method in our work, motivated by its efficiency and effectiveness, demonstrated by pang2020learninglatentspaceenergybased and further supported by the findings of nijkamp2019learningnonconvergentnonpersistentshortrun; nijkamp2019anatomymcmcbasedmaximumlikelihood, with regards to how short-run, biased MCMC methods can still be highly effective for sampling from EBMs.
 MLE and SE are compared in Sec.
@@
 As a result, forward passes through KAEM and the gradient evaluations required for ULA were found to be slow, necessitating reduced experimental complexity.
+1.3 Scope of experiments
+KAEM was compiled and differentiated using Julia’s Reactant (Reactantjl2025) and Enzyme (moses2020enzyme; moses2021enzyme_gpu; moses2022enzyme_parallel) packages, which optimizes and lowers Julia code to Multi-Level Intermediate Representation (MLIR) (lattner2020mlircompilerinfrastructureend), thus offering best-in-class performance.
+However, Gaussian Radial Basis Functions (RBF KANs) (li2024kolmogorovarnoldnetworksradialbasis), were adopted instead of the cubic B-spline bases proposed by liu2024kankolmogorovarnoldnetworks, as RBFs are Reactant-compatible and more efficient for GPU implementation.
+They have also been proven to provide an optimal solution to the regularized approximation/interpolation problem by 58326.
 The experiments in Sec.
 3.2 demonstrate that our mixture prior model is able to generate coherent, (although low-quality), samples after a few training epochs on RGB images, while still supporting efficient, exact sampling via component-wise ITS.
-In contrast, annealing incurred substantial computational overhead, but produced little to no improvements in sample quality, rendering its use difficult to justify.
-However, this result may change as a larger number of power posteriors, longer training durations, and more advanced posterior sampling algorithms become practical with the hardware advances proposed by Zetta (2024).
-Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting KAEM at scale, while improving the potential of univariate priors with its function fusion features.
+In contrast, annealing incurred substantial computational overhead, but produced small improvements in sample quality, rendering its recommendation context-dependent.
+However, this may change as a larger number of power posteriors, longer training durations, and more advanced posterior sampling algorithms become practical with the hardware advances proposed by exa2024.
+Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting KAEM at scale.
+Their XPUs also provide function/layer fusion capabilities that may improve the potential of univariate priors and adherence to KART.
 We present KAEM as an initial step to inspire further adaptations of KART in machine learning.
@@
 2.5 to improve multimodal sampling during training.
-The Kolmogorov-Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of KAEM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
+The Kolmogorov-Arnold Representation theorem (KART) (Arnold2009) forms the basis of KAEM. A structured proof of the theorem has been provided by dzhenzher2022structuredproofkolmogorovssuperposition.
 The Kolmgorov-Arnold Representation theorem 1
@@
 2.1 are Eqs.
-57.
+46.
 2.1 Univariate energy-based prior
-We interpret $u_{p}\sim\mathcal{U}(u_{p};\,0,1)$ in Eq.
-1 as a uniform random variable in the probability space, $\left([0,1],\mathcal{B}([0,1]),\mu\right)$, where $\mathcal{B}\left([0,1]\right)$ is the Borel $\sigma$-algebra on $[0,1]$ and $\mu$ is the Lebesgue measure on $[0,1]$.
-To ensure that the transformation $u_{p}\mapsto\psi_{q,p}(u_{p})$ is well defined in a probabilistic setting, we require it to be a Markov kernel.
-For all Borel measurable sets $A\subseteq\mathcal{Z}$, where $\mathcal{Z}\subset\mathbb{R}$ is introduced as the latent space of KAEM, the map
-$u_{p}\ \mapsto\ K(u_{p},A)=\delta_{\psi_{q,p}(u_{p})}(A)$ (2)
-is measurable.
-Here, $\delta_{\psi_{q,p}(u_{p})}$ denotes the Dirac measure centred at $\psi_{q,p}(u_{p})$, ensuring the kernel is deterministic: i.e., once $u_{p}$ is sampled, the outcome in $\mathcal{Z}$ is fixed.
-This ensures the pushforward,
-$\left([0,1],\mathcal{B}\left([0,1]\right),\mu\right)\;\to\;\left(\mathcal{Z},\mathcal{B}\left(\mathcal{Z}\right),\pi\right),$ (3)
-is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$.
-Here, $\mathcal{B}(\mathcal{Z})$ is the Borel $\sigma$-algebra on the latent space, $\mathcal{Z}$, and $\pi$ is the latent measure.
-The functions $\psi_{q,p}$ can therefore be conveniently interpreted as inverse cumulative distribution functions (CDFs) corresponding to independent univariate priors:
-$\psi_{q,p}(u_{p})=F^{-1}(\pi_{q,p})(u_{p}).$ (4)
-Here, $F^{-1}(\pi_{q,p}):[0,1]\to\mathcal{Z}$ is the generalized inverse CDF associated with a parameterized latent distribution, $\pi_{q,p}$.
-This is valid under standard regularity assumptions on $F$, since it is a measurable, monotone, increasing function, and thus $F^{-1}(u)=\inf\left\{z\in\mathcal{Z}:F(z)\geq u\right\}$ is a measurable mapping between the two probability spaces.
-We parameterize the latent distribution $\pi_{q,p}$ via exponential tilting of a base prior $\pi_{0}$:
-$\pi_{q,p}(z)=\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),$ (5)
-where $f_{q,p}(z)$ is a learned energy function and $Z_{q,p}$ is the finite normalizing constant (partition function) that ensures $\pi_{q,p}$ defines a valid probability.
-The inspiration for this exponential tilting form can be traced to the energy-based prior introduced by Pang et al.
-(2020), adapted to the univariate case.
-The base prior $\pi_{0}$ serves as a reference measure that guides the form of the learned distribution $\pi_{q,p}$, allowing for flexible reweighting via the energy function $f_{q,p}$.
-For $\pi_{q,p}$ to define a valid probability distribution, it must satisfy countable additivity while being non-negative and normalized.
-Assuming $f_{q,p}:\mathcal{Z}\to\mathbb{R}$ is $\mathcal{B}(\mathcal{Z})$-measurable with respect to the reference density $\pi_{0}$, this requires exponential integrability, i.e.,
+The functions $\psi_{q,p}$ can be conveniently translated to the framework of inverse transform sampling (ITS) (devroye1986nonuniform).
+We interpret $\psi_{q,p}$ as the application of an inverse cumulative density function (CDF) to a uniform random variable to recover a sample from a univariate latent density, $p_{q,p}$:
+$z_{q,p}=\psi_{q,p}(u_{p})=F^{-1}_{\pi_{q,p}}(u_{p})\quad\Leftrightarrow\quad z_{q,p}\sim p_{q,p}(z).$ (2)
+Here, $F^{-1}_{\pi_{q,p}}:[0,1]\to\mathcal{Z}_{q,p}$ is the generalized inverse CDF associated with a parameterized latent measure, $\pi_{q,p}$ in the latent space, $\mathcal{Z}_{q,p}\subset\mathbb{R}$.
+We have therefore interpreted $u_{p}\sim\text{Unif}(0,1)$ in Eq.
+1 as a uniform random variable in the probability space, $\left([0,1],\mathcal{B}([0,1]),\mu\right)$, where $\mathcal{B}\left([0,1]\right)$ is the Borel $\sigma$-algebra on $[0,1]$ and $\mu$ is the uniform probability measure on $[0,1]$.
+This is valid under standard regularity assumptions on $F$, since it is a measurable, monotone, increasing function, and thus $F_{\pi_{q,p}}^{-1}(u)=\inf\left\{z\in\mathcal{Z}_{q,p}:F_{\pi_{q,p}}(z)\geq u\right\}$ is a measurable pushforward (or Markov kernel) between the two probability spaces:
+$\left([0,1],\mathcal{B}\left([0,1]\right),\mu\right)\;\underset{F_{\pi_{q,p}}^{-1}}{\longrightarrow}\;\left(\mathcal{Z}_{q,p},\mathcal{B}\left(\mathcal{Z}_{q,p}\right),\pi_{q,p}\right),$ (3)
+We parameterize the latent measure $\pi_{q,p}$ via exponential tilting of a base measure $\pi_{0}$:
+$d\pi_{q,p}(z)=\frac{\exp(f_{q,p}(z))}{Z_{q,p}}d\pi_{0}(z),$ (4)
+where $f_{q,p}(z)$ is a learned energy function and $Z_{q,p}$ is the finite normalizing constant (partition function) that ensures $\pi_{q,p}$ induces a valid probability density, $p_{q,p}$, for each latent feature.
+The inspiration for this exponential tilting form can be traced to the energy-based prior introduced by pang2020learninglatentspaceenergybased, adapted to the univariate case.
+The base prior $\pi_{0}$ serves as a reference that guides the form of the learned $\pi_{q,p}$, allowing for flexible reweighting via the energy function $f_{q,p}$.
+For $\pi_{q,p}$ to define a valid density, which we denote $p_{q,p}$, we require exponential integrability:
 $\exp(f_{q,p}(z))$ $\in L^{1}\left(\pi_{0}\right)$
-$\text{such that}\quad Z_{q,p}$ $=\int_{\mathcal{Z}}\exp(f_{q,p}(z))\,d\pi_{0}(z)<\infty.$ (6)
-In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024).
-For instance, smoothing splines, such as those used by KANs or deep spline networks, (Bohra et al., 2020).
-Splines are typically defined on bounded domains, allowing the support of $\pi_{q,p}$ to be explicitly controlled by the user.
+$\text{such that}\quad Z_{q,p}$ $=\int_{\mathcal{Z}_{q,p}}\exp(f_{q,p}(z))\,\pi_{0}(dz)<\infty.$ (5)
+In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov-Arnold Networks (KANs) (liu2024kankolmogorovarnoldnetworks).
+For instance, smoothing splines, such as those used by KANs or deep spline networks, (9264754).
+Splines are typically defined on bounded domains, allowing the support of $\pi_{q,p}$ to be explicitly controlled by the user using the KAN grids.
+These grids must be periodically updated during training according to input activations, following the implementation described by liu2024kankolmogorovarnoldnetworks.
 The partition function $Z_{q,p}$ is often intractable in energy-based models.
-However, in this univariate setting, where the energy function is smooth and defined on a bounded domain $\left[z_{\text{min}},z_{\text{max}}\right]$, it can be accurately approximated using Gauss-Kronrod quadrature, (Laurie, 1997):
+However, in this univariate setting, where the energy function is smooth and defined on a bounded domain $\left[z_{\text{min}},z_{\text{max}}\right]$, it can be accurately approximated using Gauss-Kronrod quadrature, (gauss-quad):
 $Z_{q,p}$ $=\int_{z_{\text{min}}}^{z_{\text{max}}}\exp(f_{q,p}(z))\pi_{0}(z)dz\approx\sum_{i=1}^{N_{\text{quad}}}w_{i},G(z_{i}^{\text{node}}),$ (6)
 where $G(z)=\exp(f_{q,p}(z))\pi_{0}(z)$, and $\{z_{i}^{\text{node}},w_{i}\}_{i=1}^{N_{\text{quad}}}$ are the quadrature nodes and weights for the interval $\left[z_{\text{min}},z_{\text{max}}\right]$.
-The smoothness of $G$, together with its bounded support, makes this approximation highly accurate.
-With a valid form for the normalized CDF, inverse transform sampling (ITS) can be used to draw samples from $\pi_{q,p}$ as deterministic transformations of uniform noise.
-This preserves the deterministic structure of KART while enabling flexible, learnable latent distributions.
-Inverse transform sampling remains effective even for functions that are not Riemann integrable, provided they are integrable in the Lebesgue sense, typical for the functions in KAN literature.
-As a result, the model retains a coherent probabilistic interpretation of KART that is consistent with the energy-based framework proposed by Pang et al.
-(2020), which is well suited to maximum likelihood training.
+The availability of $Z_{q,p}$ enables ITS, and the smoothness of $G$, together with its bounded support, makes this approximation accurate.
+With this valid form for the normalized CDF, inverse transform sampling (ITS) can be used to draw samples from $p_{q,p}$ as measurable transformations of uniform noise.
+This preserves the deterministic structure of KART while enabling flexible, learnable latent distributions, well-founded so long as the KAN functions admit exponential integrability.
+As a result, the model retains a coherent probabilistic interpretation of KART that is consistent with the energy-based framework proposed by pang2020learninglatentspaceenergybased, which is well suited to maximum likelihood training.
 However, this approach is significantly limited, as it ignores inter-dimensional dependencies in the latent space and results in an axis-aligned joint distribution.
@@
 2.3.2 by relaxing our adherence to KART.
+2.1.1 Multivariate prior notation
+From this point forward, we formulate our derivations over the multivariate probability space, $\bigl(\mathcal{Z},\mathcal{B}(\mathcal{Z}),\mathbb{P}_{\bm{\theta}}\bigr),$ where $\mathbb{P}_{\bm{\theta}}$ denotes a product measure parameterized by abstract model parameters, $\theta$, and the latent space is defined as a Cartesian product:
+$\mathbb{P}_{\bm{\theta}}\;:=\;\bigotimes_{q=1}^{2n_{z}+1}\;\bigotimes_{p=1}^{n_{z}}\pi_{q,p},\quad\quad\quad\mathcal{Z}\;=\;\prod_{q=1}^{2n_{z}+1}\prod_{p=1}^{n_{z}}\mathcal{Z}_{j,k}\;\subset\;\mathbb{R}^{(2n_{z}+1)\times n_{z}},$ (7)
+Each $\pi_{q,p}$ is a univariate probability measure on $\mathcal{Z}_{q,p}$ defined via exponential tilting, as introduced in Eq.
+4.
+Under $\mathbb{P}_{\bm{\theta}}$, the latent variables $\{z_{q,p}\}_{q,p}$ are mutually independent, and the full latent variable
+$\bm{z}\;=\;\bigl(z_{q,p}\bigr)_{q=1,\ldots,2n_{z}+1;\;p=1,\ldots,n_{z}}$
+is a matrix-valued random variable whose entries correspond to independent univariate features.
 2.2 Kolmogorov-Arnold Network generator
-Having sampled the intermediary latent variable with ITS such that $z_{q,p}=F^{-1}(\pi_{q,p})(u_{p})$, the generator can be realized flexibly as $\mathcal{G}:\mathcal{Z}\to\mathcal{X}$:
-$\bm{\tilde{x}}^{(s)}$ $=\mathcal{G}\left(\bm{z}^{(s)}\right)+\bm{\epsilon}^{(s)},\quad\text{where}\quad z_{q,p}^{(s)}\sim\pi_{q,p}(z),\quad\bm{\epsilon}^{(s)}\sim\mathcal{N}(\bm{\epsilon};\;\bm{0},\sigma_{\epsilon}^{2}\bm{I}),$ (8)
+Having sampled the intermediary latent variable with ITS such that $z_{q,p}=F^{-1}_{\pi_{q,p}}(u_{p})$, the generator can be realized flexibly as $\mathcal{G}:\mathcal{Z}\to\mathcal{X}$:
+$\bm{\tilde{x}}^{(s)}$ $=\mathcal{G}\left(\bm{z}^{(s)}\right)+\bm{\epsilon}^{(s)},\quad\text{where}\quad z_{q,p}^{(s)}\sim p_{q,p}(z),\quad\bm{\epsilon}^{(s)}\sim\mathcal{N}(\bm{0},\sigma_{\epsilon}^{2}\bm{I}),$ (8)
 where $\mathbf{\tilde{x}}^{(s)}=\{\tilde{x}_{o}^{(s)}\}_{o=1}^{n_{x}}$ represents a generated data sample corresponding to a single latent sample indexed by $s$, and $n_{x}$ is used to specify the dimension of $\bm{\tilde{x}}^{(s)}$.
@@
 $\tilde{x}_{o}^{(s)}$ $=\sum_{q=1}^{2n_{z}+1}\Phi_{o,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)+\epsilon_{o}^{(s)},\quad\mathbf{\tilde{x}}^{(s)}=\begin{pmatrix}\sum_{q=1}^{2n_{z}+1}\Phi_{1,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)\\[5.0pt] \sum_{q=1}^{2n_{z}+1}\Phi_{2,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)\\[5.0pt] \vdots\\[5.0pt] \sum_{q=1}^{2n_{z}+1}\Phi_{n_{x},q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)\end{pmatrix}+\bm{\epsilon}^{(s)}.$ (9)
-The functions, $\Phi_{o,q}:\mathbb{R}\to\mathbb{R}$, are realizable as basis or KAN functions similar to $f_{q,p}:\mathcal{Z}\to\mathbb{R}$.
+The functions, $\Phi_{o,q}:\mathbb{R}\to\mathbb{R}$, are realizable as basis or KAN functions similar to $f_{q,p}:\mathcal{Z}_{q,p}\to\mathbb{R}$.
 2.3 Scaling KAEM
@@
 where $n_{x}$ denotes the dimensionality of the observed data.
-From this point forward, we formulate our derivations over a multivariate probability space $(\mathcal{Z},\mathcal{B}(\mathcal{Z}),P)$, where $\mathcal{Z}$ no longer represents a subinterval of the real line, but rather a subspace of $\mathbb{R}^{(2n_{z}+1)\times n_{z}}$.
-The measure $P$ defines a product distribution over the latent variables:
-$P(\bm{z}\mid\bm{f})=\pi_{1,1}\otimes\cdots\otimes\pi_{2n_{z}+1,n_{z}},$
-where each component $\pi_{j,k}$ is defined via exponential tilting, as introduced in Sec.
-2.1.
-The full latent vector $\bm{z}$ is a matrix-valued random variable composed of the independent latent features.
 2.3.1 Generator
@@
 1 is not required.
-Deepening this generator, as guided by Liu et al.
-(2024) for the Kolmogorov-Arnold Network (KAN), does not replace KART as the innermost foundation of KAEM:
+Deepening this generator, as guided by liu2024kankolmogorovarnoldnetworks for the Kolmogorov-Arnold Network (KAN), does not replace KART as the innermost foundation of KAEM:
 $x_{o}^{(s)}=$ $\sum_{l_{n}=1}^{L_{n}}h_{o,l_{n}}\Bigg(\dots h_{l_{3},l_{2}}\Bigg(\sum_{l_{1}=1}^{L_{1}}h_{l_{2},l_{1}}\Bigg(\sum_{q=1}^{2n_{z}+1}\Phi_{l_{1},q}\Bigg(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\Bigg)\Bigg)\Bigg)\Bigg)+\epsilon_{o}^{(s)}.$ (11)
@@
 However, using sigmoid may lead to vanishing gradients and we suggest that function domains are less critical to KART than its structural properties and should not be prioritized over maintaining gradient flow and avoiding saturation.
-Adopting a KAN-based generator also allows for training with L1 regularization, followed by pruning with the method proposed by Liu et al.
-(2024).
+Adopting a KAN-based generator also allows for training with L1 regularization, followed by pruning with the method proposed by liu2024kankolmogorovarnoldnetworks.
 This enables further model compression beyond the baseline structure defined by KART.
@@
 To address this, particularly for more complex generative tasks such as those in Sec.
-3.2, we relax the strict adherence to KART by reformulating the prior as a mixture distribution, similar to the approach described by Rivera (2024).
-This involves moving the summation over $p$ out of the generator and absorbing it into the prior, yielding a mixture model:
-$\sum_{p=1}^{n_{z}}\pi_{q,p}(z)$ $=\sum_{p=1}^{n_{z}}\alpha_{q,p}\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),\quad\text{where}\quad\sum_{p=1}^{n_{z}}\alpha_{q,p}=1,\quad\alpha_{q,p}\geq 0.$ (13)
+3.2, we relax the strict adherence to KART by reformulating the prior as a mixture distribution, similar to the approach described by rivera2024trainvae.
+This involves moving the summation over p{p} from 1 to $n_{z}$ out of the generator and absorbing it into the prior, yielding a mixture density:
+$p_{\bm{f},q}(z)$ $=\sum_{p=1}^{n_{z}}\alpha_{q,p}\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),\quad\text{where}\quad\sum_{p=1}^{n_{z}}\alpha_{q,p}=1,\quad\alpha_{q,p}\geq 0.$ (13)
 Here, $\alpha_{q,p}$ represent the mixture proportions, which must satisfy non-negativity and unit-sum constraints.
-In this study, they were fixed uniformly as $1/n_{z}$, guided by Rivera (2024).
+In this study, they were fixed uniformly as $1/n_{z}$, guided by rivera2024trainvae.
 They could alternatively be learned by passing trainable parameters through a softmax function, (to enforce the constraints), however this approach was not used for the experiments in Sec.
 3.
-The resulting distribution defines a mixture of energy-based priors.
-Sampling becomes more efficient than in the previous formulation, as it proceeds component-wise using ITS for discrete variables, (Devroye, 2006).
-However, this violates adherence to KART, since $F$ now corresponds to a discrete rather than continuous CDF.
+The result is a mixture of energy-based priors.
+Sampling becomes more efficient than in the previous formulation, as it proceeds component-wise using ITS for discrete variables, (DEVROYE200683).
+However, this violates KART, since $F$ now corresponds to a discrete rather than continuous CDF.
 The mixture proportions are learned by replacing the log-prior in Eq.
@@
 Component selection follows a categorical distribution over $p=1,\dots,n_{z}$, determined by $\alpha_{q,p}$.
-Let $u_{\text{mix}}\sim\mathcal{U}(u_{\text{mix}};\;0,1)$.
-Guided by Devroye (2006) regarding how ITS applied to discrete variables, the selected component $p^{*}$ is given by:
+Let $u_{\text{mix}}\sim\text{Unif}(0,1)$.
+Guided by DEVROYE200683 regarding how ITS applied to discrete variables, the selected component $p^{*}$ is given by:
 $p^{*}=\min\left\{j\mid\sum_{p=1}^{j}\alpha_{q,p}\geq u_{\text{mix}}\right\}.$ (14)
 This ensures that each component $p^{*}$ is selected with probability $\alpha_{q,p^{*}}$. ITS can then be used to sample from the chosen component, (see  2.4.3).
-2.3.3 Deep KAN prior
-If required, the energy functions $f_{q,p}$ can be deepened using the KAN approach instead of the mixture approach in Sec.
-2.3.2.
-In fact, each $f_{q,p}$ can itself be a realization of Eq.
-1 when its domain is restricted to $[0,1]$.
-This would introduce new challenges:
-  1. 1.
-    Prior interpretability is reduced, complicating the recovery of trained priors.
-  2. 2.
-    The component-wise structure is compromised, undermining the validity of ITS and KART.
-Instead, we propose removing the summation over $p$ in Eq. 9 and reparameterizing the prior using a multilayer KAN energy function, which can be deepened as needed:
-$P\left(\bm{z}\mid\bm{f}\right)=\frac{1}{\bm{Z}}\cdot\exp\left(\sum_{q=1}^{2n_{z}+1}\Psi_{q,p}\left(\sum_{p=1}^{n_{z}}f_{q,p}(z)\right)\right)\cdot\pi_{0}(\bm{z}),$ (15)
-where $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{n_{z}}$, and is constrained to $0,1]^{n_{z}}$ when strict adherence to KART is required.
-Sampling from this prior can be carried out using the Unadjusted Langevin Algorithm (ULA), as described in Sec.
-2.5.2.
-This is not investigated in the experiments, since it is simply equivalent to the model of Pang et al.
-(2020), reformulated as a KAN.
 2.4 Maximum likelihood training
 With $N_{x}$ training examples in a batch, the optimization objective is the minimization of the negative log-marginal likelihood:
-$\min_{\bm{f},\bm{\Phi}}\mathcal{L}(\bm{x}^{(1)},\dots,\bm{x}^{(N_{x})})=\min_{\bm{f},\bm{\Phi}}\left[-\sum_{b=1}^{N_{x}}\log\left(P(\bm{x}^{(b)}\mid\bm{f},\mathbf{\Phi})\right)\right].$ (16)
-The conditional dependency on the collection, $\bm{f},\mathbf{\Phi}$, represents all learnable elements of KAEM.
+$\min_{\bm{f},\bm{\Phi}}\mathcal{L}(\bm{x}^{(1)},\dots,\bm{x}^{(N_{x})})=\min_{\bm{f},\bm{\Phi}}\left[-\sum_{b=1}^{N_{x}}\log\left(p_{\bm{f},\mathbf{\Phi}}(\bm{x}^{(b)})\right)\right].$ (15)
+The dependency on the collection, $\bm{f},\mathbf{\Phi}$, represents all learnable elements of KAEM.
 For each sample of $\bm{x}^{(b)}$, the log-marginal likelihood in Eq.
-16 can be separated out into two log-distributions by first introducing the latent vector, $\bm{z}$, otherwise absent upon marginalization over the latent distribution.
+15 can be separated out into two log-densities by first introducing the latent vector, $\bm{z}$, otherwise absent upon marginalization over the latent space.
 This is shown under Sec.
@@
 A.4.3:
-$\nabla_{\bm{f},\bm{\Phi}}\left[\log P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right]$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\left[\log P\left(\bm{z}\mid\bm{f}\right)\right]+\nabla_{\bm{\Phi}}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\Bigg],$ (17)
+$\nabla_{\bm{f},\bm{\Phi}}\left[\log p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right]$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\Bigg[\nabla_{\bm{f}}\left[\log p_{\bm{f}}\left(\bm{z}\right)\right]+\nabla_{\bm{\Phi}}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\Bigg],$ (16)
 2.4.1 Log-prior
 Given the independence of the prior’s parameters, the log-prior can be modeled as a sum over the univariate cases:
-$\log\Big(P\left(\bm{z}\mid\bm{f}\right)\Big)$ $=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(\pi_{q,p}(z_{q,p})\right)=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)$ (18)
+$\log\Big(p_{\bm{f}}\left(\bm{z}\right)\Big)$ $=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(\pi_{q,p}(z_{q,p})\right)=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)$ (17)
 Tractable normalization is available through Eq.
-7, however it may be more efficient to discard normalization in the learning gradient.
+6, however it may be more efficient to discard normalization in the learning gradient.
 In Sec.
 A.4.4, we derive the contrastive divergence learning gradient for KAEM, which faster to train and more typical of energy-based modeling:
-$\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\left[\log P\left(\bm{z}\mid\bm{f}\right)\right]\Bigg]=\;$ $\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]$
--\; $\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg].$ (19)
-Here, $\bm{z}$ is sampled from $P(\bm{z}\mid\bm{f})$ and $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$ and reflects a univariate element of the prior model.
+$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\bm{f}}\left[\log p_{\bm{f}}\left(\bm{z}\right)\right]\Bigg]=\;$ $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]-\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg].$ (18)
+Here, $\bm{z}$ is sampled from $p_{\bm{f}}(\bm{z})$ and $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$ and reflects a univariate element of the prior model.
 The expectation with respect to the posterior can be approximated using Importance Sampling (IS), outlined under Secs.
-2.4.4 &  A.6.1, or a standard Monte Carlo estimator when using Langevin sampling methods.
+2.4.4 &  A.5.1, or a standard Monte Carlo estimator when using Langevin sampling methods.
 The prior expectation may always be approximated with a Monte Carlo estimator taken across samples:
-$\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\quad\text{where}\quad z_{q,p}^{(s)}\sim\pi_{q,p}(z).$ (20)
+$\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\quad\text{where}\quad z_{q,p}^{(s)}\sim p_{q,p}(z).$ (19)
 The gradient of Eq.
@@
 For example, a Gaussian model might be adopted for images, and an autoregressive model for text.
-It can also be realized into a multimodal generative modeling framework using the principles outlined by Yuan et al.
-(2024).
-$\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)$ $=\begin{cases}-\frac{\left\lVert\bm{x}^{(b)}-\bm{\tilde{x}}\right\rVert^{2}_{2}}{2\sigma_{\text{llhood}}^{2}}+\text{const},&\text{(static continuous data)}\\[5.0pt] \sum_{o=1}^{n_{x}}\sum_{c=1}^{K}x_{o,c}^{(b)}\log\tilde{x}_{o,c}+\text{const},&\text{(static discrete data)}\\[5.0pt] \sum_{t=1}^{T}\log\left(P(\bm{x}^{(b,t)}\mid\bm{x}^{(b,1)},\dots,\bm{x}^{(b,t-1)},\bm{z},\bm{\Phi})\right),&\text{(sequential data)}\end{cases}$ (21)
+It can also be realized into a multimodal generative modeling framework using the principles outlined by yuan2024learningmultimodallatentgenerative.
+$\log\left(p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)$ $=\begin{cases}-\frac{\left\lVert\bm{x}^{(b)}-\bm{\tilde{x}}\right\rVert^{2}_{2}}{2\sigma_{\text{llhood}}^{2}}+\text{const},&\text{(static continuous data)}\\[5.0pt] \sum_{o=1}^{n_{x}}\sum_{c=1}^{K}x_{o,c}^{(b)}\log\tilde{x}_{o,c}+\text{const},&\text{(static discrete data)}\\[5.0pt] \sum_{t=1}^{T}\log\left(p_{\bm{\Phi}}(\bm{x}^{(b,t)}\mid\bm{x}^{(b,1)},\dots,\bm{x}^{(b,t-1)},\bm{z})\right),&\text{(sequential data)}\end{cases}$ (20)
 Here, $\bm{\tilde{x}}$ is a generated sample from Eq.
@@
 The code in Sec.
-A.1 also provides an option to augment the first likelihood model with an additional perceptual loss term (Johnson et al., 2016) using features extracted from the pretrained VGG-16 CNN (Simonyan & Zisserman, 2015).
+A.1 also provides an option to augment the first likelihood model with an additional perceptual loss term (johnson2016perceptuallossesrealtimestyle) using features extracted from the pretrained VGG-16 CNN (simonyan2015deepconvolutionalnetworkslargescale).
 This could have yielded higher-fidelity images for Sec.
@@
 In Eqs.
-9 &  20, $z_{q,p}$ is drawn from $\pi_{q,p}$ with inverse transform sampling, (ITS) (Devroye, 1986).
-ITS can be used to sample from a component of the latent distribution, $\pi_{q,p}$, made possible given the unique characteristics of KAEM.
-Specifically, the univariate, finite-grid structure of the constituent functions enables quadrature/interpolating integral approximations.
+9 &  19, $z_{q,p}$ is drawn from $p_{q,p}$ with inverse transform sampling, (ITS) (devroye1986nonuniform).
+ITS can be used to sample from a component of the latent density, $p_{q,p}$, made possible given the unique characteristics of KAEM.
+Specifically, the univariate, finite-grid structure of the KAN-based energy functions enables quadrature/interpolating integral approximations.
   1. 1.
-    CDF inversion: Sampling occurs using the inverse cumulative distribution function (CDF): The CDF for a single component, $F_{\pi_{q,p}}$, is:
+    CDF inversion: Sampling occurs using the inverse cumulative density function (CDF): The CDF for a single component, $F_{\pi_{q,p}}$, is:
     $F_{\pi_{q,p}}(z^{\prime})$ $=\frac{\int_{z_{\text{min}}}^{z^{\prime}}\exp(f_{q,p}(z))\pi_{0}(z)\,dz}{Z_{q,p}},\quad Z_{q,p}=\int_{z_{\text{min}}}^{z_{\text{max}}}\exp(f_{q,p}(z))\pi_{0}(z)\,dz.$ (21)
-    These two integrals can be conveniently and efficiently approximated using Gauss-Kronrod quadrature, (Laurie, 1997), given that $f_{q,p}$ is univariate and defined on a bounded grid:
+    These two integrals can be conveniently and efficiently approximated using Gauss-Kronrod quadrature, (gauss-quad), given that $f_{q,p}$ is univariate and defined on a bounded grid:
     $G(z)$ $=\exp(f_{q,p}(z))\pi_{0}(z),\quad\int_{z_{\text{min}}}^{z^{\prime}}G(z)\,dz\approx\sum_{i=1}^{N_{\text{quad}}}w_{i}\,G(z_{i}^{\text{node}}),$ (22)
     where $z_{i}^{\text{node}}$ and $w_{i}$ are the Gauss-Legendre quadrature nodes and weights for the interval $z_{\text{min}},z^{\prime}]$.
-A random variable is uniformly drawn and shared across $q$, $u_{p}\sim\mathcal{U}(u_{p};\;0,1)$.
+A random variable is uniformly drawn and shared across $q$, $u_{p}\sim\text{Unif}(0,1)$.
 The first interval where $F_{\pi_{q,p}}(z)$ is greater than $u_{p}$ scaled by $Z_{q,p}$ is identified:
@@
     $z_{p}$ $\approx z_{k^{*}-1}^{\text{node}}+\left(z_{k^{*}}^{\text{node}}-z_{k^{*}-1}^{\text{node}}\right)\cdot\frac{u_{p}-F_{\pi_{q,p}}(z_{k^{*}-1}^{\text{node}})}{F_{\pi_{q,p}}(z_{k^{*}}^{\text{node}})-F_{\pi_{q,p}}(z_{k^{*}-1}^{\text{node}})}.$ (24)
-The procedure ensures an efficient, exact and unbiased generation of samples, $z_{q,p}$, distributed according to $\pi_{q,p}$. It can be vectorized to generate samples from all prior elements.
+The procedure ensures an efficient, exact and unbiased generation of samples, $z_{q,p}$, distributed according to $p_{q,p}$. It can be vectorized to generate samples from all prior elements.
 2.4.4 Posterior Expectations with Importance Sampling
 The log-marginal likelihood and its learning gradient presented in Eqs.
-16 &  17, involve expectations with respect to the posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is potentially multimodal and complex.
-We propose Importance Sampling (IS) (Goertzel, 1949) as an efficient method for estimating its expectation when prior task knowledge can be incorporated into the model to improve prior-posterior alignment and the quality of the IS estimator.
+15 &  16, involve expectations with respect to the posterior $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$, which is potentially multimodal and complex.
+We propose Importance Sampling (IS) (goertzel1949quota) as an efficient method for estimating its expectation when datasets are simple or when task knowledge can be incorporated into the model to improve prior-posterior alignment and the quality of the IS estimator.
 IS provides a quick approximation of the posterior expectation without requiring extensive exploration.
@@
   1. 1.
-    Proposal distribution: $N_{s}$ independent samples, $\{\bm{z}^{(s)}\}_{s=1}^{N_{s}}$, are drawn from the proposal distribution, $Q(\bm{z})=P(\bm{z}\mid\bm{f})$, which the target distribution, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$ must be absolutely continuous with respect to.
-Sampling from $Q$ is tractable with the procedure outlined in Sec.
+    Proposal distribution: $N_{s}$ independent samples, $\{\bm{z}^{(s)}\}_{s=1}^{N_{s}}$, are drawn from the proposal, $q(\bm{z})=p_{\bm{f}}(\bm{z})$, which the target, $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$ must be absolutely continuous with respect to.
+Sampling from the prior is tractable with the procedure outlined in Sec.
 2.4.3.
@@
     Importance weights: The normalized importance weights are evaluated:
-    $w_{\text{norm}}(\bm{z}^{(s)})=\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big(\log(P(\bm{x}^{(b)}\mid\bm{z}^{(s)},\bm{\Phi}))\Big).$ (26)
+    $w_{\text{norm}}(\bm{z}^{(s)})=\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big(\log(p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)}))\Big).$ (25)
   3. 3.
@@
     Resample: If required, the population, $\{\bm{z}^{(s)},w_{\text{norm}}(\bm{z}^{(s)})\}_{s=1}^{N_{s}}$, is resampled from.
-There are multiple ways to accomplish this, (Douc et al., 2005).
+There are multiple ways to accomplish this, (douc2005comparisonresamplingschemesparticle).
 Systematic, stratified, and residual resampling are provided in the codebase under Sec.
-A.1, parallelized on the CPU.
+A.1, parallelized across samples.
 In this study, we adopted residual resampling, (explained under Sec.
-A.6.2).
+A.5.2).
   5. 5.
-    Expectation estimation: The posterior expectation of an arbitrary function $\rho(\bm{z})$, (e.g., the log-likelihood), is estimated with a Monte Carlo estimator:
-    $\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}[\rho(\bm{z})]\approx{\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})\cdot w_{\text{norm}}(\bm{\bar{z}_{\text{resampled}}}^{(s)})}.$ (28)
+    Expectation estimation: After resampling, IS weights are uniformly distributed and the posterior expectation of an arbitrary function or density, $\rho(\bm{z})$, (e.g., the log-likelihood), is estimated with a Monte Carlo estimator:
+    $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}[\rho(\bm{z})]$ $\approx\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)})\cdot w_{\text{norm}}(\bm{\bm{z}^{(s)})}\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})$.
+(27)
 This is used to estimate the log-marginal likelihood and its gradient in Eqs.
-16 and 17.
+15 and 16.
 However, unlike ULA (discussed in Sec.
@@
 We inform our derivations using the Thermodynamic Integral. In place of the log-marginal likelihood outlined in Eq. 15, this uses the following representation derived in Sec. A.4.5:
-$\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)\right]\,dt$ (29)
-Here, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)$ has been introduced as the power posterior, which can be considered a representation of the original posterior distribution, with a likelihood tempered by a temperature parameter, $t$.
-This is introduced by Friel & Pettitt (2008) and organized for our purposes as follows:
-$P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)$ $=\frac{P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\cdot P(\bm{z}\mid\bm{f})}{\bm{Z}_{t}},$
-$\quad\text{where}\quad\bm{Z}_{t}=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]$ $=\int_{\mathcal{Z}}P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\,\,dP(\bm{z}\mid\bm{f})$ (30)
+$\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log\left(p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)\right]\,dt$ (28)
+Here, $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)$ has been introduced as the power posterior, which can be considered a representation of the original posterior, with a likelihood tempered by a temperature parameter, $t$.
+This is introduced by 2d186535-310b-3428-8c6f-e4ef298483d7 and organized for our purposes as follows:
+$p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)$ $=\frac{p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\cdot p_{\bm{f}}(\bm{z})}{\bm{Z}_{t}},$
+$\quad\text{where}\quad\bm{Z}_{t}=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]$ $=\int_{\mathcal{Z}}p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\,\,\mathbb{P}_{\bm{f}}(d\bm{z})$ (29)
 Given the form of Eq.
-30, setting $t=0$ results in the power posterior distribution assuming the form of the prior distribution.
-As $t$ increments, the posterior gradually adopts a shape more closely resembling the true Bayesian posterior measure.
-At $t=1$, the posterior recovers its original form and converges to the original posterior distribution.
+29, setting $t=0$ results in the power posterior density assuming the form of the prior.
+As $t$ increments, the power posterior gradually adopts a shape more closely resembling the true Bayesian posterior.
+At $t=1$, the power posterior converges fully on the original posterior.
 The effect of tempering is illustrated in Fig.
@@
 In Sec.
-A.4.7, the Steppingstone estimator (SE) introduced by Annis et al.
-(2019) is derived from the discretized Thermodynamic Integral presented by Calderhead & Girolami (2009), (and provided under Sec.
+A.4.7, the Steppingstone estimator (SE) introduced by Annis2019 is derived from the discretized Thermodynamic Integral presented by CALDERHEAD20094028, (and provided under Sec.
 A.4.6).
 This establishes a connection between the two, which we use to derive the learning gradient and inform our approach to the clustering the temperatures in Eq.
-31:
+30:
 […]
 The log-partition ratios are estimated with Monte Carlo estimators using all annealed samples:
-$\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)=\log(\bm{Z}_{t_{N_{t}}})^{\text{SS}}=\sum_{k=1}^{N_{t}}\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
+$\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)=\log(\bm{Z}_{t_{N_{t}}})^{\text{SS}}=\sum_{k=1}^{N_{t}}\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
 […]
-SE and its variants, such as Annealed Importance Sampling (AIS) (Neal, 1998), are commonly used to estimate a model’s evidence without a tractable way to learn the prior, as they lack dependence on prior gradients beyond sampling.
+SE and its variants, such as Annealed Importance Sampling (AIS) (neal1998annealedimportancesampling), are commonly used to estimate a model’s evidence without a tractable way to learn the prior, as they lack dependence on prior gradients beyond sampling.
 However, by recognizing the direct connection between the SE and MLE learning gradients, as proven in Sec.
@@
 2.5.2 Unadjusted Langevin Algorithm (ULA) and Replica Exchange
-In their implementation of Thermodynamic Integration, Calderhead & Girolami (2009) estimated the Thermodynamic Integral using the discretization in Sec.
-A.4.6, using population-based MCMC and Parallel Tempering (PT) (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996) to simultaneously sample from all power posteriors.
-Specifically, parallel chains were maintained at each temperature, (or Replica), using Metropolis-Hastings (MH) (Metropolis & Ulam, 1949), to drive local moves for each power posterior.
+In their implementation of Thermodynamic Integration, CALDERHEAD20094028 estimated the Thermodynamic Integral using the discretization in Sec.
+A.4.6, using population-based MCMC and Parallel Tempering (PT) (PT; Marinari1992; Hukushima1996) to simultaneously sample from all power posteriors.
+Specifically, parallel chains were maintained at each temperature, (or Replica), using Metropolis-Hastings (MH) (13ab5b5e-0237-33fb-a7a8-6f6e4e0d4e0f), to drive local moves for each power posterior.
 They also proposed a geometric path between the prior and posterior, enabling global swaps between adjacent temperatures.
 This approach allowed higher temperatures to leverage the efficient mixing of lower temperatures, thereby facilitating a more thorough exploration of the posterior landscape.
-We adopt a similar strategy but follow Pang et al.
-(2020) in using the Unadjusted Langevin Algorithm (ULA) (Rossky et al., 1978b; Brooks et al., 2011) without MH correction.
-Specifically, we update samples using the forward Euler-Maruyama discretization of the Langevin SDE, as detailed by Roberts & Stramer (2002), with the transition kernel described in Sec.
-A.6.3:
+We adopt a similar strategy but follow pang2020learninglatentspaceenergybased in using the Unadjusted Langevin Algorithm (ULA) (Rossky1978; Brooks2011) without MH correction.
+Specifically, we update samples using the forward Euler-Maruyama discretization of the Langevin SDE, as detailed by Roberts2002:
 $\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{(i,\;t_{k})}+\eta\nabla_{\bm{z}^{(i,\;t_{k})}}\log\gamma\left(\bm{z}^{(i,\;t_{k})}\right)+\sqrt{2\eta}\bm{\epsilon}^{(i,\;t_{k})},$ (33)
 where $\eta$ is the step size, $\bm{\epsilon}^{(i,\;t_{k})}\sim\mathcal{N}(\bm{0},\bm{I})$, and $\gamma_{t_{k}}$ denotes the target power posterior:
-$\log\gamma\left(\bm{z}^{(i,\;t_{k})}\right)=\log P(\bm{z}^{(i,\;t_{k})}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\propto t_{k}\log P(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})},\bm{\Phi})+\log P(\bm{z}^{(i,\;t_{k})}\mid\bm{f}).$
+$\log\gamma\left(\bm{z}^{(i,\;t_{k})}\right)=\log p_{\bm{f},\bm{\Phi}}(\bm{z}^{(i,\;t_{k})}\mid\bm{x}^{(b)},t_{k})\propto t_{k}\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})})+\log p_{\bm{f}}(\bm{z}^{(i,\;t_{k})}).$
 ULA is best suited for smooth, unimodal distributions, but annealing with power posteriors can improve sampling from multimodal distributions.
 Mixing can be further improved by periodically swapping populations between randomly selected adjacent temperatures, subject to:
-$r=\frac{P\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k+1})},\bm{\Phi}\right)^{t_{k}}P\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})},\bm{\Phi}\right)^{t_{k+1}}}{P\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})},\bm{\Phi}\right)^{t_{k}}P\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k+1})},\bm{\Phi}\right)^{t_{k+1}}},$ (35)
+$r=\frac{p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k+1})}\right)^{t_{k}}p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})}\right)^{t_{k+1}}}{p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})}\right)^{t_{k}}p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k+1})}\right)^{t_{k+1}}},$ (34)
 where the proposed swap, $\bm{z}^{(i,\;t_{k+1})}\leftrightarrow\bm{z}^{(i,\;t_{k})}$, is accepted with probability: $\min\left(1,r\right)$.
 3 Experiments
-Our experiments are designed to highlight the unique advantages of KAEM and to compare the proposed strategies for scaling and training.
+Our experiments are designed to highlight the unique advantages of our contributions and compare the proposed strategies for scaling and training KAEM.
+Radial basis functions (RBFs) (li2024kolmogorovarnoldnetworksradialbasis) were used as the univariate functions in place of the cubic B-spline interpolations proposed by liu2024kankolmogorovarnoldnetworks.
+Benchmarks are provided in Sec.
+A.2.
 In Sec.
-3.1, we demonstrate the effects of introducing different inductive biases through the choice of reference distributions in Eq.
-5, as well as through the KAN basis functions used for $f_{q,p}$ and $\Phi_{o,q}$ with comparisons between Radial Basis Functions (RBFs) (Li, 2024) and Fourier bases (FFTs) (Xu et al., 2024).
-We show that, when appropriate biases are encoded, importance sampling becomes a viable and highly efficient training strategy.
+3.1, we show that importance sampling becomes a viable and highly efficient training strategy for simple datasets, due to the low dimensionality of the latent space.
 We also demonstrate that the model remains effective when strictly adhering to KART with independent energy-based priors, thereby validating the use of KART as a source of structural inductive bias.
 In Sec.
-3.2, we evaluate the mixture EBM prior, sampled via component-wise ITS and paired with CNN-decoder architectures, which are smaller variants of those used by Pang et al.
-(2020).
+3.2, we evaluate the mixture EBM prior, sampled via component-wise ITS and paired with CNN-decoder architectures, which are slightly smaller variants of those used by pang2020learninglatentspaceenergybased.
 We additionally compare KAEM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion.
-These comparisons assess the practicality of using mixture distributions for scaling KAEM and using annealing to improve training for high-dimensional RGB images.
-To quantify the fidelity of generated data in a more reliable manner than previous studies in the literature, we used effectively unbiased metrics based on the principles outlined by Chong & Forsyth (2020).
+These comparisons assess the practicality of using mixture densities for scaling KAEM and using annealing to improve training for high-dimensional RGB images.
+To quantify the fidelity of generated data in a more reliable manner than previous studies in the literature, we used effectively unbiased metrics based on the principles outlined by chong2020effectively.
 By applying linear regression to extrapolate the values of standard metrics to an infinitely sized sample set, Monte Carlo error was mitigated, (otherwise associated with finite sample sizes).
 For Sec.
-3.2, we gathered $\overline{\text{FID}}_{\infty}$ and $\overline{\text{KID}}_{\infty}$, which are effectively unbiased estimators of the Fréchet Inception Distance and Kernel Inception Distance, introduced by Heusel et al.
-(2018); Bińkowski et al.
-(2021) respectively.
+3.2, we gathered $\overline{\text{FID}}_{\infty}$ and $\overline{\text{KID}}_{\infty}$, which are effectively unbiased estimators of the Fréchet Inception Distance and Kernel Inception Distance, introduced by heusel2018gans; binkowski2021demystifying respectively.
 The set of sample sizes used for the linear regression were: 1000, 1200, 1400, 1600, 1800, 2000, 2200, 2400, 2600, 2800, 3000.
@@
 Furthermore, no metrics were applied in Secs.
-3.1.1 &  3.1.1, as vision metrics rely on the Inception V3 model trained on ImageNet (Szegedy et al., 2015), potentially introducing bias against other datasets, especially non-RGB images.
+3.1.1 &  3.1.1, as vision metrics rely on the Inception V3 model trained on ImageNet (szegedy2015rethinking), potentially introducing bias against other datasets, especially non-RGB images.
 3.1 Importance Sampling
 3.1.1 MNIST & FMNIST
-The MNIST (Deng, 2012) and FMNIST (Xiao et al., 2017) datasets are grayscale datasets that can be modeled with the first likelihood model from Eq.
-21.
+The MNIST (mnist) and FMNIST (fmnist) datasets are grayscale datasets that can be modeled with the first likelihood model from Eq.
+20.
 The first collection of benchmarks presented in Sec.
@@
 (a) Real samples
-(b) No ref. prior, FFT basis
-(c) Gaussian ref. prior, RBF basis
-Figure 2: Generated MNIST examples using different bases and exponentially-tilted priors
+(b) No ref. prior
+(c) Gaussian ref. prior
+(d) Uniform ref. prior
+Figure 2: Generated MNIST examples using different reference priors
 (a) Real samples
-(b) No ref. prior, RBF basis
-(c) Gaussian ref. prior, RBF basis
-Figure 3: Generated FMNIST examples using different bases and exponentially-tilted priors
-KAEM generated a diverse set of samples when a non-restrictive reference prior was used, (Gaussian or no reference), thus validating importance sampling and independent energy-based priors for simple datasets.
-Subjectively, the best MNIST samples arose with no reference prior and RBF bases, whereas the best FMNIST samples came from a Gaussian reference prior with RBF bases, potentially because FMNIST has smoother features than MNIST.
-The additional samples plotted in Sec.
-A.7.1 illustrate that the choice and combination of bases and reference priors are important.
-Poorly chosen combinations led to posterior collapse and low-quality samples, thereby showcasing the effect of inductive bias in guiding sample quality.
-In Sec.
-A.8, examples of trained univariate latent distributions have been plotted to demonstrate KAEM’s interpretability.
-Learning can be understood by how KAEM reshaped its latent representation relative to the initial reference prior.
-For instance, lognormal priors showed little change, indicating posterior collapse, and FFTs caused much sharper relationships than the smoother RBFs.
-3.1.2 2D Darcy flow pressures
-The dataset used in this experiment was sourced from Li et al.
-(2021).
-The Darcy flow equation is a second-order elliptic partial differential equation (PDE) that models various physical phenomena, including fluid flow through porous media and heat conduction.
-At steady state, its two-dimensional form on the unit square with Dirichlet boundary conditions is given by:
-$-\nabla\cdot\left(k(\bm{z}),\nabla u(\bm{z})\right)$ $=f(\bm{z}),$ $\bm{z}$ $\in[0,1]^{2},$
-$u(\bm{z})$ $=0,$ $\bm{z}$ $\in\partial[0,1]^{2},$ (36)
-where $k(\bm{z})$ is the 2D permeability field, $u(\bm{z})$ represents the 2D pressure field, and $f(\bm{z})$ is a forcing function evaluated at $\bm{z}$.
-Generating new flow-pressure samples is a challenging task in machine learning.
-One approach is to use input permeability samples while employing the Fourier Neural Operator (FNO) (Li et al., 2021), to learn the solution operator mapping between permeability and flow pressure.
-However, in this study, we use the purely generative task to illustrate how weak inductive biases can influence what KAEM learns.
-KAEM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=100,N_{s}=100$, and the first likelihood model in Eq.
-21.
-Training spanned 1,000 epochs with 1,000 training examples, resulting in 10,000 parameter updates.
-All other hyperparameters are provided in Tab.
-5.
-Prior sampling was conducted with ITS and posterior expectations were estimated with IS.
-The architecture was guided by KART and independent priors were used.
-(a) Real samples
-(b) Gaussian ref. prior, FFT basis
-(c) Uniform ref. prior, RBF basis
-Figure 4: Generated Darcy flow examples using different bases and exponentially-tilted priors
-RBFs tended to oversmooth the solution space while FFTs were more effective at capturing large-scale spatial periodicity.
-This likely stems from their suitability in representing periodic spatial relationships, a property that aligns well with the structure of the Darcy flow, as previously showcased by the success of the FNO, which also uses Fourier methods.
-This is somewhat reflected by Fig.
-32, which depicts some periodicity in the learned, exponentially-tilted distributions on $z_{q,p}$.
-However, FFT-parameterized energy functions were also found to be non-smooth and potentially discontinuous, which is also illustrated by the distributions plotted in Fig.
-32.
-This might have compromised adherence to KART and future work might explore incorporating regularization to improve their smoothness.
-Overall, the results show that the choice of basis functions serves as an additional mechanism for encoding inductive bias, complementing the role of the reference prior.
-Once again, importance sampling only proved to be viable when appropriate inductive biases were encoded, and some combinations of bases and reference priors failed to learn meaningful features, (demonstrated by the extra samples of Sec.
-A.7.3).
+(b) No ref. prior
+(c) Gaussian ref. prior
+(d) Uniform ref. prior
+Figure 3: Generated FMNIST examples using different reference priors
+KAEM consistently generated a diverse set of samples when non-restrictive reference priors were used, (Gaussian, uniform, or no reference), thus validating importance sampling, independent energy-based priors, and strict adherence to KART for simple datasets.
+Fig. 4 depicts a single latent component of KAEM trained on FMNIST with a Gaussian reference prior, demonstrating KAEM’s potential for latent prior discovery.
+Generally, priors were modified minimally when trained with IS, and remained largely aligned with their reference distributions.
+This suggests that the reference prior used to initialize KAEM guides the structure of the resulting latent space and may provide a useful source of inductive bias or reusable structure.
+Figure 4: Component $p_{q=1,p=1}(z)$ after training on FMNIST. The density has been normalized by $Z_{q=1,p=1}$ using quadrature, however there is a small discretization bias preventing exactness.
 3.2 Langevin Monte Carlo and Steppingstone estimator
 KAEM, when adherent to KART, can be interpreted as a dense or spline-based network.
-However, for spatial data, convolutional neural networks (CNNs) (LeCun et al., 2015), are more suitable as the generator, $\mathcal{G}$.
+However, for spatial data, convolutional neural networks (CNNs) (LeCun2015Deep), are more suitable as the generator, $\mathcal{G}$.
 In addition to their computational efficiency through weight sharing, CNNs offer a key advantage over dense networks due to their spatial inductive bias: sequences of convolutional layers implement shift-equivariant transformations that are well-suited for images.
-We evaluated KAEM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015).
+We evaluated KAEM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (37648) and CelebA ($64\times 64$) (liu2015deeplearningfaceattributes).
 The first likelihood model in Eq.
@@
 Due to limited computational resources during this study, (detailed in Sec.
-A.2), training was shortened to 8,000 parameter updates, (70,000 were used in the study by Pang et al.
-(2020)), and significantly smaller versions of the CNN decoders used by Pang et al.
-(2020) were adopted as the generator architectures.
+A.2), training was shortened to 12,000 parameter updates, (70,000 were used in the study by pang2020learninglatentspaceenergybased), and the CNN decoder architectures used by pang2020learninglatentspaceenergybased were adopted as the generators.
 A single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator.
-FFT bases were used for the prior energy functions, and full architectural details are provided under Sec.
-A.5.1.
-Posterior sampling was conducted with 40 steps of ULA, and no reference prior was used, $\pi_{0}=1$.
-KAEM was trained for 40 epochs with a batch size of $N_{x}=100$ on 20,000 training examples.
+Full architectural details are provided under Sec.
+A.6.1.
+Posterior sampling was conducted with 40 steps of ULA, and a standard Gaussian reference prior was used, $\pi_{0}=\mathcal{N}(0,1)$.
+KAEM was trained for 30 epochs with a batch size of $N_{x}=50$ on 20,000 training examples.
 All remaining hyperparameters are in Tab.
-6.
+5.
 Annealing was conducted with $N_{t}=5$ power posteriors and the $p$ scheduling outlined in Sec.
@@
 MLE SE
-$\overline{\text{FID}}_{\infty}=108.40$ $\overline{\text{FID}}_{\infty}=131.85$
-$\overline{\text{KID}}_{\infty}=0.0955$ $\overline{\text{KID}}_{\infty}=0.1188$
+$\overline{\text{FID}}_{\infty}=139.37$ $\overline{\text{FID}}_{\infty}=213.26$
+$\overline{\text{KID}}_{\infty}=0.1597$ $\overline{\text{KID}}_{\infty}=0.2047$
 Table 2: Effectively unbiased FID and KID scores (lower is better) for CelebA ($64\times 64$).
 MLE SE
-$\overline{\text{FID}}_{\infty}=122.18$ $\overline{\text{FID}}_{\infty}=127.54$
-$\overline{\text{KID}}_{\infty}=0.1349$ $\overline{\text{KID}}_{\infty}=0.1420$
+$\overline{\text{FID}}_{\infty}=139.37$ $\overline{\text{FID}}_{\infty}=160.18$
+$\overline{\text{KID}}_{\infty}=0.1597$ $\overline{\text{KID}}_{\infty}=0.17044$
 (a) Real samples
@@
 Despite being trained with reduced complexity for a small iteration count, our mixture prior model was able to generate low quality samples of SVHN and CelebA.
-This demonstrates the potential of component-wise ITS for exact and highly efficient sampling in replacement to the approach of Pang et al.
-(2020), which relied on ULA to sample from the prior, thus introducing bias and inefficiency.
+This demonstrates the potential of component-wise ITS for exact and highly efficient sampling in replacement to the approach of pang2020learninglatentspaceenergybased, which relied on ULA to sample from the prior, thus introducing bias and inefficiency.
 Our proposed SE training strategy resulted in worse image metrics for both datasets.
-However, from our subjective perspective, the generated images appeared more well-defined, yet less diverse under SE than under MLE.
+However, from our subjective perspective, the generated CelebA images appeared more well-defined, yet less diverse under SE than under MLE.
+Future work could investigate when SE might be preferred over MLE, as SVHN performed well under the MLE criterion, whereas CelebA subjectively benefited from SE to produce human-perceptible features.
+For example, a more optimally tuned annealing schedule might have sharpened SVHN features in a manner similar to its effect on CelebA.
 Ultimately, we cannot justify the Thermodynamic training strategy due to its substantial computational overhead.
-With more VRAM than available in this study, it would be possible to parallelize a greater number of temperatures than $N_{t}=5$, which could potentially improve quality, as suggested by the results of Calderhead & Girolami (2009).
+With more VRAM than available in this study, it would be possible to parallelize a greater number of temperatures than $N_{t}=5$, which could potentially improve quality, as suggested by the results of CALDERHEAD20094028.
 More posterior sampling options, such as modified annealing schedules, a larger number of ULA steps, or the use of autoMALA with bias correction and temperature-adaptive step size tuning may further improve results.
-Nevertheless, exploring these options is not worthwhile until the hardware described by Zetta (2024) becomes available, given the added algorithmic complexity, hyperparameter sensitivity, and extended training times involved.
+Nevertheless, exploring these options might not be worthwhile until the hardware described by exa2024 becomes available, given the added algorithmic complexity, hyperparameter sensitivity, and extended training times involved.
 4 Conclusions
-KAEM provides a principled and scalable framework for structured and informed generative modeling.
-Our results demonstrate that correctly chosen basis functions and reference priors can enable more efficient training on simple or well-understood datasets through importance sampling.
+KAEM provides a principled, scalable framework for structured and informed generative modeling.
+Our results demonstrate that importance sampling can enable more efficient training on simple or low-dimensional datasets.
 We also show how the Kolmogorov-Arnold representation theorem can inform architectural design and introduce structural bias, allowing KAEM’s architecture to be guided entirely by the size of its latent space.
-Moreover, we find that the inverse transform method offers clear advantages over Langevin Monte Carlo sampling for energy-based priors in terms of speed and exact sampling.
+Moreover, we find that the inverse transform method offers advantages over Langevin Monte Carlo sampling for energy-based priors in terms of speed and exact sampling.
 However, the benefits of annealing and population-based MCMC remain inconclusive, suggesting directions for further work.
-Future investigations into the practicality of our annealing methods depend on continued hardware advancements, such as those introduced by Zetta (2024).
+Future investigations into the practicality of our annealing methods depend on continued hardware advancements, such as those introduced by exa2024.
 4.1 Future work
 4.1.1 Hardware acceleration
-Benchmarks are provided in Sec.
-A.2.
-The univariate, localized nature of spline interpolations complicates their parallelization on GPUs, since they involve sparse and recursive computations.
+Benchmarks are provided in Sec. A.2, compiled to MLIR/XLA using Reactant (Reactantjl2025) and EnzymeMLIR for automatic differentiation (moses2020enzyme; moses2021enzyme_gpu; moses2022enzyme_parallel).
+The univariate, localized nature of B-spline interpolations complicates their parallelization on GPUs, since they involve sparse and recursive computations, so other RBF bases were used in replacement.
 The search operations in Eqs.
-1424, and 70 were also difficult to parallelize on GPUs, which are generally ill-suited for search algorithms, as they often rely on conditional logic.
+1423, and 71 are also difficult to parallelize with GPU kernels, since GPUs are generally ill-suited for search algorithms, as they often rely on conditional logic.
 Parallelizing conditionals can lead to thread divergence, making GPUs inefficient for these operations.
+To compile the model with Reactant, these algorithms were instead written using sparse multiplications for conditional masking, which was efficient but suboptimal on GPUs.
 Additionally, while ITS enables fast inference and scales when the prior is a mixture model, KAEM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA.
@@
 Nonetheless, the unique advantages of KAEM justify continued research and development.
-Moreover, Zetta (2024) introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control that can mitigate the challenges of sparsity, recursion, and irregular memory access patterns.
-The LFU presents a promising avenue for more efficient implementations of ULA, and improved posterior sampling, potentially including bias correction and adaptive step size tuning with autoMALA, (Sec.
-A.6.4).
+Moreover, exa2024 introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control that can mitigate the challenges of sparsity, recursion, and irregular memory access patterns.
+The LFU presents a promising avenue for more efficient implementations of ULA, and improved posterior sampling, potentially including bias correction and adaptive step size tuning with autoMALA (bironlattes2024automalalocallyadaptivemetropolisadjusted) to assist when a single step size is insufficient for all power posterior geometries.
 Even without annealing, replacing ULA with autoMALA may yield a more robust probabilistic model with stronger theoretical guarantees.
-Overall, KAEM may prove more scalable than conventional generative models, owing to its reliance on finite representations of continuous functions and its compatibility with the LFU.
+KAEM might prove more scalable than conventional generative models, owing to its reliance on finite representations of continuous functions and its compatibility with the LFU.
 4.1.2 More opportunities for importance sampling
@@
 3.2.
-This may be attributed to prior-posterior mismatch, which rendered the IS estimator ineffective.
+IS is limited by estimator variance, whereas ULA is often limited by mixing and when sampling multimodal distributions, (addressed in this study with annealing).
+Therefore, the failure of IS for the more complex datasets might be attributed to prior-posterior mismatch, which rendered the IS estimator ineffective.
 Moreover, relying on IS restricts the prior to its initial domain, since proposals are generated from the prior using ITS and finite-domain KAN energy functions.
-In contrast, ULA can explore values of $\bm{z}$ beyond this domain, enabling adaptation of the prior domain within the grid-updating scheme for KANs proposed by Liu et al.
-(2024).
-A possible remedy for the curse of dimensionality is to learn in a reduced-order space using Principal Component Analysis (PCA) (Jolliffe & Cadima, 2016), similar to the approach of Kingma et al.
-(2014), who also applied PCA to train VAEs.
+In contrast, ULA can explore values of $\bm{z}$ beyond this domain, enabling adaptation of the prior domain within the grid-updating scheme for KANs proposed by liu2024kankolmogorovarnoldnetworks.
+A possible remedy for the curse of dimensionality is to learn in a reduced-order space using Principal Component Analysis (PCA) (Jolliffe2016PCA), similar to the approach of kingma2014semisupervisedlearningdeepgenerative, who also applied PCA to train VAEs.
 To address prior-posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins.
 Additionally, efforts should be directed toward selecting appropriate initial domains or adapting the domains dynamically.
+Additionally, KAEM could be used to learn domain-specific priors that transfer across generator architectures.
+In this setting, IS may serve as an efficient strategy for rapidly retraining a new generator on a prior that is already aligned with the posterior before retraining.
 4.1.3 Expressivity
@@
 Additionally, it is often desirable to enrich the latent space with more expressive distributions.
-For example, Normalizing Flows (Papamakarios et al., 2021) achieve this by applying a sequence of reversible transformations that progressively map a simple base distribution into a complex one.
+For example, Normalizing Flows (papamakarios2021normalizingflowsprobabilisticmodeling) achieve this by applying a sequence of reversible transformations that progressively map a simple base distribution into a complex one.
 This might be achievable in KAEM by introducing more latent distributions, or revising Eq.
-13 into a Mixture of Experts framework (Cai et al., 2024) where $\alpha_{q,p}$ could be realized as a more flexible kernel to enable different inductive biases.
-However, this might be expensive for the model by Pang et al.
-(2020) since its use of Langevin sampling is computationally intense.
-For example, the previous work of Cui & Han (2024), which divides their latent EBM space using the diffusion strategy by Ho et al.
-(2020), states that sampling is expensive due to their use of MCMC methods.
+13 into a Mixture of Experts framework (cai2024surveymixtureexperts) where $\alpha_{q,p}$ could be realized as a more flexible kernel to enable different inductive biases.
+However, this might be expensive for the model by pang2020learninglatentspaceenergybased since its use of Langevin sampling is computationally intense.
+For example, the previous work of cui2024learninglatentspacehierarchical, which divides their latent EBM space using the diffusion strategy by ho2020denoisingdiffusionprobabilisticmodels, states that sampling is expensive due to their use of MCMC methods.
 KAEM could potentially circumvent the additional cost by using IS as a fast sampler, though this would require addressing its current limitations.
@@
 A.2 Computational benchmarks
-These benchmarks were collected for FP32 on a GPU (NVIDIA GeForce RTX 4060) with Julia’s BenchmarkTools package, (JuliaCI, 2024).
-The index search operations in Eqs.
-2414, & 70 were parallelized across 20 threads on a CPU (i9-13900H @ 5.40GHz).
+These benchmarks were collected for FP32 on a GPU (NVIDIA GeForce RTX 4060) with Julia’s BenchmarkTools package, (BenchmarkTools).
+Compilation and reverse differentiation were performed using Julia’s Reactant (Reactantjl2025) and Enzyme (moses2020enzyme; moses2021enzyme_gpu; moses2022enzyme_parallel) packages.
 Figure 7: Benchmarks for varying latent dimensions, $n_{z}$, while sampling from independent EBM priors with ITS and estimating posterior expectations with IS.
@@
 They reflect the performance of a reverse autodifferentiation of the MLE criterion, Eq.
-17.
+16.
 Figure 8: Benchmarks for varying amounts of temperatures, $N_{t}$, using a mixture of EBM priors (Tab.
-7) with ITS sampling.
+6) with ITS sampling.
 Posterior expectations were estimated with SE and population-based ULA.
@@
 They reflect the computational cost of performing a single reverse autodifferentiation of the SE criterion, Eq.
-32.
+31.
 A.3 Guidance on tempering
 A.3.1 Cyclic annealing
-Calderhead & Girolami (2009) demonstrated that uniformly scheduling the discretized temperatures in Eq.
+CALDERHEAD20094028 demonstrated that uniformly scheduling the discretized temperatures in Eq.
 30 results in unequal contributions of adjacent power posteriors to the Thermodynamic Integral.
@@
 This overlap ensures similarity between adjacent power posteriors, as introduced in Sec.
-A.6.1, which influences the Monte Carlo estimator in Eq.
-33, given that SE is an application of importance sampling, (Annis et al., 2019).
+A.5.1, which influences the Monte Carlo estimator in Eq.
+32, given that SE is an application of importance sampling, (Annis2019).
 This overlap is quantified by the KL divergence term in Eq.
-52.
-Maintaining small KL divergences between adjacent power posteriors can be achieved by clustering temperatures in regions where $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$ changes rapidly.
-To accomplish this, we follow Calderhead & Girolami (2009) in adopting a power-law relationship to schedule the temperatures of Eq.
-31:
+50.
+Maintaining small KL divergences between adjacent power posteriors can be achieved by clustering temperatures in regions where $\mathbb{P}(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$ changes rapidly.
+To accomplish this, we follow CALDERHEAD20094028 in adopting a power-law relationship to schedule the temperatures of Eq.
+30:
 $t_{k}=\left(\frac{k}{N_{t}}\right)^{p},\quad k=0,1,\ldots,N_{t},$ (35)
@@
 1.
-The study by Calderhead & Girolami (2009) derives an analytic term for an optimal tempering distribution in the context of linear regression models, where the power posterior in Eq.
+The study by CALDERHEAD20094028 derives an analytic term for an optimal tempering distribution in the context of linear regression models, where the power posterior in Eq.
 29 is used to evaluate the plausibility of model parameters in relation to the data, rather than reflecting a latent variable posterior.
@@
 Figure 11: Cosine $p$ schedule, increasing from 1 to 8
-Initializing $p\leq 1$ focuses learning on high-variance, data-informed likelihoods during the early stages of the cycle.
-Here, features emerge closer to the true posterior, and learning requires exploration and higher variance gradient estimates.
-As training progresses, KAEM will have learned global features sufficiently well, reducing the meaningful change of measure between adjacent power posteriors of a $p\leq 1$ schedule.
-At this stage, $p(i)$ will have transitioned to $p>1$, thereby shifting focus towards exploitation and sculpting the prior with lower variance gradient estimates.
+Initializing $p>1$ focuses learning on low-variance, smoother power posteriors during the early stages of the cycle, where global features emerge between adjacent power posteriors.
+As training progresses, KAEM will have learned the global features sufficiently well, reducing the change of density between adjacent power posteriors of a $p>1$ schedule.
+At this stage, $p(i)$ will have transitioned to $p\leq 1$, thereby shifting KAEM’s focus towards capturing more high-variance, data-informed features of the posterior landscape, where training is expected to enhance its ability to generate finer details.
 Guidance on selecting optimal values for the tempering hyperparameters is provided in Tab. 3.
@@
 Although increasing this value improves the quality of samples at each stage, it also raises the computational cost, much like $N_{t}$.
-It can be kept at a small value, such as $N_{\text{local}}=40$, (based on the recommendations of Nijkamp et al.
-(2019b; a) regarding short-run Markov chain Monte Carlo).
+It can be kept at a small value, such as $N_{\text{local}}=40$, (based on the recommendations of nijkamp2019learningnonconvergentnonpersistentshortrun; nijkamp2019anatomymcmcbasedmaximumlikelihood regarding short-run Markov chain Monte Carlo).
 $N_{\text{unadjusted}}>1$ Controls the number of autoMALA iterations performed without MH adjustments, reflecting the level of exploration at each power posterior.
-This can be set to 1, as guided by Biron-Lattes et al.
-(2024).
-$p_{\text{init}}\leq 1,$ $p_{\text{end}}>1,$ $N_{\text{cycles}}\geq 0$ Controls the evolution of tempering during training.
+This can be set to 1, as guided by bironlattes2024automalalocallyadaptivemetropolisadjusted.
+$p_{\text{init}}>1,$ $p_{\text{end}}\leq 1,$ $N_{\text{cycles}}\geq 0$ Controls the evolution of tempering during training.
 If the data is expected to have smooth relationships, reduce both $p_{\text{init}}$ and $p_{\text{end}}$.
 For data with finer details, increase both values to allow for better focus on capturing intricate features.
-The values used for this study were $p_{\text{init}}=1$, and $p_{\text{end}}=8$, since our experience suggests that power posteriors change rapidly closer to the prior.
-The number of cycles, $N_{\text{cycles}}$, controls the periodicity of cosine scheduling similar to cyclic annealing (Fu et al., 2019).
+The values used for this study were $p_{\text{init}}=2$, and $p_{\text{end}}=0.5$, since our experience suggests that power posteriors initially change rapidly closer to the prior.
+The number of cycles, $N_{\text{cycles}}$, controls the periodicity of cosine scheduling similar to cyclic annealing (fu2019cyclicalannealingschedulesimple).
 A.4 Derivations and other identities
@@
 The following short proof for a learning gradient with respect to an arbitrary set of parameters, $\bm{\theta}$, proves useful for other derivations in this study:
-$\mathbb{E}_{P(\bm{x}\mid\bm{\theta})}\left[\nabla_{\bm{\theta}}\log P(\bm{x}\mid\bm{\theta})\right]$ $=\mathbb{E}_{P(\bm{x}\mid\bm{\theta})}\left[\frac{\nabla_{\bm{\theta}}P(\bm{x}\mid\bm{\theta})}{P(\bm{x}\mid\bm{\theta})}\right]=\int_{\mathcal{X}}\frac{\nabla_{\bm{\theta}}P(\bm{x}\mid\bm{\theta})}{P(\bm{x}\mid\bm{\theta})}\,\,\,dP(\bm{x}\mid\bm{\theta})$
-$=\int_{\mathcal{X}}\nabla_{\bm{\theta}}P(\bm{x}\mid\bm{\theta})\,d\bm{x}=\nabla_{\bm{\theta}}\int_{\mathcal{X}}P(\bm{x}\mid\bm{\theta})\,d\bm{x}=\nabla_{\bm{\theta}}1=0.$ (38)
+$\mathbb{E}_{p_{\bm{\theta}}(\bm{x})}\left[\nabla_{\bm{\theta}}\log p_{\bm{\theta}}(\bm{x})\right]$ $=\mathbb{E}_{p_{\bm{\theta}}(\bm{x})}\left[\frac{\nabla_{\bm{\theta}}p_{\bm{\theta}}(\bm{x})}{p_{\bm{\theta}}(\bm{x})}\right]=\int_{\mathcal{X}}\frac{\nabla_{\bm{\theta}}p_{\bm{\theta}}(\bm{x})}{p_{\bm{\theta}}(\bm{x})}\,\,\,dp_{\bm{\theta}}(\bm{x})$
+$=\int_{\mathcal{X}}\nabla_{\bm{\theta}}p_{\bm{\theta}}(\bm{x})\,d\bm{x}=\nabla_{\bm{\theta}}\int_{\mathcal{X}}p_{\bm{\theta}}(\bm{x})\,d\bm{x}=\nabla_{\bm{\theta}}1=0.$ (36)
 A.4.2 Marginal likelihood
-$\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\left[\log\left(P(\bm{x}^{(b)},\bm{z}\mid\bm{f},\bm{\Phi})\right)\right]$
-$=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\left[\log\left(P\left(\bm{z}\mid\bm{f}\right)\cdot P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)\right]$
-$=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\left[\log\left(P\left(\bm{z}\mid\bm{f}\right)\right)+\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)\right].$ (39)
+$\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\left[\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)},\bm{z})\right)\right]$
+$=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\left[\log\left(p_{\bm{f}}\left(\bm{z}\right)\cdot p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)\right]$
+$=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\left[\log\left(p_{\bm{f}}\left(\bm{z}\right)\right)+\log\left(p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)\right].$ (37)
 A.4.3 Maximum likelihood learning gradient
@@
 The prior model is solely dependent on $\bm{f}$ in Eq.
-5.
+4.
 We can rewrite Eq.
 17 as:
-$\nabla_{\bm{f}}\log\left(P\left(\bm{z}\mid\bm{f}\right)\right)$ $=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(\pi_{q,p}(z_{q,p})\right)=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)$ (41)
+$\nabla_{\bm{f}}\log\left(p_{\bm{f}}\left(\bm{z}\right)\right)$ $=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(\pi_{q,p}(z_{q,p})\right)=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)$ (39)
 where $Z_{q,p}=\int_{\mathcal{Z}}\exp(f_{q,p}(z_{p}))\,d\pi_{0}(z_{p})$ is the normalization constant. We can simplify the constant using Eq. 36:
-$\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg[\nabla_{\bm{f}}\log P\left(\bm{z}\mid\bm{f}\right)\Bigg]$ $=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)\Bigg]=0$
-$=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]-\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(Z_{q,p}\right)=0$
-$\therefore\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(Z_{q,p}\right)$ $=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]$ (42)
+$\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\log p_{\bm{f}}\left(\bm{z}\right)\Bigg]$ $=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)\Bigg]=0$
+$=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]-\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(Z_{q,p}\right)=0$
+$\therefore\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(Z_{q,p}\right)$ $=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]$ (40)
 Therefore the log-prior gradient of Eq. 39 can be evaluated without quadrature normalization as:
-$\nabla_{\bm{f}}\log P\left(\bm{z}\mid\bm{f}\right)$ $=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})-\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg],$ (43)
+$\nabla_{\bm{f}}\log p_{\bm{f}}\left(\bm{z}\right)$ $=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})-\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg],$ (41)
 and the learning gradient attributed to the prior is finalized into a contrastive divergence format:
-$\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\left[\log P\left(\bm{z}\mid\bm{f}\right)\right]\Bigg]=\;$ $\mathbb{E}_{P(\bm{z}\mid\bm{x},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]$
--\; $\mathbb{E}_{P(\bm{z}\mid\bm{f})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg].$ (44)
+$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\bm{f}}\left[\log p_{\bm{f}}\left(\bm{z}\right)\right]\Bigg]=\;$ $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]$
+-\; $\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg].$ (42)
 A.4.5 Deriving the Thermodynamic Integral
 Starting from the power posterior:
-$P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)=\frac{P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\,\,P(\bm{z}\mid\bm{f})}{\bm{Z}_{t}}$ (45)
+$p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)=\frac{p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\,\,p_{\bm{f}}(\bm{z})}{\bm{Z}_{t}}$ (43)
 The partition function $\bm{Z}_{t}$ is given by:
-$\bm{Z}_{t}=\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]$ (46)
+$\bm{Z}_{t}=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]$ (44)
 Now, differentiating the log-partition function with respect to $t$:
-$\frac{\partial}{\partial t}\log(\bm{Z}_{t})$ $=\frac{1}{\bm{Z}_{t}}\frac{\partial}{\partial t}\bm{Z}_{t}=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[\frac{\partial}{\partial t}P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]$
-$=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\cdot P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{t}\right]\$
-$=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]$ (47)
+$\frac{\partial}{\partial t}\log(\bm{Z}_{t})$ $=\frac{1}{\bm{Z}_{t}}\frac{\partial}{\partial t}\bm{Z}_{t}=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[\frac{\partial}{\partial t}p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]$
+$=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\cdot p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]\$
+$=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]$ (45)
 Integrating both sides from 0 to 1:
-$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\,dt$ (48)
-Using the fundamental theorem of calculus, (Larson & Edwards, 2008):
-$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\,dt$ (49)
+$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\,dt$ (46)
+Using the fundamental theorem of calculus, (larson2008calculus):
+$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\,dt$ (47)
 The limits of the integrals can be expressed as:
-$\log(\bm{Z}_{t=0})$ $=\log\Big(\mathbb{E}_{P(\bm{z}\mid\bm{f})}\left[P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)^{0}\right]\Big)=\log(1)=0$
-$\log(\bm{Z}_{t=1})$ $=\log\Big(\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\Big)=\log P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})$ (50)
+$\log(\bm{Z}_{t=0})$ $=\log\Big(\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{0}\right]\Big)=\log(1)=0$
+$\log(\bm{Z}_{t=1})$ $=\log\Big(\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\Big)=\log\mathbb{P}(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})$ (48)
 Therefore, we obtain the log-marginal likelihood expressed as the Thermodynamic Integral:
-$\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\log P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})=\int_{0}^{1}\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t)}\left[\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right]\,dt$ (51)
+$\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\log\mathbb{P}(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\,dt$ (49)
 A.4.6 Discretizing the Thermodynamic Integral
-Here, $t_{k}$ denotes the tempering at the $k$-th index of the schedule, and $N_{\text{t}}$ is the number of temperatures. As derived by Calderhead & Girolami (2009), Eq. 29 then be evaluated using:
-$\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(P_{t_{k-1}}||P_{t_{k}})-D_{\text{KL}}(P_{t_{k}}||P_{t_{k-1}})$ (52)
+Here, $t_{k}$ denotes the tempering at the $k$-th index of the schedule, and $N_{\text{t}}$ is the number of temperatures. As derived by CALDERHEAD20094028, Eq. 28 then be evaluated using:
+$\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(\mathbb{P}_{t_{k-1}}||\mathbb{P}_{t_{k}})-D_{\text{KL}}(\mathbb{P}_{t_{k}}||\mathbb{P}_{t_{k-1}})$ (50)
 Where:
-$\Delta t_{k}$ $=t_{k}-t_{k-1},\quad\quad E_{k}=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\left[\log\left(P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\right)\right],$
-$D_{\text{KL}}(P_{t_{k-1}}||P_{t_{k}})$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}\Bigg[\log\Bigg(\frac{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Bigg)\Bigg]$ (53)
+$\Delta t_{k}$ $=t_{k}-t_{k-1},\quad\quad E_{k}=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\left[\log\left(p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)\right],$
+$D_{\text{KL}}(\mathbb{P}_{t_{k-1}}||\mathbb{P}_{t_{k}})$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}\Bigg[\log\Bigg(\frac{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Bigg)\Bigg]$ (51)
 This reflects the trapezium rule for numerical integration, supported by bias correction to provide an estimate of the Riemann integral that is especially robust to error.
 The learning gradient derived from the discretized Thermodynamic Integral is unchanged from the MLE gradient in Eq.
-17, as demonstrated under Sec.
+16, as demonstrated under Sec.
 A.4.7.
@@
 Starting from the definitions attributed to the KL divergence terms:
-$D_{\text{KL}}(P_{t_{k-1}}\|P_{t_{k}})$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}\Big[\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\Big]$
-$\text{and}\quad D_{\text{KL}}(P_{t_{k}}\|P_{t_{k-1}})$ $=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})\Big]$
-$=-\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\Big]$ (54)
+$D_{\text{KL}}(\mathbb{P}_{t_{k-1}}\|\mathbb{P}_{t_{k}})$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}\Big[\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})\Big]$
+$\text{and}\quad D_{\text{KL}}(\mathbb{P}_{t_{k}}\|\mathbb{P}_{t_{k-1}})$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})\Big]$
+$=-\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})\Big]$ (52)
 We can leverage Eq. 29 to quantify the log-difference in terms of the change in likelihood contribution plus some normalization constant:
-$\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})$ $-\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})=-\Delta t_{k}\cdot\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$ (55)
+$\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})$ $-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})=-\Delta t_{k}\cdot\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$ (53)
 The definitions in Eq. 51 can be used to express the KL divergences in terms of the expected likelihoods:
-$\therefore D_{\text{KL}}(P_{t_{k-1}}\|P_{t_{k}})$ $=-\Delta t_{k}\cdot\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k-1})}\Big[\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\Big]+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
+$\therefore D_{\text{KL}}(\mathbb{P}_{t_{k-1}}\|\mathbb{P}_{t_{k}})$ $=-\Delta t_{k}\cdot\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}\Big[\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\Big]+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
 $=-\Delta t_{k}\cdot E_{k-1}+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$ (54)
-$\text{and}\quad D_{\text{KL}}(P_{t_{k}}\|P_{t_{k-1}})$ $=\Delta t_{k}\cdot\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\Big]-\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
+$\text{and}\quad D_{\text{KL}}(\mathbb{P}_{t_{k}}\|\mathbb{P}_{t_{k-1}})$ $=\Delta t_{k}\cdot\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\Big]-\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
 $=\Delta t_{k}\cdot E_{k}-\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$ (55)
 Substituting back into Eq. 50, we eliminate the expected likelihood terms:
-$\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(P_{t_{k-1}}\|P_{t_{k}})$ $-D_{\text{KL}}(P_{t_{k}}\|P_{t_{k-1}})=-\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
-$\therefore\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)$ $=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(P_{t_{k-1}}||P_{t_{k}})-D_{\text{KL}}(P_{t_{k}}||P_{t_{k-1}})$
+$\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(\mathbb{P}_{t_{k-1}}\|\mathbb{P}_{t_{k}})$ $-D_{\text{KL}}(\mathbb{P}_{t_{k}}\|\mathbb{P}_{t_{k-1}})=-\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
+$\therefore\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)$ $=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(\mathbb{P}_{t_{k-1}}||\mathbb{P}_{t_{k}})-D_{\text{KL}}(\mathbb{P}_{t_{k}}||\mathbb{P}_{t_{k-1}})$
 $=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-\Bigg[\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)\Bigg]$
@@
 Following a similar method to Sec. A.4.4 we can use Eq. 36 to express the learning gradient attributed to each normalization constant in a simpler manner:
-$\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[\nabla_{\bm{f},\bm{\Phi}}\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\Big]$
-$=\,$ $\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})+\nabla_{\bm{f}}\log P(\bm{z}\mid\bm{f})\Big]-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=0$
-$\therefore$ $\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})+\nabla_{\bm{f}}\log P(\bm{z}\mid\bm{f})\Big]$ (59)
+$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\nabla_{\bm{f},\bm{\Phi}}\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})\Big]$
+$=\,$ $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})+\nabla_{\bm{f}}\log p_{\bm{f}}(\bm{z})\Big]-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=0$
+$\therefore$ $\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})+\nabla_{\bm{f}}\log p_{\bm{f}}(\bm{z})\Big]$ (57)
 Telescoping the sum in Eq. 56, and noting that $k=1$ corresponds to the power posterior equaling the prior, and $k=N_{t}$ to the posterior, Eq. 56 yields KAEM’s learning gradient:
-$\nabla_{\bm{f},\bm{\Phi}}\log\left(P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})\right)$ $=\sum_{k=1}^{N_{t}}\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k-1}})$
+$\nabla_{\bm{f},\bm{\Phi}}\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)$ $=\sum_{k=1}^{N_{t}}\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k-1}})$
 $=\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{N_{t}}})-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{0}})$
-$=\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}\Bigg[\nabla_{\bm{f}}\log P\left(\bm{z}\mid\bm{f}\right)+\nabla_{\bm{\Phi}}\log P\left(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi}\right)\Bigg]$
+$=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\Bigg[\nabla_{\bm{f}}\log p_{\bm{f}}\left(\bm{z}\right)+\nabla_{\bm{\Phi}}\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\Bigg]$
 […]
-A.5 Hyperparameters and training details
+A.5 Sampling theory
+This section provides more theoretical detail regarding KAEM’s sampling procedures.
+A.5.1 Importance Sampling
+Importance Sampling (IS) is a Monte Carlo that enables the estimation of expectations with respect to one probability measure using samples drawn from another.
+For a detailed exposition, see 910e8e7d-2e70-317c-adf6-e21cd50755a1.
+The Radon-Nikodym derivative
+IS involves rewriting the expectation of an arbitrary function or density, $\rho(\bm{z})$, under the target measure, $\mathbb{P}(\bm{z})$, in terms of a more tractable proposal measure, $q(\bm{z})$.
+Specifically, we use the Radon-Nikodym derivative $\frac{d\mathbb{P}}{d\mathbb{Q}}$ to rewrite the expectation of an arbitrary function of $\bm{z}$, denoted as $\rho(\bm{z})$:
+$\mathbb{E}_{\mathbb{P}(\bm{z})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\,\mathbb{P}(d\bm{z})=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\cdot\frac{d\mathbb{P}}{d\mathbb{Q}}\,\mathbb{Q}(d\bm{z}).$ (59)
+Here, $\frac{d\mathbb{P}}{d\mathbb{Q}}$ acts as a weight that adjusts the contribution of each sample to account for the difference between the target and proposal measures.
+This requires that the proposal measure $\mathbb{Q}$ is absolutely continuous with respect to the target measure $\mathbb{P}$, i.e.
+$\mathbb{Q}(A)=0$ implies $\mathbb{P}(A)=0$, for every measurable subset $A\in\mathcal{Z}$.
+In the context of KAEM, the target distribution is the posterior, $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$, which is proportional to the product of the likelihood and the prior:
+$p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})\propto p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\cdot p_{\bm{f}}(\bm{z}).$ (60)
+The exact normalization constant, (the marginal likelihood $p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})$), is intractable, but Importance Sampling circumvents the need to compute it explicitly.
+Instead, the expectation of $\rho(\bm{z})$ under the posterior can be expressed as:
+$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})w(\bm{z})\,\mathbb{Q}(d\bm{z}),$ (61)
+where the importance weights $w(\bm{z})$ are given by:
+$w(\bm{z})=\frac{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}{q(\bm{z})}=\frac{p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\cdot p_{\bm{f}}(\bm{z})}{q(\bm{z})}.$ (62)
+Proposal distribution and importance weights
+A practical choice for the proposal density $q(\bm{z})$ is the prior, $p_{\bm{f}}(\bm{z})$, given the availability of draws from the prior using the method outlined in Sec.
+2.4.3.
+This simplifies the importance weights, as the prior has a tractable form and covers the support of the posterior.
+This yields:
+$w(\bm{z}^{(s)})=p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)})\propto\exp(\log(p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)}))).$ (63)
+This reflects that the likelihood now directly informs how much weight to assign to each sample $\bm{z}$.
+Intuitively, samples that better explain the observed data $\bm{x}^{(b)}$, (as measured by the likelihood), are given higher importance.
+Using the definition of softmax, the normalized weights are given by:
+$\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big(\log(p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)}))\Big).$ (64)
+Resampling
+The suitability of these weights depends on how well $q(\bm{z})$ matches the posterior.
+In KAEM, there is likely to be significant mismatch due to the complexity of the likelihood models described in Eq.
+20.
+Consequently, the weights will exhibit high variance, resulting in many samples contributing minimally.
+This reduces the Effective Sample Size (ESS) and introduces bias into the estimates.
+To address this, the weights are resampled before proceeding with the posterior expectation.
+This is accomplished by using one of the methods outlined by douc2005comparisonresamplingschemesparticle, whenever the ESS falls below a certain threshold determined by Eq.
+26.
+Resampling redistributes the sample weights, creating a new population with uniformly distributed weights while preserving the statistical properties of the original distribution.
+As training progresses and the prior becomes a better match for the posterior, the need for resampling decreases.
+After resampling, the corresponding weights will be uniformly distributed and a standard Monte Carlo estimator can be applied instead of the IS estimator.
+Importance Sampling estimator
+To estimate the expectation, we draw $N_{s}$ independent samples $\{\bm{z}^{(s)}\}_{s=1}^{N_{s}}\sim q(\bm{z})$ using the procedure outlined in Sec. 2.4.3 and compute a weighted sum:
+$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}[\rho(\bm{z})]$ $\approx\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)})\cdot w_{\text{norm}}(\bm{\bm{z}^{(s)})}$ (65)
+$\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})$. (66)
+The estimator is unbiased, implying that its expectation matches the true posterior expectation.
+As $N_{s}\to\infty$, the approximation converges almost surely to the true value, provided that $q(\bm{z})$ sufficiently covers the posterior to ensure a well-conditioned Radon-Nikodym derivative.
+This estimator is used for the log-marginal likelihood and its gradient in Eqs.
+15 &  16, since it provides a means of estimating expectations with respect to the posterior distribution.
+A.5.2 Residual resampling
+In this study, we adopted residual resampling to filter the population of latent samples whenever the effective sample size (ESS) criterion in Eq.
+26 is satisfied.
+Given a set of normalized weights $\{w_{\text{norm}}(\bm{z}^{(s)})\}_{s=1}^{N}$, the method deterministically replicates samples according to the integer part of $Nw_{\text{norm}}$, and uses multinomial sampling to allocate the remaining draws according to the residual mass.
+  1. 1.
+    Integer replication: Each sample $s$ is deterministically replicated
+    $r_{s}=\left\lfloor N\,w_{\text{norm}}(\bm{z}^{(s)})\right\rfloor,$ (67)
+    resulting in $\sum_{s=1}^{N}r_{s}$ allocated indices.
+  2. 2.
+    Residual weights: The number of remaining draws is
+    $R=N-\sum_{s=1}^{N}r_{s}.$ (68)
+    The unallocated probability mass (fractional remainder) is computed as
+    $m_{s}=N\,w_{\text{norm}}(\bm{z}^{(s)})-r_{s}.$ (69)
+    If $R>0$, the residual weights are obtained by normalizing this mass:
+    $w_{\text{residual}}(\bm{z}^{(s)})=\frac{m_{s}}{R},$ (70)
+    otherwise $w_{\text{residual}}(\bm{z}^{(s)})=0$ for all $s$.
+  3. 3.
+    Multinomial allocation: The remaining $R$ indices are sampled according to the cumulative distribution function of the residual weights. For each draw,
+    $k^{*}=\min\left\{j\mid\sum^{j}_{s=1}w_{\text{residual}}(\bm{z}^{(s)})\geq u_{i}\right\}\quad\text{where}\quad u_{i}\sim\mathcal{U}([0,1]).$ (71)
+    The final resampled index set consists of the deterministic replications given by $\{r_{s}\}$ and the stochastic residual allocations drawn from Eq.
+71.
+After resampling, weights are reset to a uniform distribution and a standard Monte Carlo estimator can be used.
+A.6 Hyperparameters and training details
 Table 4: Hyperparameters for Sec. 3.1.1.
@@
 layer_widths $p,q$ 40, 81
-spline_type (-) variable
-spline_activation (-) Silu
+spline_type (-) RBF
+spline_activation (-) Relu
 layer_norm (-) false
@@
 spline_degree $d$ 3
-grid_size $N_{\text{grid}}$ RBF: 20
-FFT: 50
+grid_size $N_{\text{grid}}$ 20
 grid_update_ratio $\alpha_{\text{grid}}$ 0.05
 Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
-Lognormal: $[0,4]$
-Gaussian: [1.2,1.2][-1.2,1.2]
+Gaussian: [1.5,1.5][-1.5,1.5]
+No ref.: [1.2,1.2][-1.2,1.2]
 Spline basis $\varepsilon$ init. $\varepsilon$ 0.1
@@
 Train basis param. $\tau_{\text{trainable}}$ true
-GaussQuad_nodes $N_{\text{quad}}$ 100
+GaussQuad_nodes $N_{\text{quad}}$ 200
 Use mixture prior (-) false
@@
 betas (-) (0.9, 0.999)
-Table 5: Hyperparameters for Sec. 3.1.2.
+Table 5: Hyperparameters for Sec. 3.2.
 EBM Prior
-layer_widths $p,q$ 40, 81
-spline_type (-) variable
-spline_activation (-) Silu
+layer_widths $p,q$ See Sec. A.6.1
+spline_type (-) RBF
+spline_activation (-) Relu
 layer_norm (-) false
@@
 spline_degree $d$ 3
-grid_size $N_{\text{grid}}$ RBF: 20
-FFT: 50
-grid_update_ratio $\alpha_{\text{grid}}$ 0.05
-Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
-Lognormal: $[0,4]$
-Gaussian: [1.2,1.2][-1.2,1.2]
-Spline basis $\varepsilon$ init. $\varepsilon$ 0.1
-Spline basis $\mu$ scaling $\mu$ 1.0
-Basis activation scaling $\sigma_{b}$ 1.0
-Basis spline scaling $\sigma_{s}$ 1.0
-Initial basis param. $\tau_{0}$ 1.0
-Train basis param. $\tau_{\text{trainable}}$ true
-GaussQuad_nodes $N_{\text{quad}}$ 100
-Use mixture prior (-) false
-KAN Likelihood
-hidden_widths $q,2q+1$ 81, 162
-All other KAN hparams Identical to EBM Prior
-generator_noise $\sigma_{\epsilon}$ 0.1
-generator_variance $\sigma_{\text{llhood}}$ 0.1
-output_activation $\sigma_{\text{act}}$ sigmoid
-Importance sampling resampler resampler residual
-ESS threshold $\gamma$ 0.5
-Grid updating
-update_prior_grid (-) true
-update_llhood_grid (-) true
-num_grid_updating_samples (-) 100
-grid_update_frequency (-) 100
-grid_update_decay (-) 0.999
-Training
-batch_size $N_{x}$ 100
-importance_sample_size $N_{s}$ 100
-Training examples $N_{\text{train}}$ 1,000
-Testing examples $N_{\text{test}}$ 1,000
-num_generated_samples (-) 3,000
-Number of epochs $N_{\text{epochs}}$ 1,000
-eps $\epsilon$ 0.001
-Optimizer
-type (-) adam
-learning_rate $\eta_{LR}$ 0.001
-betas (-) (0.9, 0.999)
-Table 6: Hyperparameters for Sec. 3.2.
-EBM Prior
-layer_widths $p,q$ See Sec. A.5.1
-spline_type (-) FFT
-spline_activation (-) LeakyReLU
-layer_norm (-) false
-quadrature_method (-) gausslegendre
-spline_degree $d$ 3
-grid_size $N_{\text{grid}}$ 1
+grid_size $N_{\text{grid}}$ 20
 grid_update_ratio $\alpha_{\text{grid}}$ N/A
-Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}] [$1.5,1.5][-1.5,1.5]
+Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}] [$1,1][-1,1]
 Spline basis $\varepsilon$ init. $\varepsilon$ 0.1
@@
 grid_update_frequency (-) 100
-grid_update_decay (-) 0.9999
+grid_update_decay (-) 0.999
 Training
@@
 iters $N_{\text{local}}$ 40
-step_size $\eta$ 0.005
+step_size $\eta$ 0.002
 THERMODYNAMIC_INTEGRATION
@@
 num_temps $N_{t}$ 5
-Initial exponent in power law schedule $p_{\text{start}}$ 1
-Final exponent in power law schedule $p_{\text{end}}$ 8
+Initial exponent in power law schedule $p_{\text{start}}$ 2
+Final exponent in power law schedule $p_{\text{end}}$ 0.5
 Number of cosine cycles $N_{\text{cycles}}$ 0
-Global swap frequency, replica_exchange_frequency (-) 1
-A.5.1 SVHN & CelebA Architectures
-Table 7: The dense Chebyshev KAN network used as the mixture prior model
+A.6.1 SVHN & CelebA Architectures
+Table 6: The dense RBF KAN network used as the mixture prior model
 Layer Operation Output Dim.
@@
 Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
-1 $\times4$ convT($128$), LeakyReLU $4\times 4\times 128$ 1
-2 $\times4$ convT($64$), LeakyReLU $8\times 8\times 64$ 2
-3 $\times4$ convT($32$), LeakyReLU $16\times 16\times 32$ 2
+1 $\times4$ convT($256$), LeakyReLU $4\times 4\times 256$ 1
+2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
+3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
 4 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
@@
 Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
-1 $\times4$ convT($256$), LeakyReLU $4\times 4\times 256$ 1
-2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
-3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
-4 $\times4$ convT(32), LeakyReLU $32\times 32\times 32$ 2
+1 $\times4$ convT($512$), LeakyReLU $4\times 4\times 512$ 1
+2 $\times4$ convT($256$), LeakyReLU $8\times 8\times 256$ 2
+3 $\times4$ convT($128$), LeakyReLU $16\times 16\times 128$ 2
+4 $\times4$ convT(64), LeakyReLU $32\times 32\times 64$ 2
 5 $\times4$ convT(3), Sigmoid $64\times 64\times 3$ 2
 Config: kernel=4,4,4,4,4 strides=1/2/2/2/2, paddings=0/1/1/1/1, activation=LeakyReLU, batchnorm=False
-A.6 Sampling theory
-This section provides more theoretical detail regarding KAEM’s sampling procedures.
-A.6.1 Importance Sampling
-Importance Sampling (IS) is a Monte Carlo method rooted in the measure-theoretic formulation of probability.
-It enables the estimation of expectations with respect to one probability measure using samples drawn from another.
-For a detailed exposition, see Kloek & van Dijk (1978).
-The Radon-Nikodym derivative
-IS involves rewriting the expectation of an arbitrary function, $\rho(\bm{z})$, under the target measure, $P(\bm{z})$, in terms of a more tractable proposal measure, $Q(\bm{z})$.
-Specifically, we use the Radon-Nikodym derivative $\frac{dP}{dQ}$ to rewrite the expectation of an arbitrary function of $\bm{z}$, denoted as $\rho(\bm{z})$:
-$\mathbb{E}_{P(\bm{z})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\,dP(\bm{z})=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\cdot\frac{dP}{dQ}\,dQ(\bm{z}).$ (61)
-Here, $\frac{dP}{dQ}$ acts as a weight that adjusts the contribution of each sample to account for the difference between the target and proposal measures.
-This requires that the proposal measure $Q$ is absolutely continuous with respect to the target measure $P$, i.e.
-$Q(A)=0$ implies $P(A)=0$, for every measurable subset $A\in\mathcal{Z}$.
-In the context of KAEM, the target distribution is the posterior, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is proportional to the product of the likelihood and the prior:
-$P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})\propto P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\cdot P(\bm{z}\mid\bm{f}).$ (62)
-The exact normalization constant, (the marginal likelihood $P(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})$), is intractable, but Importance Sampling circumvents the need to compute it explicitly.
-Instead, the expectation of $\rho(\bm{z})$ under the posterior can be expressed as:
-$\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})w(\bm{z})\,dQ(\bm{z}),$ (63)
-where the importance weights $w(\bm{z})$ are given by:
-$w(\bm{z})=\frac{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}{Q(\bm{z})}=\frac{P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})\cdot P(\bm{z}\mid\bm{f})}{Q(\bm{z})}.$ (64)
-Proposal distribution and importance weights
-A practical choice for the proposal distribution $Q(\bm{z})$ is the prior, $P(\bm{z}\mid\bm{f})$, given the availability of draws from the prior using the method outlined in Sec.
-2.4.3.
-This simplifies the importance weights, as the prior has a tractable form and covers the support of the posterior.
-This yields:
-$w(\bm{z}^{(s)})=P(\bm{x}^{(b)}\mid\bm{z}^{(s)},\bm{\Phi})\propto\exp(\log(P(\bm{x}^{(b)}\mid\bm{z}^{(s)},\bm{\Phi}))).$ (65)
-This reflects that the likelihood now directly informs how much weight to assign to each sample $\bm{z}$.
-Intuitively, samples that better explain the observed data $\bm{x}^{(b)}$, (as measured by the likelihood), are given higher importance.
-Using the definition of softmax, the normalized weights are given by:
-$\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big(\log(P(\bm{x}^{(b)}\mid\bm{z}^{(s)},\bm{\Phi}))\Big).$ (66)
-Resampling
-The suitability of these weights depends on how well $Q(\bm{z})$ matches the posterior.
-In KAEM, there is likely to be significant mismatch due to the complexity of the likelihood models described in Eq.
-21.
-Consequently, the weights will exhibit high variance, resulting in many samples contributing minimally.
-This reduces the Effective Sample Size (ESS) and introduces bias into the estimates.
-To address this, the weights are resampled before proceeding with the posterior expectation.
-This is accomplished by using one of the methods outlined by Douc et al.
-(2005), whenever the ESS falls below a certain threshold determined by Eq.
-27.
-Resampling redistributes the sample weights, creating a new population with uniformly distributed weights while preserving the statistical properties of the original distribution.
-As training progresses and the prior becomes a better match for the posterior, the need for resampling decreases.
-Importance Sampling estimator
-To estimate the expectation, we draw $N_{s}$ independent samples $\{\bm{z}^{(s)}\}_{s=1}^{N_{s}}\sim Q(\bm{z})$ using the procedure outlined in Sec. 2.4.3 and compute a weighted sum:
-$\mathbb{E}_{P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})}[\rho(\bm{z})]$ $\approx{\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})\cdot w_{\text{norm, resampled}}(\bm{\bar{z}_{\text{resampled}}}^{(s)})}.$ (67)
-The estimator is unbiased, implying that its expectation matches the true posterior expectation.
-As $N_{s}\to\infty$, the approximation converges almost surely to the true value, provided that $Q(\bm{z})$ sufficiently covers the posterior to ensure a well-conditioned Radon-Nikodym derivative.
-This estimator is used for the log-marginal likelihood and its gradient in Eqs.
-16 &  17, since it provides a means of estimating expectations with respect to the posterior distribution.
-A.6.2 Residual resampling
-In this study, we adopted residual resampling to redraw weights.
-This redistributes the population of latent samples according to the normalized weights of Sec.
-A.6.1.
-The procedure is only conducted when Eq.
-27 regarding ESS is satisfied:
-  1. 1.
-    Integer replication: The sample, $s$, is replicated $r_{s}$ times, where:
-    $r_{s}=\left\lfloor N_{s}\cdot w_{\text{norm}}(\bm{z}^{(s)})\right\rfloor$ (68)
-  2. 2.
-    Residual weights: The total number of replicated samples after the previous stage will be $\sum_{s=1}^{N_{s}}r_{s}$.
-Therefore, $N_{s}-\sum_{s=1}^{N_{s}}r_{s}$ remain, which must be resampled using the residuals:
-    $w_{\text{residual}}(\bm{z}^{(s)})$ $=w_{\text{norm}}(\bm{z}^{(s)})\cdot(N_{s}-r_{s})$
-    $w_{\text{norm},\ \text{residual}}(\bm{z}^{(s)})$ $=\frac{w_{\text{residual}}(\bm{z}^{(s)})}{\sum_{s=1}^{N_{s}}w_{\text{residual}}(\bm{z}^{(s)})}$ (69)
-    If a sample has been replicated at the previous stage, then its corresponding $w_{\text{norm},\,\text{residual}}$ will be reduced, leading to a lower probability of it being resampled again.
-  3. 3.
-    Resample: The remaining samples are drawn with multinomial resampling, based on the cumulative distribution function of the residuals:
-    $k^{*}=\min\left\{j\mid\sum^{j}_{s=1}w_{\text{norm},\,\text{residual}}(\bm{z}^{(s)})\geq u_{i}\right\}\quad\text{where}\quad u_{i}\sim\mathcal{U}([0,1]).$ (70)
-    where $k^{*}$ represents the index of the sample to keep.
-A.6.3 Metropolis-adjusted Langevin algorithm (MALA)
-The Metropolis-Adjusted Langevin Algorithm (MALA) (Rossky et al., 1978a; Zhu & Mumford, 1998), is a Markov Chain Monte Carlo (MCMC) method designed to sample from complex distributions by leveraging gradient information.
-It constructs a proposal mechanism that efficiently explores the target measure by combining stochastic perturbations with information from the target’s local geometry to propose local moves, thereby refining the standard Metropolis-Hastings (MH) algorithm (Hastings, 1970).
-MALA operates on the target (power) posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$.
-The algorithm constructs a sequence of random variables $\{\bm{z}^{(i,\;t_{k})}\}_{i=1}^{\infty}$ that forms a Markov chain converging in distribution to the target measure.
-The target posterior measure, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$, is:
-$\log P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})=\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})^{t_{k}}+\log P(\bm{z}\mid\bm{f})-\text{const},$ (71)
-MALA
-Starting with the initial state of the first chain, $\bm{z}^{(1,\;{t}_{1})}\sim Q(\bm{z})=P(\bm{z}\mid\bm{f})$, the local state within a temperature is updated as follows:
-  1. 1.
-    Langevin diffusion: A new state $\bm{z}^{\prime\;(i,\;t_{k})}$ is proposed for a local chain, operating with a specific $t_{k}$, using a transition kernel inspired by (overdamped) Langevin dynamics (Roberts & Stramer, 2002; Brooks et al., 2011):
-    Target: $\;\,\log\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)\propto\log P(\bm{x}^{(b)}\mid\bm{z},\bm{\Phi})^{t_{k}}+\log P(\bm{z}\mid\bm{f})$
-    Proposal: $\;\,\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})}\sim q\left(\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})}\right)$
-    Kernel: $\;\,q(\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})})=\mathcal{N}\Bigg(\bm{z}^{\prime\;(i,\;t_{k})};\;\bm{z}^{(i,\;t_{k})}+\frac{\eta}{2}\bm{C}\cdot\nabla_{\bm{z}}\log\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right),\;\left(\sqrt{\eta}\cdot\bm{C}^{\frac{1}{2}}\right)^{2}\Bigg),$ (72)
-    where $\eta>0$ is the step size, $\nabla_{\bm{z}}\log\gamma$ is the log-posterior gradient with respect to $\bm{z}^{(i,\;t_{k})}$, and $\bm{C}$ is a pre-conditioning, positive-definite matrix, (for example, the identity matrix, $\bm{I}$).
-  2. 2.
-    MH criterion: Once $N_{\text{unadjusted}}$ iterations have elapsed, Metropolis-Hastings adjustments, (MH) (Metropolis & Ulam, 1949), are introduced. The criterion for the local proposal is:
-    $r_{\text{local}}=\frac{\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)\,q\left(\bm{z}^{(i,\;t_{k})}\mid\bm{z}^{\prime\;(i,\;t_{k})}\right)}{\gamma\left(\bm{z}^{(i,\;t_{k})}\right)\,q\left(\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})}\right)},$ (73)
-  3. 3.
-    Acceptance: The proposal is accepted with probability $\min(1,r_{\text{local}})$.
-If accepted, $\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{\prime\;(i,\;t_{k})}$; otherwise, the current state is retained: $\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{(i,\;t_{k})}$.
-  4. 4.
-    Global Swaps: Global swaps are proposed and accepted subject to the criterion outlined in Eq. 35
-Convergence
-Under mild regularity conditions, the local Markov chain $\{\bm{z}^{(i,\;t_{k})}\}$ generated by MALA converges in distribution to $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$ as $N_{\text{local}}\to\infty$.
-The proposal mechanism in Eq.
-72 ensures that the chain mixes efficiently, particularly in high-dimensional spaces, provided that the step size $\eta$ is tuned appropriately.
-The algorithm provides theoretical guarantees of the target, due to its satisfaction of detailed balance under Eq.
-73:
-$\gamma\left(\bm{z}^{(i,\;t_{k})}\right)\,q\left(\bm{z}^{\prime\;(i,\;t_{k})}\mid\bm{z}^{(i,\;t_{k})}\right)$ $=\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)\,q\left(\bm{z}^{(i,\;t_{k})}\mid\bm{z}^{\prime\;(i,\;t_{k})}\right),$ (74)
-ensuring that the stationary distribution of the local chain matches the local target.
-A.6.4 autoMALA
-More details on the Metropolis-adjusted Langevin Algorithm (MALA) are provided in Sec. A.6.3.
-Selecting an appropriate step size for unaltered MALA presents a challenge, particularly for SE, as each power posterior exhibits distinct local curvature.
-To address this, we adaptively tune the step size with autoMALA (Biron-Lattes et al., 2024), which reinterprets MALA as Hamiltonian Monte Carlo (HMC) with a single leapfrog step (Duane et al., 1987).
-The algorithm operates locally on a target power posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$.
-It constructs a sequence of random variables per temperature, $\{\bm{z}^{(i,\;t_{k})}\}_{i=1}^{\infty}$, that forms a Markov chain converging in distribution to the local target:
-$\log\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)=\log P(\bm{z}^{(i,\;t_{k})}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})\propto\log P(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})},\bm{\Phi})^{t_{k}}+\log P(\bm{z}^{(i,\;t_{k})}\mid\bm{f}).$
-Starting with the initial state of the first chain, $\bm{z}^{(1,\;t_{1})}\sim Q(\bm{z})=P(\bm{z}\mid\bm{f})$, the local state within a temperature is updated as follows:
-  1. 1.
-    Acceptance thresholds: Two bounding acceptance thresholds are sampled uniformly:
-    $a,b$ $\sim\mathcal{U}\Big(\left(a,b\right);\bm{0},\bm{1}\Big),\quad(a,b)\in[0,1]^{2},\quad b>a$ (75)
-  2. 2.
-    Mass matrix: Random pre-conditioning matrices are initialized per latent dimension:
-    $\varepsilon^{(i,\;t_{k})}$ $\sim\text{Beta}_{\;01}\left(1,1;\left[\frac{1}{2},\frac{2}{3}\right]\right),$
-    $M_{p,p}^{1/2\;^{(i,\;t_{k},\;q)}}$ $=\varepsilon^{(i,\;t_{k})}\cdot\Sigma_{p,p}^{-1/2\;{(i,\;t_{k},\;q)}}+(1-\varepsilon^{(i,\;t_{k})}),$ (76)
-    where $\Sigma_{p,p}^{(i,\;t_{k},\;q)}=\text{VAR}_{\,s}\left[\bar{z}_{q,p}^{(i,\;t_{k},\,s)}\right]$.
-Here, $\text{Beta}_{\;01}\left[a,b;c,d\right]$ denotes a zero-one-inflated Beta distribution.
-The parameters $a,b>0$ represent the shape parameters of the Beta distribution, while $c,d\in[0,1]^{2}$ specify the probabilities governing the mixture.
-The positive definite mass matrix, $\bm{M}$, can be related to Eq.
-72 as $\bm{M}=\bm{C}^{-1}$, and is efficient to invert as a diagonal matrix.
-  3. 3.
-    Leapfrog proposal: The following transition is proposed for a local chain, $t_{k}$, with an adaptive step size, $\eta^{(i,t_{k})}$:
-    $\bm{p}^{(i,\;t_{k})}$ $\sim\mathcal{N}\left(\bm{p};\;\bm{0},\bm{M}^{(i,\;t_{k})}\right)$
-    $\bm{p}^{\prime\;(i,\;t_{k})}_{1/2}$ $=\bm{p}^{(i,\;t_{k})}+\frac{\eta^{(i,t_{k})}}{2}\nabla_{\bm{z}}\log\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)$
-    $\bm{z}^{\prime\;(i,\;t_{k})}$ $=\bm{z}^{(i,\;t_{k})}+\eta^{(i,t_{k})}\left(\bm{M}^{-1}\bm{p}^{\prime}_{1/2}\right)^{(i,\;t_{k})}$
-    $\bm{\hat{p}}^{(i,\;t_{k})}$ $=\bm{p}^{\prime\;(i,\;t_{k})}_{1/2}+\frac{\eta^{(i,t_{k})}}{2}\nabla_{\bm{z}}\log\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)$
-    $\bm{p}^{\prime\;(i,\;t_{k})}$ $=-\bm{\hat{p}}^{(i,\;t_{k})},$ (77)
-  4. 4.
-    MH criterion: Once $N_{\text{unadjusted}}$ iterations have elapsed, Metropolis-Hastings (MH) adjustments are introduced. The MH acceptance criterion for the local proposal is:
-    $r_{\text{local}}=\frac{\gamma\left(\bm{z}^{\prime\;(i,\;t_{k})}\right)\,\mathcal{N}\left(\bm{p}^{\prime\;(i,\;t_{k})};\;\bm{0},\bm{M}^{(i,\;t_{k})}\right)}{\gamma\left(\bm{z}^{(i,\;t_{k})}\right)\,\mathcal{N}\left(\bm{p}^{(i,\;t_{k})};\;\bm{0},\bm{M}^{(i,\;t_{k})}\right)},$ (78)
-  5. 5.
-    Step size adaptation: Starting with an initial estimate, $\eta^{(i,t_{k})}_{\text{init}}$, set as the average accepted step size from the previous training iteration, (used as a simple alternative to the round-based tuning algorithm proposed by Biron-Lattes et al.
-(2024)), the step size is adjusted as follows:
-    • If $a<r_{\text{local}}<b$, then $\eta^{(i,t_{k})}=\eta^{(i,t_{k})}_{\text{init}}$
-    • If $r_{\text{local}}\leq a$, then $\eta^{\prime\;(i,t_{k})}=\eta^{\;(i,t_{k})}_{\text{init}}/\Delta\eta^{k}$, where $k\leftarrow k+1$ is chosen recursively until $r_{\text{local}}>a$, and the step size is accepted, $\eta^{(i,t_{k})}=\eta^{\prime\;(i,t_{k})}$
-    • If $r_{\text{local}}\geq b$, then $\eta^{\prime\;(i,t_{k})}=\eta^{(i,t_{k})}_{\text{init}}\cdot\Delta\eta^{k}$, where $k\leftarrow k+1$ is chosen recursively until $r_{\text{local}}<b$, and the step size is accepted, $\eta^{(i,t_{k})}=\eta^{\prime\;(i,t_{k})}/\Delta\eta$
-    Here, $\Delta\eta$ is a tunable hyperparameter that controls the rate of step size adaptation.
-To improve efficiency in our implementation, tuning is prematurely terminated when $\eta^{(i,t_{k})}$ falls out of a user-defined range, $\eta_{\text{min}},\eta_{\text{max}}]$.
-Excessively large or small step sizes are not beneficial, so preventing the algorithm from searching extreme values saves time without compromising performance.
-  6. 6.
-    Reversibility check: If the step size was modified, the reversibility of the update in Eq.
-77 is verified before proceeding with MH acceptance.
-If $\eta^{(i,t_{k})}_{\text{init}}$ cannot be recovered from the reversed step-size adjustment process with the proposed state, $\left\{\bm{z}^{\prime\;(i,\;t_{k})},\;{p}^{\prime\;(i,\;t_{k})},\eta^{(i,t_{k})}\right\}$, the proposal is rejected regardless of the outcome of the MH adjustment.
-  7. 7.
-    Acceptance: The proposal is accepted with probability $\min(1,r_{\text{local}})$.
-If accepted, $\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{\prime\;(i,\;t_{k})}$; otherwise, the current state is retained: $\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{(i,\;t_{k})}$.
-  8. 8.
-    Global Swaps: Global swaps are proposed and accepted subject to Eq. 35
-A.7 Extra image samples
-A.7.1 MNIST
-(a) Real samples
-(b) No ref. prior, RBF basis
-(c) No ref. prior, FFT basis
-(d) Gaussian ref. prior, RBF basis
-(e) Gaussian ref. prior, FFT basis
-(f) Uniform ref. prior, RBF basis
-(g) Uniform ref. prior, $F$ basis
-(h) Lognormal prior, RBF basis
-(i) Lognormal ref. prior, FFT basis
-Figure 12: Generated MNIST examples using different bases and exponentially-tilted priors (Sec. 3.1.1).
-A.7.2 FMNIST
-(a) Real samples
-(b) No ref. prior, RBF basis
-(c) No ref. prior, FFT basis
-(d) Gaussian ref. prior, RBF basis
-(e) Gaussian ref. prior, FFT basis
-(f) Uniform ref. prior, RBF basis
-(g) Uniform ref. prior, $F$ basis
-(h) Lognormal prior, RBF basis
-(i) Lognormal ref. prior, FFT basis
-Figure 13: Generated FMNIST examples using different bases and exponentially-tilted priors (Sec. 3.1.1).
-A.7.3 Darcy flow
-(a) Real samples
-(b) No ref. prior, RBF basis
-(c) No ref. prior, FFT basis
-(d) Gaussian ref. prior, RBF basis
-(e) Gaussian ref. prior, FFT basis
-(f) Uniform ref. prior, RBF basis
-(g) Uniform ref. prior, $F$ basis
-(h) Lognormal prior, RBF basis
-(i) Lognormal ref. prior, FFT basis
-Figure 14: Generated Darcy flow examples using different bases and exponentially-tilted priors (Sec. 3.1.2).
-A.7.4 SVHN
-(a) Maximum Likelihood
-(b) Steppingstone estimator
-Figure 15: Generated SVHN examples by KAEM trained with MLE and SE.
-A.7.5 CelebA
-(a) Maximum Likelihood
-(b) Steppingstone estimator
-Figure 16: Generated CelebA examples by KAEM trained with MLE and SE.
-A.8 Trained latent priors
-The following priors were extracted from trained KAEM models, each corresponding to the first three components of prior $q=1$.
-A.8.1 MNIST RBF
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 17: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 18: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 19: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
-A.8.2 MNIST FFT
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 20: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 21: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 22: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
-A.8.3 FMNIST RBF
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 23: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 24: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 25: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
-A.8.4 FMNIST FFT
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 26: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 27: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 28: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
-A.8.5 Darcy RBF
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 29: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and Gaussian constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 30: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and lognormal constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 31: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and uniform constrained priors.
-A.8.6 Darcy FFT
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 32: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and Gaussian constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 33: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and lognormal constrained priors.
-(a) $q=1,p=1$
-(b) $q=1,p=2$
-(c) $q=1,p=3$
-Figure 34: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and uniform constrained priors.